### v1

# Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

Abstract
We propose denoising diffusion variational inference (DDVI), an approximate inference algorithm for latent variable models which relies on diffusion models as expressive variational posteriors. Our method augments variational posteriors with auxiliary latents, which yields an expressive class of models that perform diffusion in latent space by reversing a user-specified noising process. We fit these models by optimizing a novel lower bound on the marginal likelihood inspired by the wake-sleep algorithm. Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks. When applied to deep latent variable models, our method yields the denoising diffusion VAE (DD-VAE) algorithm. We use this algorithm on a motivating task in biology—inferring latent ancestry from human genomes—outperforming strong baselines on the Thousand Genomes dataset.
1 Introduction
Latent variable methods often rely on variational inference to fit an approximate model of the posterior distribution (Vahdat & Kautz, 2020; Maaløe et al., 2016). The expressivity of this model has a significant impact on the performance of variational inference (Kingma et al., 2016), which motivates research that leverages modern generative models—including normalizing flows (Rezende & Mohamed, 2015) and generative adversarial networks (Goodfellow et al., 2014; Makhzani et al., 2015)—to represent expressive approximate posteriors.
This work seeks to improve variational inference via expressive posteriors based on diffusion models (Ho et al., 2020; Song et al., 2020). Diffusion methods are defined via a noising process, which maps data into Gaussian noise; a diffusion model generates data by reversing this noising process, which yields high-quality samples and accurate density estimates (Kingma et al., 2021). Here, we argue for using diffusion models in latent space, where we gradually map a simple (e.g., Gaussian) latent representation of the data into one that is more complex via an iterative diffusion-like procedure. This procedure yields an expressive approximate posterior trained with a denoising objective that does not involve adversarial training (Makhzani et al., 2015) or constrained invertible architectures (Kingma et al., 2016).
Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm that introduces auxiliary latent variables into the approximate posterior via a user-specified noising process. This process transforms the latent variable we seek to model into a simple (e.g., Gaussian) auxiliary latent; during inference, we fit the approximate posterior by reversing this noising process. Our learning objective is a variational lower bound inspired by the wake-sleep algorithm (Hinton et al., 1995) that can be interpreted as a form of regularized variational inference. We also derive extensions of our method to semi-supervised learning and clustering.
Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks. When applied to deep latent variable models, our method yields the denoising diffusion VAE (DD-VAE) algorithm. We evaluate DD-VAEs on a real problem in biological data analysis—inferring human ancestry from genetic data. Our method outperforms strong baselines on the Thousand Genomes dataset (Siva, 2008) and learns a low-dimensional latent space that preserves semantically meaningful structure (Haghverdi et al., 2015).
Contributions.
In summary, this work introduces denoising diffusion variational inference, an approximate inference algorithm that features three key components: auxiliary latent variables, a user-specified noising process over these variables, and a lower bound on the marginal likelihood inspired by wake-sleep. Our method can be used to fit deep latent variable models, which yields the DD-VAE algorithm. This algorithm is especially effective at dimensionality reduciton and representation learning, where it outperforms alternative methods based on adversarial training.
Figure 1: Denoising diffusion variational inference in a VAE. Between the encoder and decoder, we have a diffusion model to map a simple distribution into a complex distribution over latents.
2 Background
Deep Latent Variable Models
Latent variable models (LVMs) $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ are usually fit by optimizing the evidence lower bound (ELBO)
$\log p_{\bm{\theta}}(\mathbf{x})\leq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}))$
which serves as a tractable surrogate for the marginal log-likelihood (MLL). The gap between the MLL and the ELBO equals precisely $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO (Kingma & Welling, 2013).
Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods (Maaløe et al., 2016), MCMC-based methods (Salimans et al., 2015), normalizing flows (Rezende & Mohamed, 2015)—or improved learning objectives—e.g., adversarial or sample-based losses (Makhzani et al., 2015; Zhao et al., 2017; Si et al., 2022; 2023).
The wake-sleep algorithm (Hinton et al., 1995) optimizes an alternative objective
$\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))$
, in which the KL divergence term is reversed. The learning procedure for wake-sleep involves alternating between ”wake” phases where the recognition model is updated and ”sleep” phases where the generative model is refined.
Denoising Diffusion Models
A diffusion model is defined via a user-specified noising process $q$ that maps data $\mathbf{x}_{0}$ into a sequence of $T$ variables $\mathbf{y}_{1:T}=\mathbf{y}_{1},...,\mathbf{y}_{T}$ that represent increasing levels of corruption to $\mathbf{x}_{0}$. We obtain $\mathbf{y}_{1:T}$ by applying a Markov chain $q(\mathbf{y}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where we define $\mathbf{y}_{0}=\mathbf{x}_{0}$ for convenience. When $\mathbf{x}_{0}$ is a continuous vector, a standard choice of transition kernel is $q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{{\bm{\alpha}}_{t}}\mathbf{y}_{t-1},\sqrt{1-{\bm{\alpha}}_{t}}\mathbf{I})$, which is a Gaussian centered around a copy of $\mathbf{y}_{t-1}$ to which we added noise following a schedule $10<{\bm{\alpha}}_{1}<{\bm{\alpha}}_{2}<...<{\bm{\alpha}}_{T}=$ 1.
A diffusion model can then be represented as a latent variable distribution $p(\mathbf{x}_{0},\mathbf{y}_{1:T})$ that factorizes as $p(\mathbf{x}_{0},\mathbf{y}_{1:T})=p(\mathbf{y}_{T})\prod_{t=0}^{T-1}p_{\bm{\theta}}(\mathbf{y}_{t}\mid\mathbf{y}_{t+1})$ (again using $\mathbf{y}_{0}$ as shorthand for $\mathbf{x}_{0}$). This model seeks to approximate the reverse of the forward diffusion $q$ and map noise $\mathbf{y}_{T}$ into data $\mathbf{x}_{0}$.
The true reverse of the process $q$ cannot be expressed in closed form; as such, we parameterize $p_{\bm{\theta}}$ with parameters ${\bm{\theta}}$ trained by maximizing the ELBO:
$\log p_{\bm{\theta}}(\mathbf{x}_{0})\geq\mathbb{E}_{q}\left[\log p_{\bm{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})-\sum_{t=2}^{T}D_{\mathrm{KL}}(q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})||p_{\bm{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t}))\right]-D_{\mathrm{KL}}(q(\mathbf{x}_{T}|\mathbf{x}_{0})||p(\mathbf{x}_{T}))$
Visualization and Dimensionality Reduction
There exist two important types of dimensionality reduction methods, which (1) emphasize the preservation of pairwise distance structures among all data samples, e.g., PCA (Wold et al., 1987) and LDA (Balakrishnama & Ganapathiraju, 1998), and (2) local distances over global ones, e.g., t-SNE (Van der Maaten & Hinton, 2008) and UMAP (McInnes et al., 2018). Latent variable models (LVMs) (Kingma et al., 2016; Makhzani et al., 2015) are another class of techniques for dimensionality reduction. They represent high-dimensional data in terms of latent, or hidden, variables in a lower-dimensional space, effectively providing a compact representation of the data.
3 Variational Inference With Denoising Diffusion Models
We introduce denoising diffusion variational inference (DDVI), which enhances variational inference with diffusion-based methods and is motivated by challenges in data visualization and dimensionality reduction. Our approach consists in augmenting variational inference in a latent variable model $p(\mathbf{x},\mathbf{z})$ with auxiliary latents $\mathbf{y}\in\mathcal{Y}$ introduced via a user-specified noising process $r(\mathbf{y}|\mathbf{z})$. The $r(\mathbf{y}|\mathbf{z})$ transforms $\mathbf{z}$—which is the latent whose intractable posterior we seek to approximate—into $\mathbf{y}$, whose posterior will be easier to model. Examples of $r$ include forward diffusion processes, discrete noising processes (Austin et al., 2021), as well as custom regularizers (Section 3.3).
We then form an expressive posterior $q(\mathbf{z}|\mathbf{x})$ by fitting the reverse of the noising process $r(\mathbf{y}|\mathbf{z})$ as in a diffusion model. We define $q(\mathbf{z}|\mathbf{x})$ by sampling from a first model $q(\mathbf{y}|\mathbf{x})$—this is an easier task since we can choose $y$ to have a simple (e.g., Gaussian) posterior—and then by sampling from a denoising model $q(\mathbf{z}|\mathbf{x},\mathbf{y})$ that approximates the reverse process $r(\mathbf{z}|\mathbf{y})$. The model $q(\mathbf{z}|\mathbf{x},\mathbf{y})$ is parameterized and trained as a denoising diffusion model, and is thus a flexible posterior estimator.
When diffusion variational inference is used to fit a variational autoencoder, we refer to the resulting algorithm as VAE with diffusion encoders (DD-VAE). We define the full algorithm below.
3.1 DD-VAE: Variational Autoencoders With Diffusion Encoders
We seek to fit a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ with a potentially complex prior $p_{\bm{\theta}}(\mathbf{z})$. One perspective that can be used to define our approach consists in lifting $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ into an extended latent space by introducing auxiliary latents $\mathbf{y}\in\mathcal{Y}$ and applying variational inference in the extended space.
Specifically, we augment the model with the aforementioned user-specified noising process $r(\mathbf{y}|\mathbf{z})$. We use $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})=r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ to denote the extended probability distribution. Note that the marginalizing out $\mathbf{y}$ in $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ yields the original model: hence fitting $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ in an extended probability space solves our original task.
[…]
3.1.1 Denoising Diffusion Variational Inference (DDVI)
One possible way to apply variational inference to $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ is to apply the ELBO twice to obtain:
$\log p_{\bm{\theta}}(\mathbf{x})$ $\geq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y}))$ (1)
$\geq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{y}))]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y}))$ (2)
While Equation 2 is a valid learning objective, it does not yield a training procedure comparable to that of a diffusion model. Diffusion model training involves sampling noisy data from the forward process; here, we sample from the approximate reverse process (from $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$) and seek to match the true process $r(\mathbf{z}|\mathbf{y})$. In practice, we have found the learning signal from this procedure to be too weak to learn a good $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$ that reverses the noising process $r(\mathbf{y}|\mathbf{z})$, as illustrated in Table 7.
Instead, we adopt a learning objective $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ inspired by the wake-sleep algorithm:
$\mathcal{L}$ $=\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / recons. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]}_{\text{sleep term $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$}}$ (3)
Observe that this objective is the ELBO in Equation 2 augmented with an additional regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ that is inspired by the sleep phase of the wake-sleep algorithm. This term consists of the reverse KL divergence $D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))$. Section 3.1.2 shows that this divergence can be optimized via a diffusion-like training procedure. As in wake-sleep, we optimize $\mathcal{L}_{\text{sleep}}$ over ${\bm{\phi}}$ only.
Objective (3) poses restrictions on $p_{\bm{\theta}}$ and $r$. The prior $p_{\bm{\theta}}(\mathbf{y},\mathbf{z})$ needs to have a tractable density, although we will also present approximations and experimental results with implicit sample-based priors. Additionally, $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ must feature tractable entropy and sampling.
Lastly, note that $\mathcal{L}$ lower bounds the marginal log-likelihood $\log p_{\bm{\theta}}(\mathbf{x})$. When $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ equals the true posterior, this bound is tight, since the ELBO is tight, and the sleep term also equals zero.
3.1.2 Optimization Using Wake-Sleep in Latent Space
Next, we introduce optimization algorithms for diffusion variational inference. Maximizing $\mathcal{L}(x,{\bm{\theta}},{\bm{\phi}})$ involves optimizing the sleep term $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$. This optimization is tractable since:
$\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ $=-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})|q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x},\mathbf{z})}[\log[{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}/{p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})}]]$ (4)
$=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})}[\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})$ (5)
$\geq\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})}[\mathbb{E}_{r(\mathbf{y}|\mathbf{z})}[\log(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})/r(\mathbf{y}|\mathbf{z}))]]+\bar{H}(p_{\bm{\theta}})$ (6)
where in Equation 6 we applied the ELBO with $r(\mathbf{y}|\mathbf{z})$ as our choice of variational posterior over the latent $\mathbf{y}$ in the distribution $q_{\bm{\phi}}$; $\bar{H}(p_{\bm{\theta}})$ is the expected conditional entropy of $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on ${\bm{\phi}}$. By maximizing Equation 6, we fit $q(\mathbf{z}|\mathbf{x})$ to $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ while also fitting $q(\mathbf{z}|\mathbf{y})$ to the true reverse denoising process $r(\mathbf{z}|\mathbf{y})$ via the reconstruction term $\mathbb{E}_{r(\mathbf{y}|\mathbf{z})}[\log q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y})]$, which is one of the terms comprising Equation 6. To fit the objective $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$, we propose introducing an additional sleep step in which we sample from $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ in order to compute gradients for $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$.
This procedure mirrors wake-sleep; however, sampling $\mathbf{x}$ from $p_{\bm{\theta}}$ to obtain gradients for the sleep term introduces computational overhead. To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation ^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ of $\mathcal{L}$:
^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ $=\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / reconstr. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))}_{\text{latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$}}.$ (7)
We have replaced $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ with a latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$, in which $\mathbf{x}$ is given, and we only seek to fit the true reverse noising process $r(\mathbf{z}|\mathbf{y})$ independently of $\mathbf{x}$. We can similarly show that
$\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$ $=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})}[\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})\geq\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})r(\mathbf{y}|\mathbf{z})}[\log({q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}/r(\mathbf{y}|\mathbf{z}))]+\bar{H}(p_{\bm{\theta}})$ (8)
$=-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})}[D_{\mathrm{KL}}(r(\mathbf{y}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{z},\mathbf{x}))]-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q(\mathbf{z}|\mathbf{x})),$ (9)
where $\bar{H}(p_{\bm{\theta}})$ is an entropy term constant in ${\bm{\phi}}$. Thus, we minimize the forward KL divergence by sampling $\mathbf{z}$, and applying the noising process to get $\mathbf{y}$; the $q_{\bm{\phi}}$ is fit to denoise $\mathbf{z}$ from $\mathbf{y}$ as in Equation 6.
We optimize our bound on ^$\hat{\mathcal{L}}(x,{\bm{\theta}},{\bm{\phi}})$ end-to-end using minibatch gradient descent over ${\bm{\theta}},{\bm{\phi}}$. While the wake term is a reconstruction loss as in wake-sleep, the sleep term generates latent samples $\mathbf{z},\mathbf{y}$ from $r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ (by analogy with $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ in normal wake-sleep); the denoiser $q_{\bm{\phi}}$ is trained to recover $\mathbf{z}$ from $\mathbf{y}$. Thus, we perform wake-sleep in latent space, which obviates the need for alternating wake and sleep phases, and allows efficient end-to-end training. A limitation of this approximation is that the sleep term does not fit $q_{\bm{\phi}}$ to the true $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x},\mathbf{y})$, and as a consequence $\hat{L}$ is not a tight lower bound on $\log p_{\bm{\theta}}(\mathbf{x})$. We may think of $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$ as a regularizer to the ELBO.
3.1.3 DDVI with Diffusion-Based Encoders
Our framework is naturally instantiated with diffusion models. Let $\mathbf{y}_{0}=\mathbf{z}$ and $\mathbf{y}_{T}=\mathbf{y}$. The forward noising model can be defined as $r(\mathbf{y}_{1:T}|\mathbf{z})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where the $\mathbf{y}_{0:T}$ are an extended set of latent variables that represent increasingly noised versions of $\mathbf{y}_{0}$. We parameterize the approximate reverse diffusion process as $q_{\bm{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}).$
[…]
$\mathcal{L}_{\text{diff}}=$ $\mathbb{E}_{r}\left[\log q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}D_{\mathrm{KL}}(r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0})||q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}))\right]-D_{\mathrm{KL}}(r(\mathbf{y}_{T}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})).$ (10)
This bound is an instantiation of Equation 8 when $r$ is a diffusion process. Many choices of $r,q$ fit this framework. A common type of noising process is Gaussian diffusion, where we define $r(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-{\bm{\alpha}}_{t}}\mathbf{y}_{t-1},{\bm{\alpha}}_{t}\mathbf{I})$ for a suitable schedule $({\bm{\alpha}}_{t})_{t=1}^{T}$. We then adopt the parameterization $q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})=\mathcal{N}(\mathbf{y}_{t-1};\mu_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t),\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t))$. It is then common to parameterize $q_{\bm{\phi}}$ with a noise prediction network ${\bm{\epsilon}}_{\bm{\phi}}$ (Ho et al., 2020); the sum of KL divergences can be approximated by $E_{t,{\bm{\epsilon}}_{t}\sim r(\mathbf{y}_{0},t)}||{\bm{\epsilon}}_{t}-{\bm{\epsilon}}_{\bm{\phi}}(\sqrt{\bar{{\bm{\alpha}}_{t}}}\mathbf{y}_{0}+\sqrt{1-\bar{{\bm{\alpha}}_{t}}}{\bm{\epsilon}}_{t},\mathbf{x},t)||^{2}$. By our earlier argument, this objective encourages $q_{\bm{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})$ to match $r(\mathbf{y}_{0:T})$ as well as the true posterior $p(\mathbf{z}|\mathbf{x})$.
Lastly, we need to show that this choice of $q$ admits a tractable entropy. This follows from
$H(q)$ $=-\sum_{t=1}^{T+1}\mathbb{E}_{q}[\log q(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]=\sum_{t=1}^{T+1}\mathbb{E}_{q}[\frac{d}{2}\left(1+\log(2\pi)\right)+\frac{1}{2}\log|\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x})|]$ (11)
where $d$ is the dimension of $\mathbf{y}$ and we use the notation $\mathbf{y}_{T+1}=\mathbf{x}$. The right-hand term can be approximated using Monte Carlo; it is also common to leave the variance $\Sigma_{\bm{\phi}}$ fixed (as we typically do in our experiments), in which case $H(q)$ is a constant.
3.2 Auxiliary-Variable Diffusion Encoders
Denoising diffusion variational inference can also be understood from the perspective of auxiliary-variable generative models (Maaløe et al., 2016). Our method can be seen as introducing an approximate posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})=\int_{\mathbf{y}}q_{\bm{\phi}}(\mathbf{z},\mathbf{y}|\mathbf{x})$ which itself contains latent variables $\mathbf{y}$.
The standard approach to fit auxiliary-variable generative models is to apply the ELBO twice:
$\log p_{\bm{\theta}}(\mathbf{x})$ $\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))=\text{ELBO}_{\mathbf{z}|\mathbf{x}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ (12)
$\geq\text{ELBO}_{\mathbf{z}|\mathbf{x}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})-\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$ (13)
$=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x},\mathbf{z})r(\mathbf{y}|\mathbf{x},\mathbf{z})-\log q(\mathbf{y},\mathbf{z}|\mathbf{x})],$ (14)
where in Equation 13 we applied the ELBO over $\mathbf{z}$, and in Equation 12 we applied the ELBO again over the latent $\mathbf{y}$ of $q$. The $r(\mathbf{y}|\mathbf{x},\mathbf{z})$ can be interepreted as an approximate variational posterior for the true posterior $q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})$ of the data distribution. Specifically, the gap between Equation 12 and $\log p_{\bm{\theta}}(\mathbf{x})$ is precisely $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))+\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$. Thus, if we correctly match the pairs of approximate posteriors, we will achieve a tight bound.
Additionally, notice that Equation 14 is exactly the ELBO in Equation 2 after rearranging terms. Thus the two perspectives are equivalent: using variational inference to fit an augmented model $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ is equivalent to fitting the original $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ and applying the ELBO to $p_{\bm{\theta}}$, followed by applying the ELBO again to the latent $\mathbf{y}$ in $q$. Since our objective $\mathcal{L}$ is Equation 14) augmented with $\mathcal{L}_{\text{sleep}}$, denoising VI can also be interpreted as fitting an latent-variable approximate posterior $q$. As seen in Equation 9, we implicitly minimize the reverse KL divergence $\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})}[D_{\mathrm{KL}}(r(\mathbf{y}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{z},\mathbf{x}))]$.
Note, however, that there is a crucial difference between our method and auxiliary-variable generative models: we do not optimize the parameters of $r$. Rather, $r$ is a user-defined noising process that acts as a regularizer for the model $q_{\bm{\phi}}$ and whose reverse we seek to fit.
3.3 Regularizing Variational Inference with Denoising Processes
Lastly, our approach can be understood from a third perspective: that of regularization. As we mentioned earlier, $\mathcal{L}$ is the ELBO with an additional regularization term that encourages the approximate posterior $q(\mathbf{z}|\mathbf{x},\mathbf{y})q(\mathbf{y}|\mathbf{x})$ to gradually transform $\mathbf{y}$ into $\mathbf{z}$ following the guidance of $r$. In that sense, $r$ provides data augmentation to ensure that the mapping from $\mathbf{y}$ to $\mathbf{x}$ is easier to learn, similarly to how a diffusion process provides additional learning signal for generating $\mathbf{x}$ from noise.
Our framework also supports the user defining general distributions as regularizers $r_{1}(\mathbf{z},\mathbf{y})$ and $r_{2}(\mathbf{y}|\mathbf{z},\mathbf{x})$. Here, $r_{1}$ puts a custom prior on $\mathbf{z},\mathbf{y}$, while $r_{2}$ specifies a mapping between $\mathbf{y},\mathbf{z}$ that should be followed by the encoder model. We write $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}},r_{1})$ and $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}},r_{2})$, where each term is still defined as in Equation 3 and Equation 6, except using $r_{1},r_{2}$ and we optimize
$\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}},r_{1},r_{2})=\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\beta\cdot\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}},r_{1})+\gamma\cdot\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}},r_{2}),$ (15)
where $\beta,\gamma>0$ (in particular, $\beta$ mirrors the $\beta$-VAE model (Higgins et al., 2016)).
Examples of $r_{2}$ include Gaussian diffusion, discrete noising processes, as well as noise conditioned on $\mathbf{x}$. Note that $r_{2}$ could even be fully deterministic when parameterized with a bijective flow. With a strong regularizer, we could set $q(\mathbf{z}|\mathbf{y},\mathbf{x})$ to equal the reverse $r_{2}(\mathbf{z}|\mathbf{y},\mathbf{x})$ of this flow. Thus, for suitably designed $r_{2}$, we could perfectly match the shape of any mapping. This mirrors the InteL-VAE framework (Miao et al., 2021), with support for flexible and probabilistic regularizers.
Most importantly, our method can be understood as introducing a new form of regularization. While a standard VAE introduces a prior regularization term, we introduce a new form of regularization via a function $r_{2}$. This can be advantageous when specifying or enforcing certain priors is difficult.
4 Extensions
4.1 Semi-Supervised Learning
Following Makhzani et al. (2015), we extend our algorithm to the semi-supervised learning setting where some data points are labeled and we define $p(\mathbf{z},l)=p(\mathbf{z}|c)p(l)$. Then, the model can be specified as $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\bm{\theta}}(\mathbf{z}|l)p(l)$ and the variational distributions are $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\bm{\phi}}(\mathbf{y}|\mathbf{x}),q_{\bm{\phi}}(l|\mathbf{x})$. In this setting, we consider two cases of whether the label can be observed or not (Kingma et al., 2014). We extend Equation 7 to incorporate the label $l$ corresponding to a data point as follows:
$\mathcal{L}_{\text{semi}}=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}|l))-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|l)||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},l))$
When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\bm{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\bm{{\bm{\theta}}}},{\bm{\phi}})+D_{\mathrm{KL}}(q_{\bm{\phi}}(l|\mathbf{x})||p(l))$. Therefore, we can conclude a marginal likelihood on our dataset as follows: ~$\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\bm{{\bm{\theta}}}},{\bm{{\bm{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\bm{{\bm{\theta}}}},{\bm{{\bm{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
We also want to guarantee that all model parameters can be learned in all cases, including $q_{\bm{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference. Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: ~$\tilde{\mathcal{L}}_{\text{semi}_{\bm{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\bm{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\bm{\phi}}(l|\mathbf{x})\right]$
4.2 Clustering
We have further extended our algorithm to encompass the clustering paradigm. We propose two distinct strategies. In the first approach, we simply set $p_{\bm{{\bm{\theta}}}}(\mathbf{z})$ as a mixture of desired priors. The means of these priors are characterized by ${\bm{{\bm{\theta}}}}$. From these means, cluster membership, denoted as $\mathbf{c}$ can be deduced. Remarkably, this approach requires no alteration to the existing learning objective.
Alternatively, the second method retains the original prior but introduces an additional cluster latent variable $\mathbf{c}$ where $\sum_{i}c_{i}=1$. Thus, the model can be specified as $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{c})=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})p(\mathbf{c})$ with $p(\mathbf{c})=Dir({\bm{{\bm{\epsilon}}}})$. Consequently, the variational distributions become $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{c},\mathbf{x}),q_{\bm{\phi}}(\mathbf{y},\mathbf{c}|\mathbf{x})$. This reformulates the learning objective as:
$\mathcal{L}_{\text{clus}}(\mathbf{x})$ $=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z},\mathbf{c}))-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))$
5 Experiments
We compare DD-VAE with vanilla VAE (Kingma & Welling, 2013), IAF-VAE (Kingma et al., 2016), and AAE (Makhzani et al., 2015) on MNIST (Lecun et al., 1998) and CIFAR-10 (Krizhevsky & Hinton, 2009) in unsupervised and semi-supervised learning settings, and also on the 1000 genomes dataset Siva (2008) in clustering settings. The differences between these methods are summarized in Table 8. We discuss the computational costs of all methods in Section A.1. The priors, model architecture, and training details can be founded in Section A.2, Section A.3, and Section A.4 respectively. All results below are reported with 95% confidence interval using 5 different seeds.
5.1 Unsupervised learning
We fit a model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ on the MNIST and CIFAR-10 datasets with three priors $p(\mathbf{z})$: pinwheel, swiss roll, and square and report our results in Table 1 and Table 5. We measure a k-nearest neighbors classification accuracy of the latents (Acc). We also measure latent negative log-likelihood (Latent NLL) by fitting a kernel density estimator (KDE) on the latents produced by the model with test data as input and compute the log-likelihood of the latents sampled from the prior under the fitted KDE.
We consistently see our method perform best on both KNN accuracy and latent NLL, indicating that its output latents are the most informative of labels and best-aligned with the priors. Table 1 reports qualiatitve results that indicate that DD-VAE enforces most accurately the correct prior shape.
Figure 2: Unsupervised visualization on MNIST using three different priors (pinwhell, swiss roll, and square). Each color indicates a class.
Method Pinwheel Swiss Roll Square Acc Latent NLL Acc Latent NLL Acc Latent NLL VAE $0.58\pm 0.01$ $6.18\pm 0.93$ $0.56\pm 0.055$ $82.96\pm 99.30$ $0.35\pm 0.02$ $25.07\pm 62.41$ IAF-VAE $\bm{0.71\pm 0.00}$ $2.06\pm 0.21$ $0.67\pm 0.02$ $12.13\pm 18.28$ $0.65\pm 0.03$ $1.53\pm 0.82$ AAE $0.67\pm 0.02$ $1.94\pm 0.09$ $0.68\pm 0.03$ $3.43\pm 0.14$ $0.57\pm 0.02$ $1.97\pm 0.70$ HVAE w/ flow $\bm{0.72\pm 0.01}$ $2.18\pm 0.13$ $\bm{0.71\pm 0.02}$ $4.17\pm 1.23$ $\bm{0.70\pm 0.01}$ $\bm{1.25\pm 0.08}$ DiffVAE $\bm{0.72\pm 0.01}$ $\bm{1.57\pm 0.10}$ $0.66\pm 0.01$ $\bm{3.19\pm 0.07}$ $\bm{0.72\pm 0.03}$ $\bm{1.24\pm 0.13}$
Table 1: Unsupervised learning on MNIST. We report accuracy using KNN (K=20) classifier and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
5.2 Semi-supervised Learning
We also evaluate the performance of our method and the baselines under semi-supervised learning setting where some labels are observed (1,000 for MNIST and 10,000 for CIFAR-10) and the partitions of the priors are known. We use the same set of priors and baselines. Details on how we partition each prior into $p_{\bm{{\bm{\theta}}}}(\mathbf{z}|l)$ can be founded in Section A.2. The partitions defined for our priors are local parts of the priors. We substitute the KNN accuracy metric with the average class latent NLL, which measures how well the test latents from each class match their respective partition of the prior $p_{\bm{{\bm{\theta}}}}(\mathbf{z}|l)$. In our case, this metric implicitly measures accuracy because the metric requires that the test latents for each class be close to each other, as the partitions defined are local parts of the full prior. The results are shown in Table 2 and Table 6. DD-VAE mostly outperforms the baselines across different priors and metrics. We also show the visualization in Figure 3 where DD-VAE matches the prior almost perfectly.
Figure 3: Semi-supervised visualization on MNIST with 1,000 labels using three different priors (pinwheel, swiss roll, and square). Each a indicates one class.
Method Pinwheel Swiss Roll Square Latent NLL Avg Class Latent NLL Avg Class Latent NLL Avg Class Latent NLL Latent NLL Latent NLL VAE $7.77\pm 1.83$ $15.64\pm 8.79$ $56.91\pm 33.90$ $244.51\pm 94.19$ $1.30\pm 0.03$ $-0.71\pm 0.05$ IAF-VAE $2.71\pm 0.22$ $1.84\pm 0.46$ $8.10\pm 12.17$ $14.59\pm 31.88$ $1.04\pm 0.04$ $-1.01\pm 0.03$ AAE $1.90\pm 0.04$ $0.64\pm 0.11$ $\bm{3.35\pm 0.17}$ $1.86\pm 0.77$ $1.70\pm 0.26$ $2.50\pm 4.70$ DiffVAE $\bm{1.28\pm 0.01}$ $\bm{-0.94\pm 0.01}$ $\bm{3.35\pm 0.10}$ $\bm{1.24\pm 0.09}$ $\bm{0.96\pm 0.03}$ $\bm{-1.20\pm 0.02}$
Table 2: Semi-supervised learning on MNIST (1,000 labels). We report average class latent negative log-likelihood and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
5.3 Clustering and Visualization for Genotype Analysis
Figure 4: Visualization of genotype clusters. A color represents one ethnicity.
In this section, we report results on an real-world task in genome analysis. Visualizing genotype data reveals patterns in the latent ancestry of individuals. We compare DD-VAE against with the three strong clustering baselines using the 1000 Genomes dataset. We also report visualizations from three dimensionality reduction algorithms: PCA, TSNE, and UMAP. For each clustering algorithm, we seek to discover up to 20 clusters. We report quantitative results in terms of cluster purity, cluster completeness, and normalized mutual information (NMI). There is an inherent trade-off between cluster purity completeness. The overall clustering performance can be captured with NMI.
Method Cluster Cluster NMI
Method Purity Completeness
VAE $0.28\pm 0.02$ $\bm{0.78\pm 0.16}$ $0.59\pm 0.08$
IAF-VAE $0.29\pm 0.04$ $\bm{0.73\pm 0.06}$ $0.55\pm 0.06$
AAE $0.37\pm 0.06$ $\bm{0.76\pm 0.11}$ $\bm{0.63\pm 0.02}$
DiffVAE $\bm{0.45\pm 0.03}$ $\bm{0.75\pm 0.05}$ $\bm{0.66\pm 0.04}$
Table 3: Quantitative genotype clustering results.
In Table 3, we see that DD-VAE attains the best performance on cluster purity and NMI. For cluster completeness, VAE and AAE have better means but much larger confidence interval. Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA (Wold et al., 1987), t-SNE (Van der Maaten & Hinton, 2008), and UMAP (McInnes et al., 2018)). The legend of Figure 4 can be founded at Figure 5.
6 Related Work, Discussion, and Conclusion
There are a number of recent works merging VAE and diffusion models. Vahdat et al. (2021); Wehenkel & Louppe (2021) use diffusion with VAE priors. Recent research (Preechakul et al., 2022; Zhang et al., 2022; Wang et al., 2023b) has also melded auto-encoders with diffusion models, focusing on semantically meaningful latents. Cohen et al. (2022) crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution. These works focus on sample quality and directly fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to a complex distribution (e.g., compressed images (Rombach et al., 2022)); our work instead seeks to achieve high sample quality in latent space, and obtains complex $\mathbf{z}$ by diffusion from a simple $\mathbf{y}\sim q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})$.
Latent variable models in general are an attractive alternative to visualization methods like PCA, UMAP, and t-SNE (McInnes et al., 2018; Van der Maaten & Hinton, 2008). Domain-specific knowledge can be injected through the prior, and deep neural networks can be utilized to achieve a more expressive mapping from the data space to the latent space. Nevertheless, downsides of LVMs are that they are more computationally expensive and require careful hyperparameter tuning.
While this paper focuses on applications of DD-VAE to dimensionality reduction and visualization, there exist other tasks for the algorithm, e.g., density estimation or sample quality. Since our learning objective differs from the ELBO (it adds a regularizer), we anticipate gains on models whose training benefits from regularization, but perhaps not on all models. Also, attaining competitive likelihood estimation requires architecture improvements that are orthogonal to this paper. However, our ability to generate diverse samples and achieve class separation in latent space hints at the method’s potential on these tasks. Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making or causal effect estimation (Nguyen & Grover, 2022; Deshpande et al., 2022; Deshpande & Kuleshov, 2023; Rastogi et al., 2023).
Appendix A Appendix
A.1 Computational Cost Analysis
Method NMI values at different wall-clock training times NMI @ 10 min NMI @ 20 min NMI @ 30 min NMI @ 40 min NMI @ 50 min NMI @ 60 min VAE $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ IAF-VAE $0.54$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ AAE $0.61$ $0.57$ $0.57$ $0.57$ $0.57$ $0.57$ DiffVAE (T=5) warm up $0.63$ $0.63$ $0.66$ $0.66$ $0.66$ DiffVAE (T=10) warm up $\mathbf{0.64}$ $\mathbf{0.68}$ $\mathbf{0.70}$ $\mathbf{0.70}$ $\mathbf{0.70}$ DiffVAE (T=20) warm up $0.50$ $0.51$ $0.56$ $0.64$ $0.68$ DiffVAE (T=50) warm up $0.52$ $0.54$ $0.51$ $0.59$ $0.59$
Table 4: Computational cost trade-off on 1kgenome: NMI vs wall-clock training time
We conduct a computational cost analysis between the baselines and DD-VAE with various timesteps on the genotype clustering/visualization experiments. The table Table 4 below shows that DD-VAE outperforms baselines at all timestamps and continues to improve after the baselines have plateaued.
A.2 Priors
Below we describe the sampling process for each prior.
Pinwheel. This distribution was used in (Johnson et al., 2016). We define the number of clusters to be 10. For semi-supervised learning experiments, this prior is partitioned into 10 partitions, each partition being a cluster.
Swiss Roll. This distribution was used in Marsland (2014). We add noise $\sigma=0.1$ to the prior. For semi-supervised learning experiments, this prior is partitioned into 10 partitions. The samples from the prior, before we add noise to it, can actually be characterized by a single scalar representing how far you are long the swiss roll from the center. The paritioning is done by creating 10 equal-length intervals in this 1D space.
Square. This distribution has the shaped of a square going from -1 to 1 in both axes. Each position on the square can be characterized by a single scalar representing how far you are from the top left corner. Sampling is done by sampling the position uniformly and turn the 1D position to 2D latent. We add noise $\sigma=0.06$ to the prior. For semi-supervised learning experiments, this prior is partitioned into 10 partitions. The partitioning is done by creating 10 equal-length intervals in the 1D position space.
VAE and IAF-VAE requires that we can evaluate the prior density. To do this, for all priors, we evaluate the density by fitting a kernel density estimator with mixture of gaussian kernel with bandwidth equal to 0.005, 0.008, 0.01, 0.03, and 0.05.
A.3 Model Architecture
All methods use the same architecture for encoder $q_{\bm{\phi}}(z|x)$ and decoder $p_{\bm{\theta}}(x|z)$, excluding the extra parts specific to each method which we describe below, for the same dataset. For MNIST, the encoder and decoder are multi-layer perceptron with two hidden layers, each with 1000 hidden units. For CIFAR-10, the encoder is a 4-layer convolutional neural network with (16, 32, 64, 128) channels with a linear layer on top, and the decoder is a 4-layer tranposed convolutional neural network with (64, 32, 16, 3) channels where a linear layer is used to first turn the feature dimension from 2 to 64.
IAF-VAE employs 4 IAF transformations on top of the encoder, each is implemented with a 4-layer MADE. The number of hidden units in MADE is 24. The ordering is reversed between every other IAF transformation.
AAE has a discriminator, used in adversarial training, which is a multi-layer perceptron with two hidden layers, each with 1000 hidden units.
DD-VAE has a diffusion model on top of the encoder. The time-conditioned reverse diffusion distribution is implemented with a 5-layer time-conditioned multi-layer perceptron, each with 128 hidden units. A time-conditioned linear layer learns an additional embedding for each timestep and adds it to the output of the linear layer.
A.4 Training Details
For training, we use Adam optimizer and latent size of 2 for all of our experiments. The training details of each algorithm are detailed below:
VAE. The batch size is set to 128. The number of epochs is 200 for unsupervised and clustering experiments and 50 for semi-supervised experiments. The learning rate is 0.0001. The loss is BCE for MNIST and CIFAR-10 experiments and MSE for genotype analysis experiments. For semi-supervised MNIST experiments, the kl divergence weight is set to be 0.01, while for semi-supervised CIFAR-10 experiments, the kl divergence weight is set to be 0.01. For other experiments, the KL divergence weight is set with a schedule linear on number of epochs going from 0 to 0.01. We also have a weight of 5 multiplied to the prior density.
IAF-VAE. The batch size, number of epochs, learning rate, loss, KL divergence weight, and prior density weight are the same as VAE. The context size, i.e., the size of features used to initialize the flow layers for different datat point, is 10.
AAE. The batch size is set to 128. The number of epochs is 200 for all experiments. The learning rate is 0.0002. The loss is MSE for all experiments. To stabilize the training, we add noise to the input to the discriminator with sigma 0.3 at the start and lower it by 0.1 for every 50 epochs. The noise equals to 0 at epoch 150.
DD-VAE. The batch size is set to 128 for most experiments, except for semi-supervied experiments where the batch size is 1024. The number of epochs is 200 for unsupervised and clustering experiments and 30 for semi-supervised experiments. The learning rate is 0.0001. The loss is BCE for MNIST and CIFAR-10 experiments and MSE for genotype analysis experiments. For unsupervised MNIST and CIFAR-10 experiments, the KL divergence weight is set to 0.003. For semi-supervised MNIST experiment, we use KL divergence weight of 0.1. For semi-supervised CIFAR-10 experiment, we use KL divergence weight of 0.5. For clustering experiment, we use KL divergence weight of 0.005. The number of timesteps is 20 for unsupervised and clustering experiments and 100 for semi-supervised experiments.
A.5 Genotype Analysis Experiments Details
Figure 5: Legend showing what ethnicity each color corresponds to in the 1000 Genomes dataset
Before inputting the data points into any of the visualization methods, we first pre-process it by running a PCA and keep only the first 1000 principal components of the data points. We further divide the features by 30 for all latent variables model methods.
The legend of the 1000 Genomes Visualization plot can be found at Figure 5.
Method Pinwheel Swiss Roll Square Acc Latent NLL Acc Latent NLL Acc Latent NLL VAE $0.16\pm 0.01$ $86.9\pm 99.03$ $0.15\pm 0.01$ $360.41\pm 185.87$ $0.15\pm 0.01$ $63.3\pm 70.64$ IAF-VAE $0.19\pm 0.01$ $3.12\pm 1.16$ $0.16\pm 0.01$ $189.39\pm 57.48$ $0.18\pm 0.01$ $0.99\pm 0.11$ AAE $0.23\pm 0.00$ $2.30\pm 0.07$ $\bm{0.23\pm 0.01}$ $\bm{2.84\pm 0.08}$ $0.20\pm 0.01$ $1.46\pm 0.49$ HVAE w/ flow $0.20\pm 0.01$ $2.06\pm 0.13$ $0.20\pm 0.01$ $106.15\pm 15.30$ $0.19\pm 0.01$ $1.05\pm 0.06$ DiffVAE $\bm{0.23\pm 0.01}$ $\bm{1.37\pm 0.02}$ $\bm{0.23\pm 0.01}$ $\bm{2.80\pm 0.11}$ $\bm{0.24\pm 0.01}$ $\bm{0.86\pm 0.05}$
Table 5: Unsupervised learning on CIFAR-10. We report accuracy using KNN (K=20) classifier and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
Method Pinwheel Swiss Roll Square Latent NLL Avg Class Latent NLL Avg Class Latent NLL Avg Class Latent NLL Latent NLL Latent NLL VAE $7.67\pm 1.34$ $9.84\pm 2.87$ $49.64\pm 28.12$ $128.46\pm 47.08$ $1.51\pm 0.23$ $0.65\pm 0.37$ IAF-VAE $4.54\pm 0.59$ $4.46\pm 0.99$ $64.27\pm 73.87$ $102.16\pm 106.36$ $1.04\pm 0.07$ $0.07\pm 0.09$ AAE $2.33\pm 0.02$ $2.42\pm 0.06$ $3.76\pm 0.04$ $4.23\pm 0.13$ $1.45\pm 0.12$ $1.31\pm 0.20$ DiffVAE $\bm{1.49\pm 0.09}$ $\bm{0.35\pm 0.18}$ $\bm{3.35\pm 0.27}$ $\bm{2.28\pm 0.39}$ $\bm{0.97\pm 0.03}$ $\bm{-0.29\pm 0.04}$
Table 6: Semi-supervised learning on CIFAR-10 (10,000 labels). We report average class latent negative log-likelihood and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square VAE $6.18\pm 0.93$ $82.96\pm 99.30$ $25.07\pm 62.41$ IAF-VAE $2.06\pm 0.21$ $12.13\pm 18.28$ $1.53\pm 0.82$ AAE $1.94\pm 0.09$ $3.43\pm 0.14$ $1.97\pm 0.70$ HVAE w/ flow $2.18\pm 0.13$ $4.17\pm 1.23$ $\mathbf{1.25\pm 0.08}$ DiffVAE $\mathbf{1.57\pm 0.10}$ $\mathbf{3.19\pm 0.07}$ $\mathbf{1.24\pm 0.13}$ VAE with Diffusion (standard ELBO) $6.09$ $510.53$ $265.25$
Table 7: Unsupervised learning on MNIST, including the failed results of VAE with Diffusion
Model Training Objective Approximating Family Sample-based Prior Auxiliary Variable Tasks Simplified Graphical Illustration VAE ELBO Diagonal Gaussian No No Density estimation $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ IAF-VAE ELBO Normalizing flow No Yes Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{x}$ AAE Adversarial training Adversarial generator Yes No Visualization $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ HVAE w/o and w/ flow ELBO Factorial Normal / Normalizing flow No Yes Density estimation / High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ ADGM ELBO Non-Gaussian No Yes Density estimation $\mathbf{x}\rightarrow\mathbf{a}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ LDM ELBO Diagonal Gaussian No Yes High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ LSGM ELBO+score matching Diagonal Gaussian No Yes High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ DiffVAE ELBO+sleep term Denoising diffusion Yes Yes Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{T}(\mathbf{y})\rightarrow\mathbf{z}_{0}(\mathbf{z})\rightarrow\mathbf{x}$
Table 8: Comparison of DiffVAE to other relevant methods

### v3

@@
-Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors
+Denoising Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors
 Abstract
-We propose denoising diffusion variational inference (DDVI), an approximate inference algorithm for latent variable models which relies on diffusion models as expressive variational posteriors.
-Our method augments variational posteriors with auxiliary latents, which yields an expressive class of models that perform diffusion in latent space by reversing a user-specified noising process.
-We fit these models by optimizing a novel lower bound on the marginal likelihood inspired by the wake-sleep algorithm.
+We propose denoising diffusion variational inference (DDVI), a black-box variational inference algorithm for latent variable models which relies on diffusion models as flexible approximate posteriors.
+Specifically, our method introduces an expressive class of diffusion-based variational posteriors that perform iterative refinement in latent space; we train these posteriors with a novel regularized evidence lower bound (ELBO) on the marginal likelihood inspired by the wake-sleep algorithm.
 Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks.
-When applied to deep latent variable models, our method yields the denoising diffusion VAE (DD-VAE) algorithm.
-We use this algorithm on a motivating task in biology—inferring latent ancestry from human genomes—outperforming strong baselines on the Thousand Genomes dataset.
+We find that DDVI improves inference and learning in deep latent variable models across common benchmarks as well as on a motivating task in biology—inferring latent ancestry from human genomes—where it outperforms strong baselines on the Thousand Genomes dataset.
 1 Introduction
-Latent variable methods often rely on variational inference to fit an approximate model of the posterior distribution (Vahdat & Kautz, 2020; Maaløe et al., 2016).
-The expressivity of this model has a significant impact on the performance of variational inference (Kingma et al., 2016), which motivates research that leverages modern generative models—including normalizing flows (Rezende & Mohamed, 2015) and generative adversarial networks (Goodfellow et al., 2014; Makhzani et al., 2015)—to represent expressive approximate posteriors.
-This work seeks to improve variational inference via expressive posteriors based on diffusion models (Ho et al., 2020; Song et al., 2020).
-Diffusion methods are defined via a noising process, which maps data into Gaussian noise; a diffusion model generates data by reversing this noising process, which yields high-quality samples and accurate density estimates (Kingma et al., 2021).
-Here, we argue for using diffusion models in latent space, where we gradually map a simple (e.g., Gaussian) latent representation of the data into one that is more complex via an iterative diffusion-like procedure.
-This procedure yields an expressive approximate posterior trained with a denoising objective that does not involve adversarial training (Makhzani et al., 2015) or constrained invertible architectures (Kingma et al., 2016).
-Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm that introduces auxiliary latent variables into the approximate posterior via a user-specified noising process.
-This process transforms the latent variable we seek to model into a simple (e.g., Gaussian) auxiliary latent; during inference, we fit the approximate posterior by reversing this noising process.
-Our learning objective is a variational lower bound inspired by the wake-sleep algorithm (Hinton et al., 1995) that can be interpreted as a form of regularized variational inference.
+We are interested in amortized black-box variational inference problems of the form
+$\log\;p_{\bm{\theta}}(\mathbf{x})\geq\max_{\bm{\phi}}\;\;\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\;p_{\bm{\theta}}(\mathbf{x},\mathbf{z})-\log\;q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})\right]:=\max_{\bm{\phi}}\;\;\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}}),$ (1)
+in which we approximate the marginal likelihood $\log\;p_{\bm{\theta}}(\mathbf{x})$ of a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ with an evidence lower bound $\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ that is a function of an approximate variational posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$.
+We assume that $p_{\bm{\theta}}$ factorizes as $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\bm{\theta}}$ include Bayesian networks, topic models [4], variational autoencoders (VAEs), and broad classes of $p_{\bm{\theta}}$ defined via modern probabilistic programming frameworks [9].
+Maximizing $\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ over ${\bm{\phi}}$ yields a variational posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ that approximates $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ as well as a tight bound on $\log\;p_{\bm{\theta}}(\mathbf{x})$ that serves as a learning objective for $p_{\bm{\theta}}$.
+The approximation gap $\log\;p_{\bm{\theta}}(\mathbf{x})-\max_{\bm{\phi}}\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ equals precisely $\min_{\bm{\phi}}\text{KL}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$, which motivates the design of expressive classes of posteriors $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ that reduce this gap.
+Recent efforts leverage modern generative models—including normalizing flows [28, 16] and generative adversarial networks [8, 20]—as expressive model families for $q_{\bm{\phi}}$ that tighten the ELBO.
+This work seeks to further improve variational inference via expressive posteriors based on diffusion models [12, 37].
+Diffusion methods have become the de-facto standard for high-quality image synthesis [30].
+Here, we use diffusion in latent space to parameterize $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$.
+We train this distribution with a denoising diffusion-like objective that does not involve adversarial training [20] or constrained invertible normalizing flow architectures [16].
+Samples from $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ are obtained via iterative refinement of $\mathbf{z}$, starting from a Gaussian distribution, and gradually forming one that is multi-modal and complex.
+[…]
+Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion as well as a learning objective inspired by the wake-sleep algorithm [11] that can be interpreted as a regularized variational inference.
 We also derive extensions of our method to semi-supervised learning and clustering.
 Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks.
-When applied to deep latent variable models, our method yields the denoising diffusion VAE (DD-VAE) algorithm.
-We evaluate DD-VAEs on a real problem in biological data analysis—inferring human ancestry from genetic data.
-Our method outperforms strong baselines on the Thousand Genomes dataset (Siva, 2008) and learns a low-dimensional latent space that preserves semantically meaningful structure (Haghverdi et al., 2015).
+We evaluate DDVI on synthetic benchmarks and on a real problem in biological data analysis—inferring human ancestry from genetic data.
+Our method outperforms strong baselines on the Thousand Genomes dataset [35] and learns a low-dimensional latent space that preserves biologically meaningful structure [10].
 Contributions.
-In summary, this work introduces denoising diffusion variational inference, an approximate inference algorithm that features three key components: auxiliary latent variables, a user-specified noising process over these variables, and a lower bound on the marginal likelihood inspired by wake-sleep.
-Our method can be used to fit deep latent variable models, which yields the DD-VAE algorithm.
-This algorithm is especially effective at dimensionality reduciton and representation learning, where it outperforms alternative methods based on adversarial training.
+In summary, this work introduces denoising diffusion variational inference, an approximate inference algorithm defined by two components: a class of approximate posteriors $q(\mathbf{z}|\mathbf{x})$ parameterized by diffusion, and a lower bound on the marginal likelihood inspired by wake-sleep.
+We complement DDVI with extensions to semi-supervised learning and clustering.
+Our method is especially suited for probabilistic programming, representation learning, and dimensionality reduction, where it outperforms alternative methods based on normalizing flows and adversarial training.
 Figure 1: Denoising diffusion variational inference in a VAE. Between the encoder and decoder, we have a diffusion model to map a simple distribution into a complex distribution over latents.
@@
 Latent variable models (LVMs) $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ are usually fit by optimizing the evidence lower bound (ELBO)
-$\log p_{\bm{\theta}}(\mathbf{x})\leq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}))$
+$\log p_{\bm{\theta}}(\mathbf{x})\geq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z})),$ (2)
 which serves as a tractable surrogate for the marginal log-likelihood (MLL).
-The gap between the MLL and the ELBO equals precisely $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO (Kingma & Welling, 2013).
-Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods (Maaløe et al., 2016), MCMC-based methods (Salimans et al., 2015), normalizing flows (Rezende & Mohamed, 2015)—or improved learning objectives—e.g., adversarial or sample-based losses (Makhzani et al., 2015; Zhao et al., 2017; Si et al., 2022; 2023).
-The wake-sleep algorithm (Hinton et al., 1995) optimizes an alternative objective
-$\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))$
-, in which the KL divergence term is reversed.
-The learning procedure for wake-sleep involves alternating between ”wake” phases where the recognition model is updated and ”sleep” phases where the generative model is refined.
+The gap between the MLL and the ELBO equals precisely $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO [14].
+Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods [19], MCMC-based methods [31], normalizing flows [28]—or improved learning objectives—e.g., adversarial or sample-based losses [20, 48, 33, 34].
+The wake-sleep algorithm [11] optimizes an alternative objective
+$\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})),$ (3)
+in which the KL divergence term is reversed.
+The learning procedure for wake-sleep involves alternating between "wake" phases where the recognition model is updated and "sleep" phases where the generative model is refined.
 Denoising Diffusion Models
@@
 This model seeks to approximate the reverse of the forward diffusion $q$ and map noise $\mathbf{y}_{T}$ into data $\mathbf{x}_{0}$.
-The true reverse of the process $q$ cannot be expressed in closed form; as such, we parameterize $p_{\bm{\theta}}$ with parameters ${\bm{\theta}}$ trained by maximizing the ELBO:
-$\log p_{\bm{\theta}}(\mathbf{x}_{0})\geq\mathbb{E}_{q}\left[\log p_{\bm{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})-\sum_{t=2}^{T}D_{\mathrm{KL}}(q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})||p_{\bm{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t}))\right]-D_{\mathrm{KL}}(q(\mathbf{x}_{T}|\mathbf{x}_{0})||p(\mathbf{x}_{T}))$
-Visualization and Dimensionality Reduction
-There exist two important types of dimensionality reduction methods, which (1) emphasize the preservation of pairwise distance structures among all data samples, e.g., PCA (Wold et al., 1987) and LDA (Balakrishnama & Ganapathiraju, 1998), and (2) local distances over global ones, e.g., t-SNE (Van der Maaten & Hinton, 2008) and UMAP (McInnes et al., 2018).
-Latent variable models (LVMs) (Kingma et al., 2016; Makhzani et al., 2015) are another class of techniques for dimensionality reduction.
-They represent high-dimensional data in terms of latent, or hidden, variables in a lower-dimensional space, effectively providing a compact representation of the data.
+The true reverse of the process $q$ cannot be expressed in closed form; as such, we parameterize $p_{\bm{\theta}}$ with ${\bm{\theta}}$ trained by maximizing the ELBO:
+$\log p_{\bm{\theta}}(\mathbf{x}_{0})\geq\mathbb{E}_{q}\Bigr{[}\log p_{\bm{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})-\sum_{t=2}^{T}D_{\mathrm{KL}}(q_{t}||p_{t})\Bigr{]}-D_{\mathrm{KL}}(q(\mathbf{x}_{T}|\mathbf{x}_{0})||p(\mathbf{x}_{T}))$ (4)
+where $q_{t},p_{t}$ denote the distributions $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ and $p_{\bm{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, respectively.
 3 Variational Inference With Denoising Diffusion Models
-We introduce denoising diffusion variational inference (DDVI), which enhances variational inference with diffusion-based methods and is motivated by challenges in data visualization and dimensionality reduction.
-Our approach consists in augmenting variational inference in a latent variable model $p(\mathbf{x},\mathbf{z})$ with auxiliary latents $\mathbf{y}\in\mathcal{Y}$ introduced via a user-specified noising process $r(\mathbf{y}|\mathbf{z})$.
-The $r(\mathbf{y}|\mathbf{z})$ transforms $\mathbf{z}$—which is the latent whose intractable posterior we seek to approximate—into $\mathbf{y}$, whose posterior will be easier to model.
-Examples of $r$ include forward diffusion processes, discrete noising processes (Austin et al., 2021), as well as custom regularizers (Section 3.3).
-We then form an expressive posterior $q(\mathbf{z}|\mathbf{x})$ by fitting the reverse of the noising process $r(\mathbf{y}|\mathbf{z})$ as in a diffusion model.
-We define $q(\mathbf{z}|\mathbf{x})$ by sampling from a first model $q(\mathbf{y}|\mathbf{x})$—this is an easier task since we can choose $y$ to have a simple (e.g., Gaussian) posterior—and then by sampling from a denoising model $q(\mathbf{z}|\mathbf{x},\mathbf{y})$ that approximates the reverse process $r(\mathbf{z}|\mathbf{y})$.
-The model $q(\mathbf{z}|\mathbf{x},\mathbf{y})$ is parameterized and trained as a denoising diffusion model, and is thus a flexible posterior estimator.
-When diffusion variational inference is used to fit a variational autoencoder, we refer to the resulting algorithm as VAE with diffusion encoders (DD-VAE). We define the full algorithm below.
-3.1 DD-VAE: Variational Autoencoders With Diffusion Encoders
-We seek to fit a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ with a potentially complex prior $p_{\bm{\theta}}(\mathbf{z})$.
-One perspective that can be used to define our approach consists in lifting $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ into an extended latent space by introducing auxiliary latents $\mathbf{y}\in\mathcal{Y}$ and applying variational inference in the extended space.
-Specifically, we augment the model with the aforementioned user-specified noising process $r(\mathbf{y}|\mathbf{z})$.
-We use $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})=r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ to denote the extended probability distribution.
-Note that the marginalizing out $\mathbf{y}$ in $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ yields the original model: hence fitting $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ in an extended probability space solves our original task.
+We introduce denoising diffusion variational inference (DDVI), which improves variational inference with diffusion-based techniques.
+The goal of DDVI is to fit a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$.
+We assume that $p_{\bm{\theta}}$ factorizes as $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\bm{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) [14].
+Our approach is comprised of two key components: a family of approximate posteriors $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ based on diffusion and a learning objective that can be seen as a regularized ELBO.
+The $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ iteratively refines latents $\mathbf{z}$, starting from a Gaussian distribution.
+The learning objective trains $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to reverse a used-specified forward diffusion process.
+3.1 Diffusion-Based Variational Posteriors
 […]
-3.1.1 Denoising Diffusion Variational Inference (DDVI)
-One possible way to apply variational inference to $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ is to apply the ELBO twice to obtain:
-$\log p_{\bm{\theta}}(\mathbf{x})$ $\geq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y}))$ (1)
-$\geq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{y}))]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y}))$ (2)
-While Equation 2 is a valid learning objective, it does not yield a training procedure comparable to that of a diffusion model.
-Diffusion model training involves sampling noisy data from the forward process; here, we sample from the approximate reverse process (from $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$) and seek to match the true process $r(\mathbf{z}|\mathbf{y})$.
-In practice, we have found the learning signal from this procedure to be too weak to learn a good $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$ that reverses the noising process $r(\mathbf{y}|\mathbf{z})$, as illustrated in Table 7.
-Instead, we adopt a learning objective $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ inspired by the wake-sleep algorithm:
-$\mathcal{L}$ $=\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / recons.
-term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]}_{\text{sleep term $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$}}$ (3)
-Observe that this objective is the ELBO in Equation 2 augmented with an additional regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ that is inspired by the sleep phase of the wake-sleep algorithm.
-This term consists of the reverse KL divergence $D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))$.
-Section 3.1.2 shows that this divergence can be optimized via a diffusion-like training procedure.
-As in wake-sleep, we optimize $\mathcal{L}_{\text{sleep}}$ over ${\bm{\phi}}$ only.
-Objective (3) poses restrictions on $p_{\bm{\theta}}$ and $r$.
-The prior $p_{\bm{\theta}}(\mathbf{y},\mathbf{z})$ needs to have a tractable density, although we will also present approximations and experimental results with implicit sample-based priors.
-Additionally, $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ must feature tractable entropy and sampling.
-Lastly, note that $\mathcal{L}$ lower bounds the marginal log-likelihood $\log p_{\bm{\theta}}(\mathbf{x})$.
-When $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ equals the true posterior, this bound is tight, since the ELBO is tight, and the sleep term also equals zero.
-3.1.2 Optimization Using Wake-Sleep in Latent Space
-Next, we introduce optimization algorithms for diffusion variational inference.
-Maximizing $\mathcal{L}(x,{\bm{\theta}},{\bm{\phi}})$ involves optimizing the sleep term $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$.
-This optimization is tractable since:
-$\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ $=-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})|q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x},\mathbf{z})}[\log[{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}/{p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})}]]$ (4)
-$=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})}[\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})$ (5)
-$\geq\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})}[\mathbb{E}_{r(\mathbf{y}|\mathbf{z})}[\log(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})/r(\mathbf{y}|\mathbf{z}))]]+\bar{H}(p_{\bm{\theta}})$ (6)
-where in Equation 6 we applied the ELBO with $r(\mathbf{y}|\mathbf{z})$ as our choice of variational posterior over the latent $\mathbf{y}$ in the distribution $q_{\bm{\phi}}$; $\bar{H}(p_{\bm{\theta}})$ is the expected conditional entropy of $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on ${\bm{\phi}}$.
-By maximizing Equation 6, we fit $q(\mathbf{z}|\mathbf{x})$ to $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ while also fitting $q(\mathbf{z}|\mathbf{y})$ to the true reverse denoising process $r(\mathbf{z}|\mathbf{y})$ via the reconstruction term $\mathbb{E}_{r(\mathbf{y}|\mathbf{z})}[\log q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y})]$, which is one of the terms comprising Equation 6.
-To fit the objective $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$, we propose introducing an additional sleep step in which we sample from $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ in order to compute gradients for $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$.
-This procedure mirrors wake-sleep; however, sampling $\mathbf{x}$ from $p_{\bm{\theta}}$ to obtain gradients for the sleep term introduces computational overhead.
+We choose the latent $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ to be a vector of $T$ variables that represent progressively simplified versions of $\mathbf{z}$, with $\mathbf{y}_{T}$ corresponding to a simple distribution (e.g., a Gaussian).
+The model $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\bm{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$ transforms $\mathbf{y}_{T}$ into $\mathbf{z}$ via iterative refinement.
+To sample from $q_{\bm{\phi}}$, we first sample $\mathbf{y}_{T}$—this is an easier task since we can define $\mathbf{y}_{T}$ to have a simple (e.g., Gaussian) distribution—and then by sampling from the denoising model $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\bm{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$.
+We define the relationship between $\mathbf{y}$ and $\mathbf{z}$ via a forward diffusion process $r(\mathbf{y}|\mathbf{z})=r(\mathbf{y}_{1:T}|\mathbf{z})=r(\mathbf{y}_{1}|\mathbf{z})\prod_{t=1}^{T-1}r(\mathbf{y}_{t+1}|\mathbf{y}_{t})$, which transforms $\mathbf{z}$—the latent whose intractable posterior we seek to approximate—into $\mathbf{y}_{T}$, whose posterior is easier to model.
+Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes [2].
+The model $q_{\bm{\phi}}$ is trained to approximately reverse this forward diffusion process.
+3.2 The Wake-Sleep Regularized ELBO
+The standard approach to fit auxiliary-variable generative models [19] is to apply the ELBO twice:
+$\log p_{\bm{\theta}}(\mathbf{x})$ $\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$ (5)
+$\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))-\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$ (6)
+$=~{}\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$ (7)
+In Equation 5, we applied the ELBO over $\mathbf{z}$, and in Equation 6 we applied the ELBO again over the latent $\mathbf{y}$ of $q$.
+Notice that the gap between the ELBO and $\log p_{\bm{\theta}}(\mathbf{x})$ is $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))+\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$.
+Thus, if we correctly match $q$ and $r$, we will achieve a tight bound.
+Wake-Sleep Regularization
+Notice that optimizing $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{z},\mathbf{x})p(\mathbf{z}))$ in Equation 7 involves sampling from the approximate reverse process $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ to match the true reverse process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$: this is the opposite of diffusion training, where we would sample from $r$ to fit $q$.
+We have found the standard ELBO to insufficient to learn a good $q_{\bm{\phi}}$ that reverses the noising process $r$, as illustrated in Table 8.
+Instead, our strategy is to introduce a modified ELBO that yields diffusion-like training.
+This objective is the ELBO in Equation 7 augmented with an additional regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$.
+$\log p_{\bm{\theta}}(\mathbf{x})\geq\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / recons.
+term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]}_{\text{sleep term $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$}}$ (8)
+The optimization of the regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ is similar to the sleep phase of the wake-sleep algorithm, and closely resembles diffusion model training (see below).
+As in wake-sleep, $\mathcal{L}_{\text{sleep}}$ is optimized over ${\bm{\phi}}$ only, and the $\mathbf{x}$ are sampled from the model.
+3.3 Optimizing the Regularized ELBO
+Computing $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ still involves intractable distributions $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}),q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$.
+To optimize $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$, we introduce another lower bound $\mathcal{L}_{\text{diff}}({\bm{\phi}})$, which we call the denoising diffusion loss (for reasons that will become apparent shortly):
+$\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ $=-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x},\mathbf{z})}[\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})$ (9)
+$\geq\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x},\mathbf{z})}[\mathbb{E}_{r}[\log\frac{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{z},\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})=\mathcal{L}_{\text{diff}}({\bm{\phi}})$ (10)
+In Equation 10, we applied the ELBO with $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ playing the role of the variational posterior over the latent $\mathbf{y}$ in $q_{\bm{\phi}}$; $\bar{H}(p_{\bm{\theta}})$ is the expected conditional entropy of $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on ${\bm{\phi}}$.
+We define the DDVI learning objective $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ as the sum of the aforementioned terms:
+$\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})=\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\mathcal{L}_{\text{diff}}({\bm{\phi}})$ (11)
+Terms $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{diff}}$ may be weighted by hyper-parameters $\beta_{\text{reg}},\beta_{\text{diff}}>0$, as in the $\beta$-VAE framework.
+In our experiments, assume $\beta_{\text{reg}}=\beta_{\text{diff}}=1$ unless otherwise specified.
+Note that since $\mathcal{L}_{\text{diff}}\leq\mathcal{L}_{\text{sleep}}\leq 0$, $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ is a valid lower bound on $\log p_{\bm{\theta}}(\mathbf{x})$ that is tight when $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})=p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$.
+Wake-Sleep Optimization
+We optimize our bound on $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ using gradient descent by alternating between ELBO optimization and taking sleep steps (see Section 5).
+Note that by maximizing $\mathcal{L}_{\text{diff}}$, we fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ via the forward KL divergence; similarly, by optimizing $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{reg}}$ (the ELBO), we fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ via the reverse KL divergence.
+Thus, optimizing $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ encourages $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to approximate $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, and when the two are equal, the bound on $\log p_{\bm{\theta}}(\mathbf{x})$ is tight.
+Simplifying Wake-Sleep
+We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$.
+In this case, $\mathcal{L}_{\text{diff}}$ requires only sampling from $p(\mathbf{z})$, and the entire loss $\mathcal{L}$ can be optimized end-to-end using gradient descent.
+This algorithm is a simpler (there is no separate sleep phase); however, $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may not perfectly approximate $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, hence $\mathcal{L}$ may no longer be a tight bound.
+We report results in Appendix B.
+3.4 Combining The Wake-Sleep Regularized ELBO With Diffusion Models
+Lastly, we can further simplify $\mathcal{L}_{\text{diff}}$ by leveraging the Markov structure of the forward and reverse processes $r,q$.
+Recall that each $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ can be a vector of $T$ latents, which we also denote as $\mathbf{y}_{1:T}$, and that $r(\mathbf{y}_{1:T}|\mathbf{z})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where $\mathbf{y}_{0}=\mathbf{z}$ and the $\mathbf{y}_{1:T}$ are increasingly noised versions of $\mathbf{y}_{0}$.
+Similarly, $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ is an approximate reverse diffusion process $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ that is trained to reverse $r$.
+In order to fit $q_{\bm{\phi}}$, we form a lower bound $\mathcal{L}_{\text{diff}}(\mathbf{x},{\bm{\phi}})$ on the sleep term $\mathbb{E}_{p(\mathbf{x},\mathbf{z})}\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}$ in Equation 9.
+This bound is identical to the ELBO of a diffusion model, and has the same derivation.
+$\mathcal{L}_{\text{diff}}=$ $~{}\mathbb{E}_{r}\left[\log q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}D_{\mathrm{KL}}(r_{t}||q_{t})\right]-D_{\mathrm{KL}}(r(\mathbf{y}_{T}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})).$ (12)
+where $r_{t},q_{t}$ denote the distributions $r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0})$ and $q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$.
+Parameterizing Diffusion-Based Encoders
+A common type of noising process compatible with this bound when $\mathbf{z}$ is continuous is Gaussian diffusion, where we define $r(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-{\bm{\alpha}}_{t}}\mathbf{y}_{t-1},{\bm{\alpha}}_{t}\mathbf{I})$ for a suitable schedule $({\bm{\alpha}}_{t})_{t=1}^{T}$.
+We then adopt the parameterization $q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})=\mathcal{N}(\mathbf{y}_{t-1};\mu_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t),\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t))$.
+It is then common to parameterize $q_{\bm{\phi}}$ with a noise prediction network ${\bm{\epsilon}}_{\bm{\phi}}$ [12]; the sum of KL divergences can be approximated by $\mathbb{E}_{t,{\bm{\epsilon}}_{t}\sim r(\mathbf{y}_{0},t)}||{\bm{\epsilon}}_{t}-{\bm{\epsilon}}_{\bm{\phi}}(\sqrt{\bar{{\bm{\alpha}}_{t}}}\mathbf{y}_{0}+\sqrt{1-\bar{{\bm{\alpha}}_{t}}}{\bm{\epsilon}}_{t},\mathbf{x},t)||^{2}$.
+Lastly, in order to use diffusion-based encoders with our ELBO-based objective, we need to show that we can tractably compute $-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$, which we equivalently rewrite as:
+$-D_{\mathrm{KL}}(q_{\bm{\phi}}||r\cdot p)=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log(r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))]+H(q).$
+The term $H(q)$ denotes the entropy. We approximate the first term using Monte-Carlo; the entropy is computed as
+$H(q)=-\sum_{t=1}^{T+1}\mathbb{E}_{q}[\log q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]=\sum_{t=1}^{T+1}\mathbb{E}_{q}\left[\frac{d}{2}\left(1+\log(2\pi)\right)+\frac{1}{2}\log\left|\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x})\right|\right]$ (13)
+where $d$ is the dimension of $\mathbf{y}$ and we use the notation $\mathbf{y}_{T+1}=\mathbf{x}$.
+The right-hand term can be approximated using Monte Carlo; it is also common to leave the variance $\Sigma_{\bm{\phi}}$ fixed (as we typically do in our experiments), in which case $H(q)$ is a constant.
+4 Extensions
+4.1 Semi-Supervised Learning
+Following Makhzani et al.
+[20], we extend our algorithm to the semi-supervised learning setting where some data points have labels denoted by $l$.
+We assume the user provides a model of the form $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\bm{\theta}}(\mathbf{z}|l)p(l)$; we set the variational distributions to $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\bm{\phi}}(\mathbf{y}|\mathbf{x}),q_{\bm{\phi}}(l|\mathbf{x})$.
+In this setting, we consider two cases, depending on whether the label is observed [15].
+We extend Equation 8 to incorporate the label $l$ corresponding to a data point as follows:
+$\mathcal{L}_{\text{semi}}=$ $~{}\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}|l))$ (14)
+$-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}\left[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x},l)||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},l))\right]$
+When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\bm{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\bm{{\bm{\theta}}}},{\bm{\phi}})+D_{\mathrm{KL}}(q_{\bm{\phi}}(l|\mathbf{x})||p(l))$.
+Therefore, we can conclude a marginal likelihood on our dataset as follows: ~$\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\bm{{\bm{\theta}}}},{\bm{{\bm{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\bm{{\bm{\theta}}}},{\bm{{\bm{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
+We also want to guarantee that all model parameters can be learned in all cases, including $q_{\bm{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference.
+Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: ~$\tilde{\mathcal{L}}_{\text{semi}_{\bm{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\bm{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\bm{\phi}}(l|\mathbf{x})\right]$
+4.2 Clustering
+We have further extended our algorithm to encompass the clustering paradigm.
+We propose two distinct strategies.
+In the first approach, we simply fir a model in which $p_{\bm{{\bm{\theta}}}}(\mathbf{z})$ as a mixture of desired priors.
+The means of these priors are characterized by ${\bm{{\bm{\theta}}}}$.
+From these means, cluster membership, denoted as $\mathbf{c}$ can be deduced.
+This approach requires no alteration to the existing learning objective.
+Alternatively, the second method retains the original prior but introduces an additional cluster latent variable $\mathbf{c}$ where $\sum_{i}c_{i}=1$.
+Thus, the model can be specified as $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{c})=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})p(\mathbf{c})$ with $p(\mathbf{c})=Dir({\bm{{\bm{\epsilon}}}})$.
+Consequently, the variational distributions become $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{c},\mathbf{x}),q_{\bm{\phi}}(\mathbf{y},\mathbf{c}|\mathbf{x})$.
+This reformulates the learning objective as:
+$\mathcal{L}_{\text{clus}}(\mathbf{x})=$ $~{}\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z},\mathbf{c}))-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}\left[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))\right]$ (15)
+Figure 2: Unsupervised visualization on MNIST using three different priors (pinwheel, swiss roll, and square). Each color indicates a class.
+Method Pinwheel Swiss Roll Square ELBO MMD Latent NLL ELBO MMD Latent NLL ELBO MMD Latent NLL AEVB $-12.13\pm 0.41$ $0.77\pm 0.04$ $1.68\pm 0.31$ $-14.80\pm 0.23$ $0.78\pm 0.17$ $5.65\pm 1.58$ $-7.85\pm 0.29$ $1.10\pm 0.66$ $2.78\pm 0.61$ AEVB-IAF $-4.19\pm 0.05$ ${0.77\pm 0.00}$ $1.64\pm 0.73$ $-5.10\pm 0.30$ $0.61\pm 0.15$ $4.43\pm 1.09$ $-3.97\pm 0.22$ $0.75\pm 0.12$ $1.68\pm 0.27$ AAEB $N/A$ $0.68\pm 0.02$ $1.54\pm 0.19$ $N/A$ $\bm{0.52\pm 0.03}$ ${3.34\pm 0.16}$ $N/A$ $0.80\pm 0.02$ $2.46\pm 0.46$ H-AEVB $-7.03\pm 3.13$ ${0.74\pm 0.02}$ $2.25\pm 3.02$ $-7.21\pm 4.62$ ${0.70\pm 0.22}$ $4.04\pm 4.62$ $-5.71\pm 3.05$ ${0.76\pm 0.21}$ ${2.22\pm 2.03}$ PIS $-7.83\pm 0.64$ $0.75\pm 0.14$ $6.50\pm 1.11$ $-9.83\pm 0.61$ $0.61\pm 0.03$ $\bm{2.40\pm 1.01}$ $-7.06\pm 0.06$ $0.77\pm 0.04$ $3.67\pm 0.08$ DDVI $\bm{-3.88\pm 0.96}$ $\bm{0.67\pm 0.04}$ $\bm{1.27\pm 0.21}$ $\bm{-5.03\pm 0.58}$ $0.62\pm 0.33$ $3.86\pm 0.17$ $\bm{-3.79\pm 0.14}$ $\bm{0.66\pm 0.07}$ $\bm{1.56\pm 0.09}$
+Table 1: Unsupervised learning on MNIST. We report ELBO, MMD between generated images and test images, and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
+5 Experiments
+We compare DDVI with Auto-Encoding Variational Bayes (AEVB) [14], AEVB with inverse autoregressive flow posteriors (AEVB-IAF) [16], Adversarial Auto-Encoding Bayes (AAEB) [20], and Path Integral Sampler (PIS) [46] on MNIST [18] and CIFAR-10 [17] in unsupervised and semi-supervised learning settings, and also on the Thousand Genomes dataset [35].
+We also compare with Hierachical Auto-Encoding Variational Bayes (H-AEVB) [26, 38] in unsupervised setting.
+We discuss the computational costs of all methods in Appendix D.
+The priors, model architecture, and training details can be founded in Appendix F, Appendix G, and Appendix H respectively.
+All results below are reported with 95% confidence interval using 3 different seeds.
+5.1 Unsupervised learning
+We start with synthetic experiments that are aimed at benchmarking the expressivity of diffusion-based posteriors and their ability to improve fitting $p$, a distribution with a complex structured prior, like one might find in probabilistic programming, scientific analysis, or other applications.
+We fit a model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ on the MNIST and CIFAR-10 datasets with three priors $p(\mathbf{z})$: pinwheel, swiss roll, and square and report our results in Table 1 and Table 6.
+The model distribution $p_{\bm{\theta}}$ is instantiated by a deep Gaussian latent variable model (DGLVM) with multi-layer perceptrons (MLPs) on MNIST and convolutional neural networks (CNNs) on CIFAR-10.
+The details of model architecture are provided in Appendix G.
+Our first set of metrics (ELBO and MMD) seeks to evaluate the learned generative model $p_{\bm{\theta}}$ is good.
+In the ELBO calculation, we average the reconstruction loss across image pixels.
+We use MMD to measure sample quality: we generate images with the trained model and calculate MMD between the generated images and test images using a mixture of Gaussian kernel111with sigma equal to [2, 5, 10, 20, 40, 80].
+We only report MMD for MNIST, since CIFAR-10 generated samples are very low-quality for all methods because the latent dimension is 2.
+Our last metric seeks to directly evaluate the expressivity of the posterior.
+We measure latent negative log-likelihood (Latent NLL) by fitting a kernel density estimator (KDE) on the latents produced by the model with test data as input and compute the log-likelihood of the latents sampled from the prior under the fitted KDE.
+From Tables 1 and 6, we see our method DDVI achieve best ELBO in all but one scenario, in which it still performs competitively.
+We also see strong results in Latent NLL and Acc in many scenarios, except for Swiss Roll where AAEB does well.
+We present visualizations of MNIST using the baseline methods and our method in Figure 2.
+Figure 3: Semi-supervised visualization on MNIST with 1,000 labels using three different priors (pinwheel, swiss roll, and square). Each a indicates one class.
+Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-11.15\pm 0.53$ $0.93\pm 0.01$ $1.36\pm 0.03$ $-15.29\pm 1.33$ $0.68\pm 0.01$ $4.60\pm 0.23$ $-10.26\pm 0.25$ $0.86\pm 0.01$ $1.68\pm 0.02$ AEVB-IAF $-2.10\pm 0.26$ $\bm{0.95\pm 0.00}$ $\bm{1.06\pm 0.03}$ $-5.38\pm 1.78$ $0.90\pm 0.02$ $2.75\pm 0.14$ $-2.67\pm 0.83$ $\bm{0.91\pm 0.01}$ $\bm{0.90\pm 0.02}$ AAEB $N/A$ $0.89\pm 0.01$ $1.55\pm 0.01$ $N/A$ $0.88\pm 0.01$ $3.07\pm 0.05$ $N/A$ $1.94\pm 0.38$ $0.76\pm 0.13$ DDVI $\bm{-0.24\pm 0.13}$ $\bm{0.95\pm 0.00}$ $\bm{1.06\pm 0.01}$ $\bm{-2.89\pm 0.33}$ $\bm{0.92\pm 0.01}$ $\bm{2.09\pm 0.00}$ $\bm{0.02\pm 0.09}$ $0.90\pm 0.01$ $1.49\pm 0.03$
+Table 2: Semi-supervised learning on MNIST (1,000 labels).
+We report ELBO, accuracy using KNN (K=20) classifier (Acc), and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
+5.2 Semi-supervised Learning
+We also evaluate the performance of our method and the baselines under semi-supervised learning setting where some labels are observed (1,000 for MNIST and 10,000 for CIFAR-10) and the partitions of the priors are known.
+For this setting, we evaluate ELBO, latent negative log-likelihood (Latent NLL), and a k-nearest neighbors classification accuracy of the latents (Acc).
+We choose classification accuracy since classification is a common downstream task for semi-supervised learning.
+We use the same set of priors and baselines.
+Details on how we partition each prior into $p_{\bm{{\bm{\theta}}}}(\mathbf{z}|\mathbf{x},l)$ can be founded in Appendix F.
+The partitions defined for our priors are local parts of the priors.
+We note that unlike unsupervised learning, we use the simplified sleep term from Appendix B in our objective for this setting, since $q_{\bm{\phi}}$ already gets extra information from $l$ here.
+The results are shown in Tables 2 and 7.
+DDVI mostly outperforms the baselines across different priors and metrics, especially on CIFAR-10 where DDVI is best across the board.
+For MNIST, DDVI always achieves the best ELBO, and it also performs competitively with other baselines in classification accuracy.
+We also show the visualizations of the latents in Figure 3 where DDVI matches the prior almost perfectly.
+5.3 Clustering and Visualization for Genotype Analysis
+Figure 4: Visualization of genotype clusters. A color represents one ethnicity.
+In this section, we report results on an real-world task in genome analysis.
+Visualizing genotype data reveals patterns in the latent ancestry of individuals.
+We instantiate DDVI with a deep Gaussian latent variable model (DGLVM) and compare it against with the three strong clustering baselines using the 1000 Genomes dataset.
+We also report visualizations from three dimensionality reduction algorithms: PCA, TSNE, and UMAP.
+For each clustering algorithm, we seek to discover up to 20 clusters.
+We report quantitative results in terms of cluster purity, cluster completeness, and normalized mutual information (NMI).
+There is an inherent trade-off between cluster purity completeness.
+The overall clustering performance can be captured with NMI.
+Method Cluster Cluster NMI
+Method Purity Completeness
+AEVB $0.28\pm 0.02$ $\bm{0.78\pm 0.16}$ $0.59\pm 0.08$
+AEVB-IAF $0.29\pm 0.04$ $\bm{0.73\pm 0.06}$ $0.55\pm 0.06$
+AAEB $0.37\pm 0.06$ $\bm{0.76\pm 0.11}$ ${0.63\pm 0.02}$
+DDVI $\bm{0.45\pm 0.03}$ $\bm{0.75\pm 0.05}$ $\bm{0.66\pm 0.04}$
+Figure 5: Quantitative genotype clustering results.
+In Figure 5, we see that DDVI attains the best performance on cluster purity and NMI.
+For cluster completeness, VAE and AAE have better means but much larger confidence interval.
+Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA [44], t-SNE [40], and UMAP [23]).
+The legend of Figure 4 can be founded at Figure 6.
+6 Discussion
+Diffusion vs. Normalizing Flows
+Our approach is most similar to flow-based approximators [28, 16]; in fact when $T\to\infty$, our diffusion-based posterior effectively becomes a continuous-time normalizing flow [36].
+However, classical flow-based methods require invertible architectures for each flow layer: this constrains their expressivity and requires backpropagating through potentially a very deep network.
+Our approach, on the other hand, trains a model (a continuous-time flow when $T\to\infty$) via a denoising objective (similar to score matching) that does not require invertible architectures and effectively admits an infinite number of layers (with weight sharing).
+This model is trained not by backpropagating through the ELBO, but rather via an auxiliary diffusion loss term (effectively, a score matching objective).
+Despite training with a modified loss, we observe in Section 5 that a diffusion model with an expressive denoising architecture yields an improved ELBO relative to regular flows.
+Also, our modified loss based on the forward KL divergence reduces posterior collapse (i.e., all modes of the prior are covered well), and thus produces better samples.
+Diffusion vs. Other Generative Models
+Variational posteriors based on GANs [20] also admit expressive architectures and require only sample-based access to the prior $p(\mathbf{z})$.
+Our diffusion-based approach admits a more stable loss, and is potentially more expressive, as it effectively supports an infnite number of layers (with shared parameters when $T\to\infty$).
+Unlike GANs, our models also admit explicit likelihoods and allow us to compute the ELBO for model evaluation.
+Our approach is similar to variational MCMC [31]; however, we train with a better objective augmented with a diffusion loss, and we adopt improved architectures with shared weights across layers.
+Diffusion for Approximate Inference
+[…]
+Our algorithms are also simpler.
+For example, diffusion sampling methods require backpropagating through a sampling process to minimize the reverse $\text{KL}(q_{\bm{\phi}}||p_{\bm{\theta}})$, which poses challenges with optimization and credit assignment.
+Some methods based on Schrodinger bridges require an iterative optimization process generalizing the sinkhorn algorithm or computationally expensive on-policy or off-policy [21] trajectory-based optimization.
+In contrast, DDVI optimizes the forward $\text{KL}(p_{\bm{\theta}}||q_{\bm{\phi}})$ using simple gradient-based optimization that directly emulates diffusion-based training.
+7 Related Work
+Latent Diffusion
+Vahdat et al.
+[39], Wehenkel and Louppe [43], Rombach et al.
+[30] perform diffusion in the latent space of a VAE to improve the efficiency of image generation.
+Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors.
+Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
+Recent work [25, 47, 42] has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$.
+Cohen et al.
+[5] crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
+Diffusion for Approximate Inference
+Diffusion sampling [3, 46, 41, 45, 29, 32, 1] mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function.
+These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) [46], denoising diffusion sampler (DDS) [41], and time-reversed diffusion sampler (DIS) [3], which have been unified by Richter et al.
+[29].
+Some other works [45, 1] use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration.
+Sendera et al.
+[32] benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps.
+Dimensionality Reduction
+Latent variable models in general are an attractive alternative to visualization methods like PCA, UMAP, and t-SNE [23, 40].
+Domain-specific knowledge can be injected through the prior, and deep neural networks can be utilized to achieve a more expressive mapping from the data space to the latent space.
+Nevertheless, downsides of LVMs are that they are more computationally expensive and require careful hyperparameter tuning.
+8 Conclusion
+While this paper focuses on applications of DDVI to dimensionality reduction and visualization, there exist other tasks for the algorithm, e.g., density estimation or sample quality.
+Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making [24, 6], meta-learning [27], or causal effect estimation [7].
+Since our learning objective differs from the ELBO (it adds a regularizer), we anticipate gains on models whose training benefits from regularization, but perhaps not on all models.
+Also, attaining competitive likelihood estimation requires architecture improvements that are orthogonal to this paper.
+However, our ability to generate diverse samples and achieve class separation in latent space hints at the method’s potential on these tasks.
+Appendix A Pseudocode
+Here we provide a pseudocode to illustrate the training process of DDVI. We release anonymized code repo at https://anonymous.4open.science/r/vi-with-diffusion-D70F.
+Algorithm 1 DDVI Pseudocode
+1:(Optional) Pre-train $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})$ and $q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})$ with DDVI but with unconditional diffusion model $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$
+2:for epoch $=1,\ldots,n$ do
+3:     for $\mathbf{x}_{1},\ldots,\mathbf{x}_{k}\sim p_{\mathcal{D}}(\mathbf{x})$ do
+4:         $\mathbf{y}_{i}\sim q_{\bm{\phi}}(\mathbf{y}|\mathbf{x}_{i})$ and $\mathbf{z}_{i}\sim q(\mathbf{z}|\mathbf{y}_{i},\mathbf{x}_{i})$ for $i=1,\ldots,k$
+5:         Optimize ${\bm{\theta}},{\bm{\phi}}$ with respect to a Monte Carlo estimate of $\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\left(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\|p_{\bm{\theta}}(\mathbf{y},\mathbf{z})\right)$ for each $\mathbf{x}_{i}$triangleright Standard ELBO training part
+6:         for iteration $=1,\ldots,m$ do $\triangleright$ Do sleep for $m$ iterations
+7:              $\mathbf{z}_{1},\ldots,\mathbf{z}_{k}\sim p(\mathbf{z})$triangleright Batch-sample latents from prior
+8:              $\hat{\mathbf{x}}_{i}\sim p(\mathbf{x}|\mathbf{z}_{i})$ for $i=1,\ldots,k$triangleright Construct fantasy inputs
+9:              $\mathbf{y}_{i}\sim r(\mathbf{y}|\mathbf{z}_{i})$ for $i=1,\ldots,k$triangleright Construct fantasy inputs
+10:              Optimize ${\bm{\phi}}$ using the standard diffusion noise prediction loss on $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{i},\hat{\mathbf{x}}_{i})$
+11:         end for
+12:     end for
+13:end for
+Appendix B Simplifying Wake-Sleep
+In wake-sleep, sampling $\mathbf{x}$ from $p_{\bm{\theta}}$ to obtain gradients for the sleep term introduces computational overhead.
 To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation ^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ of $\mathcal{L}$:
 ^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ $=\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / reconstr.
-term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}-\underbrace{D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))}_{\text{latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$}}.$ (7)
+term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))}_{\text{latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$}}.$ (16)
 We have replaced $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ with a latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$, in which $\mathbf{x}$ is given, and we only seek to fit the true reverse noising process $r(\mathbf{z}|\mathbf{y})$ independently of $\mathbf{x}$.
@@
 We may think of $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$ as a regularizer to the ELBO.
-3.1.3 DDVI with Diffusion-Based Encoders
-Our framework is naturally instantiated with diffusion models.
-Let $\mathbf{y}_{0}=\mathbf{z}$ and $\mathbf{y}_{T}=\mathbf{y}$.
-The forward noising model can be defined as $r(\mathbf{y}_{1:T}|\mathbf{z})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where the $\mathbf{y}_{0:T}$ are an extended set of latent variables that represent increasingly noised versions of $\mathbf{y}_{0}$.
-We parameterize the approximate reverse diffusion process as $q_{\bm{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}).$
-[…]
-$\mathcal{L}_{\text{diff}}=$ $\mathbb{E}_{r}\left[\log q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}D_{\mathrm{KL}}(r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0})||q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}))\right]-D_{\mathrm{KL}}(r(\mathbf{y}_{T}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})).$ (10)
-This bound is an instantiation of Equation 8 when $r$ is a diffusion process.
-Many choices of $r,q$ fit this framework.
-A common type of noising process is Gaussian diffusion, where we define $r(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-{\bm{\alpha}}_{t}}\mathbf{y}_{t-1},{\bm{\alpha}}_{t}\mathbf{I})$ for a suitable schedule $({\bm{\alpha}}_{t})_{t=1}^{T}$.
-We then adopt the parameterization $q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})=\mathcal{N}(\mathbf{y}_{t-1};\mu_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t),\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t))$.
-It is then common to parameterize $q_{\bm{\phi}}$ with a noise prediction network ${\bm{\epsilon}}_{\bm{\phi}}$ (Ho et al., 2020); the sum of KL divergences can be approximated by $E_{t,{\bm{\epsilon}}_{t}\sim r(\mathbf{y}_{0},t)}||{\bm{\epsilon}}_{t}-{\bm{\epsilon}}_{\bm{\phi}}(\sqrt{\bar{{\bm{\alpha}}_{t}}}\mathbf{y}_{0}+\sqrt{1-\bar{{\bm{\alpha}}_{t}}}{\bm{\epsilon}}_{t},\mathbf{x},t)||^{2}$.
-By our earlier argument, this objective encourages $q_{\bm{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})$ to match $r(\mathbf{y}_{0:T})$ as well as the true posterior $p(\mathbf{z}|\mathbf{x})$.
-Lastly, we need to show that this choice of $q$ admits a tractable entropy. This follows from
-$H(q)$ $=-\sum_{t=1}^{T+1}\mathbb{E}_{q}[\log q(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]=\sum_{t=1}^{T+1}\mathbb{E}_{q}[\frac{d}{2}\left(1+\log(2\pi)\right)+\frac{1}{2}\log|\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x})|]$ (11)
-where $d$ is the dimension of $\mathbf{y}$ and we use the notation $\mathbf{y}_{T+1}=\mathbf{x}$.
-The right-hand term can be approximated using Monte Carlo; it is also common to leave the variance $\Sigma_{\bm{\phi}}$ fixed (as we typically do in our experiments), in which case $H(q)$ is a constant.
-3.2 Auxiliary-Variable Diffusion Encoders
-Denoising diffusion variational inference can also be understood from the perspective of auxiliary-variable generative models (Maaløe et al., 2016).
-Our method can be seen as introducing an approximate posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})=\int_{\mathbf{y}}q_{\bm{\phi}}(\mathbf{z},\mathbf{y}|\mathbf{x})$ which itself contains latent variables $\mathbf{y}$.
-The standard approach to fit auxiliary-variable generative models is to apply the ELBO twice:
-$\log p_{\bm{\theta}}(\mathbf{x})$ $\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))=\text{ELBO}_{\mathbf{z}|\mathbf{x}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ (12)
-$\geq\text{ELBO}_{\mathbf{z}|\mathbf{x}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})-\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$ (13)
-$=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x},\mathbf{z})r(\mathbf{y}|\mathbf{x},\mathbf{z})-\log q(\mathbf{y},\mathbf{z}|\mathbf{x})],$ (14)
-where in Equation 13 we applied the ELBO over $\mathbf{z}$, and in Equation 12 we applied the ELBO again over the latent $\mathbf{y}$ of $q$.
-The $r(\mathbf{y}|\mathbf{x},\mathbf{z})$ can be interepreted as an approximate variational posterior for the true posterior $q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})$ of the data distribution.
-Specifically, the gap between Equation 12 and $\log p_{\bm{\theta}}(\mathbf{x})$ is precisely $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))+\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$.
-Thus, if we correctly match the pairs of approximate posteriors, we will achieve a tight bound.
-Additionally, notice that Equation 14 is exactly the ELBO in Equation 2 after rearranging terms.
-Thus the two perspectives are equivalent: using variational inference to fit an augmented model $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ is equivalent to fitting the original $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ and applying the ELBO to $p_{\bm{\theta}}$, followed by applying the ELBO again to the latent $\mathbf{y}$ in $q$.
-Since our objective $\mathcal{L}$ is Equation 14) augmented with $\mathcal{L}_{\text{sleep}}$, denoising VI can also be interpreted as fitting an latent-variable approximate posterior $q$.
-As seen in Equation 9, we implicitly minimize the reverse KL divergence $\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})}[D_{\mathrm{KL}}(r(\mathbf{y}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{z},\mathbf{x}))]$.
-Note, however, that there is a crucial difference between our method and auxiliary-variable generative models: we do not optimize the parameters of $r$.
-Rather, $r$ is a user-defined noising process that acts as a regularizer for the model $q_{\bm{\phi}}$ and whose reverse we seek to fit.
-3.3 Regularizing Variational Inference with Denoising Processes
-Lastly, our approach can be understood from a third perspective: that of regularization.
-As we mentioned earlier, $\mathcal{L}$ is the ELBO with an additional regularization term that encourages the approximate posterior $q(\mathbf{z}|\mathbf{x},\mathbf{y})q(\mathbf{y}|\mathbf{x})$ to gradually transform $\mathbf{y}$ into $\mathbf{z}$ following the guidance of $r$.
-In that sense, $r$ provides data augmentation to ensure that the mapping from $\mathbf{y}$ to $\mathbf{x}$ is easier to learn, similarly to how a diffusion process provides additional learning signal for generating $\mathbf{x}$ from noise.
-Our framework also supports the user defining general distributions as regularizers $r_{1}(\mathbf{z},\mathbf{y})$ and $r_{2}(\mathbf{y}|\mathbf{z},\mathbf{x})$.
-Here, $r_{1}$ puts a custom prior on $\mathbf{z},\mathbf{y}$, while $r_{2}$ specifies a mapping between $\mathbf{y},\mathbf{z}$ that should be followed by the encoder model.
-We write $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}},r_{1})$ and $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}},r_{2})$, where each term is still defined as in Equation 3 and Equation 6, except using $r_{1},r_{2}$ and we optimize
-$\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}},r_{1},r_{2})=\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\beta\cdot\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}},r_{1})+\gamma\cdot\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}},r_{2}),$ (15)
-where $\beta,\gamma>0$ (in particular, $\beta$ mirrors the $\beta$-VAE model (Higgins et al., 2016)).
-Examples of $r_{2}$ include Gaussian diffusion, discrete noising processes, as well as noise conditioned on $\mathbf{x}$.
-Note that $r_{2}$ could even be fully deterministic when parameterized with a bijective flow.
-With a strong regularizer, we could set $q(\mathbf{z}|\mathbf{y},\mathbf{x})$ to equal the reverse $r_{2}(\mathbf{z}|\mathbf{y},\mathbf{x})$ of this flow.
-Thus, for suitably designed $r_{2}$, we could perfectly match the shape of any mapping.
-This mirrors the InteL-VAE framework (Miao et al., 2021), with support for flexible and probabilistic regularizers.
-Most importantly, our method can be understood as introducing a new form of regularization.
-While a standard VAE introduces a prior regularization term, we introduce a new form of regularization via a function $r_{2}$.
-This can be advantageous when specifying or enforcing certain priors is difficult.
-4 Extensions
-4.1 Semi-Supervised Learning
-Following Makhzani et al.
-(2015), we extend our algorithm to the semi-supervised learning setting where some data points are labeled and we define $p(\mathbf{z},l)=p(\mathbf{z}|c)p(l)$.
-Then, the model can be specified as $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\bm{\theta}}(\mathbf{z}|l)p(l)$ and the variational distributions are $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\bm{\phi}}(\mathbf{y}|\mathbf{x}),q_{\bm{\phi}}(l|\mathbf{x})$.
-In this setting, we consider two cases of whether the label can be observed or not (Kingma et al., 2014).
-We extend Equation 7 to incorporate the label $l$ corresponding to a data point as follows:
-$\mathcal{L}_{\text{semi}}=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}|l))-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|l)||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},l))$
-When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\bm{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\bm{{\bm{\theta}}}},{\bm{\phi}})+D_{\mathrm{KL}}(q_{\bm{\phi}}(l|\mathbf{x})||p(l))$.
-Therefore, we can conclude a marginal likelihood on our dataset as follows: ~$\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\bm{{\bm{\theta}}}},{\bm{{\bm{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\bm{{\bm{\theta}}}},{\bm{{\bm{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
-We also want to guarantee that all model parameters can be learned in all cases, including $q_{\bm{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference.
-Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: ~$\tilde{\mathcal{L}}_{\text{semi}_{\bm{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\bm{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\bm{\phi}}(l|\mathbf{x})\right]$
-4.2 Clustering
-We have further extended our algorithm to encompass the clustering paradigm.
-We propose two distinct strategies.
-In the first approach, we simply set $p_{\bm{{\bm{\theta}}}}(\mathbf{z})$ as a mixture of desired priors.
-The means of these priors are characterized by ${\bm{{\bm{\theta}}}}$.
-From these means, cluster membership, denoted as $\mathbf{c}$ can be deduced.
-Remarkably, this approach requires no alteration to the existing learning objective.
-Alternatively, the second method retains the original prior but introduces an additional cluster latent variable $\mathbf{c}$ where $\sum_{i}c_{i}=1$.
-Thus, the model can be specified as $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{c})=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})p(\mathbf{c})$ with $p(\mathbf{c})=Dir({\bm{{\bm{\epsilon}}}})$.
-Consequently, the variational distributions become $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{c},\mathbf{x}),q_{\bm{\phi}}(\mathbf{y},\mathbf{c}|\mathbf{x})$.
-This reformulates the learning objective as:
-$\mathcal{L}_{\text{clus}}(\mathbf{x})$ $=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z},\mathbf{c}))-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))$
-5 Experiments
-We compare DD-VAE with vanilla VAE (Kingma & Welling, 2013), IAF-VAE (Kingma et al., 2016), and AAE (Makhzani et al., 2015) on MNIST (Lecun et al., 1998) and CIFAR-10 (Krizhevsky & Hinton, 2009) in unsupervised and semi-supervised learning settings, and also on the 1000 genomes dataset Siva (2008) in clustering settings.
-The differences between these methods are summarized in Table 8.
-We discuss the computational costs of all methods in Section A.1.
-The priors, model architecture, and training details can be founded in Section A.2, Section A.3, and Section A.4 respectively.
-All results below are reported with 95% confidence interval using 5 different seeds.
-5.1 Unsupervised learning
-We fit a model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ on the MNIST and CIFAR-10 datasets with three priors $p(\mathbf{z})$: pinwheel, swiss roll, and square and report our results in Table 1 and Table 5.
-We measure a k-nearest neighbors classification accuracy of the latents (Acc).
-We also measure latent negative log-likelihood (Latent NLL) by fitting a kernel density estimator (KDE) on the latents produced by the model with test data as input and compute the log-likelihood of the latents sampled from the prior under the fitted KDE.
-We consistently see our method perform best on both KNN accuracy and latent NLL, indicating that its output latents are the most informative of labels and best-aligned with the priors.
-Table 1 reports qualiatitve results that indicate that DD-VAE enforces most accurately the correct prior shape.
-Figure 2: Unsupervised visualization on MNIST using three different priors (pinwhell, swiss roll, and square). Each color indicates a class.
-Method Pinwheel Swiss Roll Square Acc Latent NLL Acc Latent NLL Acc Latent NLL VAE $0.58\pm 0.01$ $6.18\pm 0.93$ $0.56\pm 0.055$ $82.96\pm 99.30$ $0.35\pm 0.02$ $25.07\pm 62.41$ IAF-VAE $\bm{0.71\pm 0.00}$ $2.06\pm 0.21$ $0.67\pm 0.02$ $12.13\pm 18.28$ $0.65\pm 0.03$ $1.53\pm 0.82$ AAE $0.67\pm 0.02$ $1.94\pm 0.09$ $0.68\pm 0.03$ $3.43\pm 0.14$ $0.57\pm 0.02$ $1.97\pm 0.70$ HVAE w/ flow $\bm{0.72\pm 0.01}$ $2.18\pm 0.13$ $\bm{0.71\pm 0.02}$ $4.17\pm 1.23$ $\bm{0.70\pm 0.01}$ $\bm{1.25\pm 0.08}$ DiffVAE $\bm{0.72\pm 0.01}$ $\bm{1.57\pm 0.10}$ $0.66\pm 0.01$ $\bm{3.19\pm 0.07}$ $\bm{0.72\pm 0.03}$ $\bm{1.24\pm 0.13}$
-Table 1: Unsupervised learning on MNIST. We report accuracy using KNN (K=20) classifier and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
-5.2 Semi-supervised Learning
-We also evaluate the performance of our method and the baselines under semi-supervised learning setting where some labels are observed (1,000 for MNIST and 10,000 for CIFAR-10) and the partitions of the priors are known.
-We use the same set of priors and baselines.
-Details on how we partition each prior into $p_{\bm{{\bm{\theta}}}}(\mathbf{z}|l)$ can be founded in Section A.2.
-The partitions defined for our priors are local parts of the priors.
-We substitute the KNN accuracy metric with the average class latent NLL, which measures how well the test latents from each class match their respective partition of the prior $p_{\bm{{\bm{\theta}}}}(\mathbf{z}|l)$.
-In our case, this metric implicitly measures accuracy because the metric requires that the test latents for each class be close to each other, as the partitions defined are local parts of the full prior.
-The results are shown in Table 2 and Table 6.
-DD-VAE mostly outperforms the baselines across different priors and metrics.
-We also show the visualization in Figure 3 where DD-VAE matches the prior almost perfectly.
-Figure 3: Semi-supervised visualization on MNIST with 1,000 labels using three different priors (pinwheel, swiss roll, and square). Each a indicates one class.
-Method Pinwheel Swiss Roll Square Latent NLL Avg Class Latent NLL Avg Class Latent NLL Avg Class Latent NLL Latent NLL Latent NLL VAE $7.77\pm 1.83$ $15.64\pm 8.79$ $56.91\pm 33.90$ $244.51\pm 94.19$ $1.30\pm 0.03$ $-0.71\pm 0.05$ IAF-VAE $2.71\pm 0.22$ $1.84\pm 0.46$ $8.10\pm 12.17$ $14.59\pm 31.88$ $1.04\pm 0.04$ $-1.01\pm 0.03$ AAE $1.90\pm 0.04$ $0.64\pm 0.11$ $\bm{3.35\pm 0.17}$ $1.86\pm 0.77$ $1.70\pm 0.26$ $2.50\pm 4.70$ DiffVAE $\bm{1.28\pm 0.01}$ $\bm{-0.94\pm 0.01}$ $\bm{3.35\pm 0.10}$ $\bm{1.24\pm 0.09}$ $\bm{0.96\pm 0.03}$ $\bm{-1.20\pm 0.02}$
-Table 2: Semi-supervised learning on MNIST (1,000 labels). We report average class latent negative log-likelihood and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
-5.3 Clustering and Visualization for Genotype Analysis
-Figure 4: Visualization of genotype clusters. A color represents one ethnicity.
-In this section, we report results on an real-world task in genome analysis.
-Visualizing genotype data reveals patterns in the latent ancestry of individuals.
-We compare DD-VAE against with the three strong clustering baselines using the 1000 Genomes dataset.
-We also report visualizations from three dimensionality reduction algorithms: PCA, TSNE, and UMAP.
-For each clustering algorithm, we seek to discover up to 20 clusters.
-We report quantitative results in terms of cluster purity, cluster completeness, and normalized mutual information (NMI).
-There is an inherent trade-off between cluster purity completeness.
-The overall clustering performance can be captured with NMI.
-Method Cluster Cluster NMI
-Method Purity Completeness
-VAE $0.28\pm 0.02$ $\bm{0.78\pm 0.16}$ $0.59\pm 0.08$
-IAF-VAE $0.29\pm 0.04$ $\bm{0.73\pm 0.06}$ $0.55\pm 0.06$
-AAE $0.37\pm 0.06$ $\bm{0.76\pm 0.11}$ $\bm{0.63\pm 0.02}$
-DiffVAE $\bm{0.45\pm 0.03}$ $\bm{0.75\pm 0.05}$ $\bm{0.66\pm 0.04}$
-Table 3: Quantitative genotype clustering results.
-In Table 3, we see that DD-VAE attains the best performance on cluster purity and NMI.
-For cluster completeness, VAE and AAE have better means but much larger confidence interval.
-Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA (Wold et al., 1987), t-SNE (Van der Maaten & Hinton, 2008), and UMAP (McInnes et al., 2018)).
-The legend of Figure 4 can be founded at Figure 5.
-6 Related Work, Discussion, and Conclusion
-There are a number of recent works merging VAE and diffusion models.
+Appendix C Comparision of Methods
+We provide a comprehensive comparison of different methods in Table 3.
 Vahdat et al.
-(2021); Wehenkel & Louppe (2021) use diffusion with VAE priors.
-Recent research (Preechakul et al., 2022; Zhang et al., 2022; Wang et al., 2023b) has also melded auto-encoders with diffusion models, focusing on semantically meaningful latents.
+[39], Wehenkel and Louppe [43], Rombach et al.
+[30] perform diffusion in the latent space of a VAE to improve the efficiency of image generation.
+Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors.
+Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
+Recent work [25, 47, 42] has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$.
 Cohen et al.
-(2022) crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
-These works focus on sample quality and directly fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to a complex distribution (e.g., compressed images (Rombach et al., 2022)); our work instead seeks to achieve high sample quality in latent space, and obtains complex $\mathbf{z}$ by diffusion from a simple $\mathbf{y}\sim q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})$.
-Latent variable models in general are an attractive alternative to visualization methods like PCA, UMAP, and t-SNE (McInnes et al., 2018; Van der Maaten & Hinton, 2008).
-Domain-specific knowledge can be injected through the prior, and deep neural networks can be utilized to achieve a more expressive mapping from the data space to the latent space.
-Nevertheless, downsides of LVMs are that they are more computationally expensive and require careful hyperparameter tuning.
-While this paper focuses on applications of DD-VAE to dimensionality reduction and visualization, there exist other tasks for the algorithm, e.g., density estimation or sample quality.
-Since our learning objective differs from the ELBO (it adds a regularizer), we anticipate gains on models whose training benefits from regularization, but perhaps not on all models.
-Also, attaining competitive likelihood estimation requires architecture improvements that are orthogonal to this paper.
-However, our ability to generate diverse samples and achieve class separation in latent space hints at the method’s potential on these tasks.
-Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making or causal effect estimation (Nguyen & Grover, 2022; Deshpande et al., 2022; Deshpande & Kuleshov, 2023; Rastogi et al., 2023).
-Appendix A Appendix
-A.1 Computational Cost Analysis
-Method NMI values at different wall-clock training times NMI @ 10 min NMI @ 20 min NMI @ 30 min NMI @ 40 min NMI @ 50 min NMI @ 60 min VAE $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ IAF-VAE $0.54$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ AAE $0.61$ $0.57$ $0.57$ $0.57$ $0.57$ $0.57$ DiffVAE (T=5) warm up $0.63$ $0.63$ $0.66$ $0.66$ $0.66$ DiffVAE (T=10) warm up $\mathbf{0.64}$ $\mathbf{0.68}$ $\mathbf{0.70}$ $\mathbf{0.70}$ $\mathbf{0.70}$ DiffVAE (T=20) warm up $0.50$ $0.51$ $0.56$ $0.64$ $0.68$ DiffVAE (T=50) warm up $0.52$ $0.54$ $0.51$ $0.59$ $0.59$
+[5] crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
+Model Training Objective Approximating Family Sample-based Prior Auxiliary Variable Tasks Simplified Graphical Illustration AEVB ELBO Diagonal Gaussian Density estimation $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ AEVB-IAF ELBO Normalizing flow Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{x}$ AAEB Adversarial training Adversarial generator Visualization $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ H-AEVB-(IAF) ELBO Factorial Normal / Normalizing flow Density estimation / High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ ADGM ELBO Non-Gaussian Density estimation $\mathbf{x}\rightarrow\mathbf{a}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ LDM ELBO Diagonal Gaussian High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ LSGM ELBO & score matching Diagonal Gaussian High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ DDVI ELBO & sleep term Denoising diffusion Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{T}(\mathbf{y})\rightarrow\mathbf{z}_{0}(\mathbf{z})\rightarrow\mathbf{x}$
+Table 3: Comparison of DDVI to other relevant methods.
+$\mathbf{x}$ represents the original data input to the model.
+$\mathbf{z}$ denotes the latent (hidden) representation of the input data.
+$\mathbf{a}$ represents an auxiliary variable introduced in some models (like ADGM) to capture additional aspects of the data distribution or to assist in the model’s learning process.
+Appendix D Computational Cost Analysis
+Method NMI values at different wall-clock training times NMI @ 10 min NMI @ 20 min NMI @ 30 min NMI @ 40 min NMI @ 50 min NMI @ 60 min AEVB $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ AEVB-IAF $0.54$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ AAEB $0.61$ $0.57$ $0.57$ $0.57$ $0.57$ $0.57$ DDVI (T=5) warm up $0.63$ $0.63$ $0.66$ $0.66$ $0.66$ DDVI (T=10) warm up $\mathbf{0.64}$ $\mathbf{0.68}$ $\mathbf{0.70}$ $\mathbf{0.70}$ $\mathbf{0.70}$ DDVI (T=20) warm up $0.50$ $0.51$ $0.56$ $0.64$ $0.68$ DDVI (T=50) warm up $0.52$ $0.54$ $0.51$ $0.59$ $0.59$
 Table 4: Computational cost trade-off on 1kgenome: NMI vs wall-clock training time
-We conduct a computational cost analysis between the baselines and DD-VAE with various timesteps on the genotype clustering/visualization experiments.
-The table Table 4 below shows that DD-VAE outperforms baselines at all timestamps and continues to improve after the baselines have plateaued.
-A.2 Priors
+We conduct a computational cost analysis between the baselines and DDVI with various timesteps on the genotype clustering/visualization experiments.
+Table 4 shows that DDVI outperforms baselines at all timestamps and continues to improve after the baselines have plateaued.
+Appendix E Connections to Diffusion Samplers
+Diffusion sampling [3, 46, 41, 45, 29, 32, 1] mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function.
+These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) [46], denoising diffusion sampler (DDS) [41], and time-reversed diffusion sampler (DIS) [3], which have been unified by Richter et al.
+[29].
+Some other works [45, 1] use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration.
+Sendera et al.
+[32] benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps.
+Overall, there are indeed some strong connections between these works and ours:
+  • They also focus on variational methods that directly fit a parametric family of tractable distributions (given by controlled SDEs) to the target density.
+  • They cast the density estimation/sampling problem into an optimization problem over a control objective, which learns control drifts (and diffusion) parameterized by neural networks.
+But we would like to clarify that there are also some clear differences between them:
+  • The diffusion-structured samplers only focus on density estimation/sampling but ignore the problem of learning a generative model, which is one of the main focuses of our work.
+We aim to perform more accurate variational inference using an auxiliary variable model augmented by diffusion models to improve generative modeling.
+In our setting, $p_{\bm{\theta}}(z|x)$ is a moving target density, as we jointly learn ${\bm{\theta}}$ with ${\bm{\phi}}$, as opposed to a static target density that diffusion-structured samplers are designed to solve.
+  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy [46] or off-policy [21] trajectory-based optimization.
+Instead, we introduce a wake-sleep optimization algorithm and its simplified version to alleviate the weak learning signal issue and optimize the evidence lower bound in a better way.
+  • In Equation 9, we are minimizing the forward KL divergence $\mathcal{D}_{KL}(p_{\bm{\theta}}||q_{\bm{\phi}})$, where diffusion samplers are minimizing the reverse $\mathcal{D}_{KL}(q_{\bm{\phi}}||p_{\bm{\theta}})$.
+We also summarize the connections and differences in the table below.
+Diffusion Samplers GFlowNet-based DDVI (ours)
+[3, 46, 41, 1] [45, 29, 32, 21]
+Tasks Sampling, density estimation Sampling, density estimation Learning, Sampling, Dimensionality reduction
+Model Family for $p$ Any energy-based Any energy-based Latent with tractable $p(x|z),p(z)$
+Model Family for $q$ Markov chain Markov chain Markov chain
+Objective $D_{KL}(q||p)$ with regularizer Trajectory balance objective $D_{KL}(p||q)$ with ELBO
+Algorithm Gradient descent (with reference process), importance sampling RL-motivated off-policy optimization (replay buffers, Thompson sampling, etc.) Gradient descent with wake-sleep
+Compatible Models Anything energy-based Anything energy-based LDA, deep latent-variable models
+Applications Sampling from physics-based models, model selection based on NLL Sampling from physics-based models, model selection based on NLL Probabilistic programming, visualization
+Table 5: Comparison of Diffusion-structured Samplers, GFlowNet-based Approaches, and DDVI
+Appendix F Priors
 Below we describe the sampling process for each prior.
 Pinwheel.
-This distribution was used in (Johnson et al., 2016).
+This distribution was used in [13].
 We define the number of clusters to be 10.
@@
 Swiss Roll.
-This distribution was used in Marsland (2014).
-We add noise $\sigma=0.1$ to the prior.
+This distribution was used in [22].
 For semi-supervised learning experiments, this prior is partitioned into 10 partitions.
-The samples from the prior, before we add noise to it, can actually be characterized by a single scalar representing how far you are long the swiss roll from the center.
+The samples from the prior can actually be characterized by a single scalar representing how far you are long the swiss roll from the center.
 The paritioning is done by creating 10 equal-length intervals in this 1D space.
@@
 The partitioning is done by creating 10 equal-length intervals in the 1D position space.
-VAE and IAF-VAE requires that we can evaluate the prior density.
+AEVB and AEVB-IAF requires that we can evaluate the prior density.
 To do this, for all priors, we evaluate the density by fitting a kernel density estimator with mixture of gaussian kernel with bandwidth equal to 0.005, 0.008, 0.01, 0.03, and 0.05.
-A.3 Model Architecture
-All methods use the same architecture for encoder $q_{\bm{\phi}}(z|x)$ and decoder $p_{\bm{\theta}}(x|z)$, excluding the extra parts specific to each method which we describe below, for the same dataset.
+Appendix G Model Architecture
+All methods use the same architecture for encoder $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ and decoder $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})$, excluding the extra parts specific to each method which we describe below, for the same dataset.
 For MNIST, the encoder and decoder are multi-layer perceptron with two hidden layers, each with 1000 hidden units.
 For CIFAR-10, the encoder is a 4-layer convolutional neural network with (16, 32, 64, 128) channels with a linear layer on top, and the decoder is a 4-layer tranposed convolutional neural network with (64, 32, 16, 3) channels where a linear layer is used to first turn the feature dimension from 2 to 64.
-IAF-VAE employs 4 IAF transformations on top of the encoder, each is implemented with a 4-layer MADE.
-The number of hidden units in MADE is 24.
+AEVB-IAF employs 4 IAF transformations on top of the encoder, each is implemented with a 4-layer MADE.
+The number of hidden units in MADE is 128.
 The ordering is reversed between every other IAF transformation.
-AAE has a discriminator, used in adversarial training, which is a multi-layer perceptron with two hidden layers, each with 1000 hidden units.
-DD-VAE has a diffusion model on top of the encoder.
+AAEB has a discriminator, used in adversarial training, which is a multi-layer perceptron with two hidden layers, each with 1000 hidden units.
+DDVI has a diffusion model on top of the encoder.
 The time-conditioned reverse diffusion distribution is implemented with a 5-layer time-conditioned multi-layer perceptron, each with 128 hidden units.
 A time-conditioned linear layer learns an additional embedding for each timestep and adds it to the output of the linear layer.
-A.4 Training Details
-For training, we use Adam optimizer and latent size of 2 for all of our experiments. The training details of each algorithm are detailed below:
-VAE.
+Appendix H Training Details
+For training, we update the parameters for each batch of inputs by alternating between the ELBO phase (optimizing ${\bm{\theta}}$ and ${\bm{\phi}}$ with respect to the ELBO, i.e., the reconstruction term and the prior matching term) and the sleep phase (optimizing ${\bm{\phi}}$ with respect to the sleep term).
+We use Adam optimizer and latent size of 2 for all of our experiments.
+Each algorithm takes roughly 2 hours on a single Nvidia GeForce RTX 3090 to complete one run of experiment.
+The training details of each algorithm are detailed below:
+AEVB.
 The batch size is set to 128.
@@
 We also have a weight of 5 multiplied to the prior density.
-IAF-VAE.
+AEVB-IAF.
 The batch size, number of epochs, learning rate, loss, KL divergence weight, and prior density weight are the same as VAE.
 The context size, i.e., the size of features used to initialize the flow layers for different datat point, is 10.
-AAE.
+AAEB.
 The batch size is set to 128.
@@
 The noise equals to 0 at epoch 150.
-DD-VAE.
+DDVI.
 The batch size is set to 128 for most experiments, except for semi-supervied experiments where the batch size is 1024.
@@
 The number of timesteps is 20 for unsupervised and clustering experiments and 100 for semi-supervised experiments.
-A.5 Genotype Analysis Experiments Details
+Appendix I Genotype Analysis Experiments Details
 Figure 6: Legend showing what ethnicity each color corresponds to in the 1000 Genomes dataset
@@
 The legend of the 1000 Genomes Visualization plot can be found at Figure 6.
-Method Pinwheel Swiss Roll Square Acc Latent NLL Acc Latent NLL Acc Latent NLL VAE $0.16\pm 0.01$ $86.9\pm 99.03$ $0.15\pm 0.01$ $360.41\pm 185.87$ $0.15\pm 0.01$ $63.3\pm 70.64$ IAF-VAE $0.19\pm 0.01$ $3.12\pm 1.16$ $0.16\pm 0.01$ $189.39\pm 57.48$ $0.18\pm 0.01$ $0.99\pm 0.11$ AAE $0.23\pm 0.00$ $2.30\pm 0.07$ $\bm{0.23\pm 0.01}$ $\bm{2.84\pm 0.08}$ $0.20\pm 0.01$ $1.46\pm 0.49$ HVAE w/ flow $0.20\pm 0.01$ $2.06\pm 0.13$ $0.20\pm 0.01$ $106.15\pm 15.30$ $0.19\pm 0.01$ $1.05\pm 0.06$ DiffVAE $\bm{0.23\pm 0.01}$ $\bm{1.37\pm 0.02}$ $\bm{0.23\pm 0.01}$ $\bm{2.80\pm 0.11}$ $\bm{0.24\pm 0.01}$ $\bm{0.86\pm 0.05}$
-Table 5: Unsupervised learning on CIFAR-10. We report accuracy using KNN (K=20) classifier and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
-Method Pinwheel Swiss Roll Square Latent NLL Avg Class Latent NLL Avg Class Latent NLL Avg Class Latent NLL Latent NLL Latent NLL VAE $7.67\pm 1.34$ $9.84\pm 2.87$ $49.64\pm 28.12$ $128.46\pm 47.08$ $1.51\pm 0.23$ $0.65\pm 0.37$ IAF-VAE $4.54\pm 0.59$ $4.46\pm 0.99$ $64.27\pm 73.87$ $102.16\pm 106.36$ $1.04\pm 0.07$ $0.07\pm 0.09$ AAE $2.33\pm 0.02$ $2.42\pm 0.06$ $3.76\pm 0.04$ $4.23\pm 0.13$ $1.45\pm 0.12$ $1.31\pm 0.20$ DiffVAE $\bm{1.49\pm 0.09}$ $\bm{0.35\pm 0.18}$ $\bm{3.35\pm 0.27}$ $\bm{2.28\pm 0.39}$ $\bm{0.97\pm 0.03}$ $\bm{-0.29\pm 0.04}$
-Table 6: Semi-supervised learning on CIFAR-10 (10,000 labels). We report average class latent negative log-likelihood and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
-Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square VAE $6.18\pm 0.93$ $82.96\pm 99.30$ $25.07\pm 62.41$ IAF-VAE $2.06\pm 0.21$ $12.13\pm 18.28$ $1.53\pm 0.82$ AAE $1.94\pm 0.09$ $3.43\pm 0.14$ $1.97\pm 0.70$ HVAE w/ flow $2.18\pm 0.13$ $4.17\pm 1.23$ $\mathbf{1.25\pm 0.08}$ DiffVAE $\mathbf{1.57\pm 0.10}$ $\mathbf{3.19\pm 0.07}$ $\mathbf{1.24\pm 0.13}$ VAE with Diffusion (standard ELBO) $6.09$ $510.53$ $265.25$
-Table 7: Unsupervised learning on MNIST, including the failed results of VAE with Diffusion
-Model Training Objective Approximating Family Sample-based Prior Auxiliary Variable Tasks Simplified Graphical Illustration VAE ELBO Diagonal Gaussian No No Density estimation $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ IAF-VAE ELBO Normalizing flow No Yes Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{x}$ AAE Adversarial training Adversarial generator Yes No Visualization $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ HVAE w/o and w/ flow ELBO Factorial Normal / Normalizing flow No Yes Density estimation / High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ ADGM ELBO Non-Gaussian No Yes Density estimation $\mathbf{x}\rightarrow\mathbf{a}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ LDM ELBO Diagonal Gaussian No Yes High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ LSGM ELBO+score matching Diagonal Gaussian No Yes High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ DiffVAE ELBO+sleep term Denoising diffusion Yes Yes Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{T}(\mathbf{y})\rightarrow\mathbf{z}_{0}(\mathbf{z})\rightarrow\mathbf{x}$
-Table 8: Comparison of DiffVAE to other relevant methods
+Method Pinwheel Swiss Roll Square ELBO Latent NLL ELBO Latent NLL ELBO Latent NLL AEVB $-12.96\pm 1.81$ $3.26\pm 0.60$ $-12.87\pm 4.55$ $6.25\pm 1.58$ $-7.91\pm 0.11$ $2.91\pm 0.17$ AEVB-IAF $-3.24\pm 0.16$ $1.71\pm 0.84$ $-4.03\pm 0.73$ $5.51\pm 0.51$ $\bm{-2.10\pm 0.31}$ $1.71\pm 0.77$ AAEB $N/A$ $1.70\pm 0.41$ $N/A$ $\bm{3.18\pm 0.22}$ $N/A$ $1.67\pm 0.17$ H-AEVB $-4.42\pm 0.46$ $\bm{1.69\pm 0.17}$ $-5.36\pm 0.77$ $5.74\pm 0.55$ $-2.86\pm 0.11$ $1.64\pm 0.09$ PIS ${-2.92\pm 1.23}$ ${3.61\pm 0.62}$ $-4.14\pm 0.49$ $7.14\pm 0.14$ $-4.85\pm 0.06$ $3.91\pm 0.06$ DDVI $\bm{-1.38\pm 0.44}$ ${1.75\pm 0.53}$ $\bm{-3.05\pm 0.65}$ ${5.66\pm 2.63}$ $-2.47\pm 0.30$ $\bm{1.58\pm 0.09}$
+Table 6: Unsupervised learning on CIFAR-10. We report ELBO and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
+Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $-17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $-13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $-5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $-5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $-4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\bm{-1.60\pm 0.29}$ $\bm{0.49\pm 0.01}$ $\bm{1.09\pm 0.05}$ $\bm{-4.13\pm 1.51}$ $\bm{0.47\pm 0.09}$ $\bm{2.29\pm 0.08}$ $\bm{-1.73\pm 0.64}$ $\bm{0.49\pm 0.01}$ $\bm{1.48\pm 0.02}$
+Table 7: Semi-supervised learning on CIFAR-10 (10,000 labels).
+We report ELBO, accuracy using KNN (K=20) classifier (Acc), and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
+Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB -- -- -- H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\bm{1.27\pm 0.21}$ $\bm{3.86\pm 1.17}$ $\bm{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
+Table 8: Unsupervised learning on MNIST, including the results of DDVI without the sleep term.
+Appendix J Data Licenses
+Yann LeCun and Corinna Cortes hold the copyright of MNIST dataset, which is a derivative work from original NIST datasets.
+CIFAR-10 is under MIT license.
+1000 Genomes dataset is under the creative commons Attribution-NonCommercial-ShareAlike 3.0 Unported licence.

### v4

@@
 Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks.
-We find that DDVI improves inference and learning in deep latent variable models across common benchmarks as well as on a motivating task in biology—inferring latent ancestry from human genomes—where it outperforms strong baselines on the Thousand Genomes dataset.
+We find that DDVI improves inference and learning in deep latent variable models across common benchmarks as well as on a motivating task in biology—inferring latent ancestry from human genomes—where it outperforms strong baselines on 1000 Genomes dataset.
 1 Introduction
 We are interested in amortized black-box variational inference problems of the form
-$\log\;p_{\bm{\theta}}(\mathbf{x})\geq\max_{\bm{\phi}}\;\;\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\;p_{\bm{\theta}}(\mathbf{x},\mathbf{z})-\log\;q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})\right]:=\max_{\bm{\phi}}\;\;\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}}),$ (1)
-in which we approximate the marginal likelihood $\log\;p_{\bm{\theta}}(\mathbf{x})$ of a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ with an evidence lower bound $\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ that is a function of an approximate variational posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$.
-We assume that $p_{\bm{\theta}}$ factorizes as $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\bm{\theta}}$ include Bayesian networks, topic models [4], variational autoencoders (VAEs), and broad classes of $p_{\bm{\theta}}$ defined via modern probabilistic programming frameworks [9].
-Maximizing $\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ over ${\bm{\phi}}$ yields a variational posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ that approximates $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ as well as a tight bound on $\log\;p_{\bm{\theta}}(\mathbf{x})$ that serves as a learning objective for $p_{\bm{\theta}}$.
-The approximation gap $\log\;p_{\bm{\theta}}(\mathbf{x})-\max_{\bm{\phi}}\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ equals precisely $\min_{\bm{\phi}}\text{KL}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$, which motivates the design of expressive classes of posteriors $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ that reduce this gap.
-Recent efforts leverage modern generative models—including normalizing flows [28, 16] and generative adversarial networks [8, 20]—as expressive model families for $q_{\bm{\phi}}$ that tighten the ELBO.
-This work seeks to further improve variational inference via expressive posteriors based on diffusion models [12, 37].
-Diffusion methods have become the de-facto standard for high-quality image synthesis [30].
-Here, we use diffusion in latent space to parameterize $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$.
-We train this distribution with a denoising diffusion-like objective that does not involve adversarial training [20] or constrained invertible normalizing flow architectures [16].
-Samples from $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ are obtained via iterative refinement of $\mathbf{z}$, starting from a Gaussian distribution, and gradually forming one that is multi-modal and complex.
+$\log\;p_{\boldsymbol{\theta}}(\mathbf{x})\geq$ $\max_{\boldsymbol{\phi}}\;\;\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\;p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})-\log\;q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})\right]$
+$:=\max_{\boldsymbol{\phi}}\;\;\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}}),$ (1)
+in which we approximate the marginal likelihood $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ of a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ with an evidence lower bound $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ that is a function of an approximate variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$.
+We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks, topic models (Blei, Ng, and Jordan 2003), variational autoencoders (VAEs), and broad classes of $p_{\boldsymbol{\theta}}$ defined via modern probabilistic programming frameworks (Gordon et al.
+2014).
+Maximizing $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ over ${\boldsymbol{\phi}}$ yields a variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that approximates $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ as well as a tight bound on $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ that serves as a learning objective for $p_{\boldsymbol{\theta}}$.
+The approximation gap $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})-\max_{\boldsymbol{\phi}}\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ equals precisely $\min_{\boldsymbol{\phi}}\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$, which motivates the design of expressive classes of posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that reduce this gap.
+Recent efforts leverage modern generative models—including normalizing flows (Rezende and Mohamed 2015; Kingma et al.
+2016) and generative adversarial networks (Goodfellow et al.
+2014; Makhzani et al.
+2015)—as expressive model families for $q_{\boldsymbol{\phi}}$ that tighten the ELBO.
+This work seeks to further improve variational inference via expressive posteriors based on diffusion models (Ho, Jain, and Abbeel 2020; Song et al.
+2020).
+Diffusion methods have become the de-facto standard for high-quality image synthesis (Rombach et al.
+2022; Gokaslan et al.
+2024).
+Here, we use diffusion in latent space to parameterize $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$.
+We train this distribution with a denoising diffusion-like objective that does not involve adversarial training (Makhzani et al.
+2015) or constrained invertible normalizing flow architectures (Kingma et al.
+2016).
+Samples from $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ are obtained via iterative refinement of $\mathbf{z}$, starting from a Gaussian distribution, and gradually forming one that is multi-modal and complex.
 […]
-Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion as well as a learning objective inspired by the wake-sleep algorithm [11] that can be interpreted as a regularized variational inference.
+Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion and a learning objective inspired by the wake-sleep algorithm (Hinton et al.
+1995) that implements regularized variational inference.
 We also derive extensions of our method to semi-supervised learning and clustering.
@@
 We evaluate DDVI on synthetic benchmarks and on a real problem in biological data analysis—inferring human ancestry from genetic data.
-Our method outperforms strong baselines on the Thousand Genomes dataset [35] and learns a low-dimensional latent space that preserves biologically meaningful structure [10].
+Our method outperforms strong baselines on 1000 Genomes dataset (Siva 2008) and learns a low-dimensional latent space that preserves biologically meaningful structure (Haghverdi, Buettner, and Theis 2015).
 Contributions.
@@
 Deep Latent Variable Models
-Latent variable models (LVMs) $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ are usually fit by optimizing the evidence lower bound (ELBO)
-$\log p_{\bm{\theta}}(\mathbf{x})\geq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z})),$ (2)
+Latent variable models (LVMs) $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ are usually fit by optimizing the evidence lower bound (ELBO)
+$\log p_{\boldsymbol{\theta}}(\mathbf{x})\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z})),$
 which serves as a tractable surrogate for the marginal log-likelihood (MLL).
-The gap between the MLL and the ELBO equals precisely $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO [14].
-Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods [19], MCMC-based methods [31], normalizing flows [28]—or improved learning objectives—e.g., adversarial or sample-based losses [20, 48, 33, 34].
-The wake-sleep algorithm [11] optimizes an alternative objective
-$\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})),$ (3)
+The gap between the MLL and the ELBO equals precisely $\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO (Kingma and Welling 2013).
+Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods (Maaløe et al.
+2016), MCMC-based methods (Salimans, Kingma, and Welling 2015), normalizing flows (Rezende and Mohamed 2015)—or improved learning objectives—e.g., adversarial or sample-based losses (Makhzani et al.
+2015; Zhao, Song, and Ermon 2017; Si, Bishop, and Kuleshov 2022; Si et al.
+2023).
+The wake-sleep algorithm (Hinton et al. 1995) optimizes an alternative objective
+$\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]-\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})),$
 in which the KL divergence term is reversed.
-The learning procedure for wake-sleep involves alternating between "wake" phases where the recognition model is updated and "sleep" phases where the generative model is refined.
+The learning procedure for wake-sleep involves alternating between ”wake” phases where the recognition model is updated and ”sleep” phases where the generative model is refined.
 Denoising Diffusion Models
@@
 We obtain $\mathbf{y}_{1:T}$ by applying a Markov chain $q(\mathbf{y}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where we define $\mathbf{y}_{0}=\mathbf{x}_{0}$ for convenience.
-When $\mathbf{x}_{0}$ is a continuous vector, a standard choice of transition kernel is $q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{{\bm{\alpha}}_{t}}\mathbf{y}_{t-1},\sqrt{1-{\bm{\alpha}}_{t}}\mathbf{I})$, which is a Gaussian centered around a copy of $\mathbf{y}_{t-1}$ to which we added noise following a schedule $10<{\bm{\alpha}}_{1}<{\bm{\alpha}}_{2}<...<{\bm{\alpha}}_{T}=$ 1.
-A diffusion model can then be represented as a latent variable distribution $p(\mathbf{x}_{0},\mathbf{y}_{1:T})$ that factorizes as $p(\mathbf{x}_{0},\mathbf{y}_{1:T})=p(\mathbf{y}_{T})\prod_{t=0}^{T-1}p_{\bm{\theta}}(\mathbf{y}_{t}\mid\mathbf{y}_{t+1})$ (again using $\mathbf{y}_{0}$ as shorthand for $\mathbf{x}_{0}$).
+When $\mathbf{x}_{0}$ is a continuous vector, a standard choice of transition kernel is $q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{I})$, which is a Gaussian centered around a copy of $\mathbf{y}_{t-1}$ to which we added noise following a schedule $10<{\boldsymbol{\alpha}}_{1}<{\boldsymbol{\alpha}}_{2}<...<{\boldsymbol{\alpha}}_{T}=$ 1.
+A diffusion model can then be represented as a latent variable distribution $p(\mathbf{x}_{0},\mathbf{y}_{1:T})$ that factorizes as $p(\mathbf{x}_{0},\mathbf{y}_{1:T})=p(\mathbf{y}_{T})\prod_{t=0}^{T-1}p_{\boldsymbol{\theta}}(\mathbf{y}_{t}\mid\mathbf{y}_{t+1})$ (again using $\mathbf{y}_{0}$ as shorthand for $\mathbf{x}_{0}$).
 This model seeks to approximate the reverse of the forward diffusion $q$ and map noise $\mathbf{y}_{T}$ into data $\mathbf{x}_{0}$.
-The true reverse of the process $q$ cannot be expressed in closed form; as such, we parameterize $p_{\bm{\theta}}$ with ${\bm{\theta}}$ trained by maximizing the ELBO:
-$\log p_{\bm{\theta}}(\mathbf{x}_{0})\geq\mathbb{E}_{q}\Bigr{[}\log p_{\bm{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})-\sum_{t=2}^{T}D_{\mathrm{KL}}(q_{t}||p_{t})\Bigr{]}-D_{\mathrm{KL}}(q(\mathbf{x}_{T}|\mathbf{x}_{0})||p(\mathbf{x}_{T}))$ (4)
-where $q_{t},p_{t}$ denote the distributions $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ and $p_{\bm{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, respectively.
-3 Variational Inference With Denoising Diffusion Models
+The true reverse of the process $q$ cannot be expressed in closed form; as such, we parameterize $p_{\boldsymbol{\theta}}$ with ${\boldsymbol{\theta}}$ trained by maximizing the ELBO:
+$\log p_{\boldsymbol{\theta}}(\mathbf{x}_{0})\geq$ $\mathbb{E}_{q}\Bigr{[}\log p_{\boldsymbol{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})-\sum_{t=2}^{T}\text{KL}(q_{t}||p_{t})\Bigr{]}$ (2)
+$-\text{KL}(q(\mathbf{x}_{T}|\mathbf{x}_{0})||p(\mathbf{x}_{T}))$
+where $q_{t},p_{t}$ denote the distributions $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ and $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, respectively.
+3 Variational Inference With
+Denoising Diffusion Models
 We introduce denoising diffusion variational inference (DDVI), which improves variational inference with diffusion-based techniques.
-The goal of DDVI is to fit a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$.
-We assume that $p_{\bm{\theta}}$ factorizes as $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\bm{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) [14].
-Our approach is comprised of two key components: a family of approximate posteriors $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ based on diffusion and a learning objective that can be seen as a regularized ELBO.
-The $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ iteratively refines latents $\mathbf{z}$, starting from a Gaussian distribution.
-The learning objective trains $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to reverse a used-specified forward diffusion process.
-3.1 Diffusion-Based Variational Posteriors
+The goal of DDVI is to fit a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$.
+We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) (Kingma and Welling 2013).
+Our approach is comprised of three components:
+  1. 1.
+    A modeling family of approximate posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ based on diffusion;
+  2. 2.
+    A learning objective formed by a regularized ELBO;
+  3. 3.
+    An optimization algorithm inspired by wake-sleep.
+The $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ iteratively refines latents $\mathbf{z}$, starting from a Gaussian distribution.
+The learning objective trains $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to reverse a used-specified forward diffusion process.
+3.1 Modeling Family: Diffusion-Based Posteriors
 […]
 We choose the latent $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ to be a vector of $T$ variables that represent progressively simplified versions of $\mathbf{z}$, with $\mathbf{y}_{T}$ corresponding to a simple distribution (e.g., a Gaussian).
-The model $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\bm{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$ transforms $\mathbf{y}_{T}$ into $\mathbf{z}$ via iterative refinement.
-To sample from $q_{\bm{\phi}}$, we first sample $\mathbf{y}_{T}$—this is an easier task since we can define $\mathbf{y}_{T}$ to have a simple (e.g., Gaussian) distribution—and then by sampling from the denoising model $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\bm{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$.
-We define the relationship between $\mathbf{y}$ and $\mathbf{z}$ via a forward diffusion process $r(\mathbf{y}|\mathbf{z})=r(\mathbf{y}_{1:T}|\mathbf{z})=r(\mathbf{y}_{1}|\mathbf{z})\prod_{t=1}^{T-1}r(\mathbf{y}_{t+1}|\mathbf{y}_{t})$, which transforms $\mathbf{z}$—the latent whose intractable posterior we seek to approximate—into $\mathbf{y}_{T}$, whose posterior is easier to model.
-Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes [2].
-The model $q_{\bm{\phi}}$ is trained to approximately reverse this forward diffusion process.
-3.2 The Wake-Sleep Regularized ELBO
-The standard approach to fit auxiliary-variable generative models [19] is to apply the ELBO twice:
-$\log p_{\bm{\theta}}(\mathbf{x})$ $\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$ (5)
-$\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))-\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$ (6)
-$=~{}\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$ (7)
-In Equation 5, we applied the ELBO over $\mathbf{z}$, and in Equation 6 we applied the ELBO again over the latent $\mathbf{y}$ of $q$.
-Notice that the gap between the ELBO and $\log p_{\bm{\theta}}(\mathbf{x})$ is $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))+\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$.
+The model $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\boldsymbol{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$ transforms $\mathbf{y}_{T}$ into $\mathbf{z}$ via iterative refinement.
+To sample from $q_{\boldsymbol{\phi}}$, we first sample $\mathbf{y}_{T}$—this is an easier task since we can define $\mathbf{y}_{T}$ to have a simple (e.g., Gaussian) distribution—and then by sampling from the denoising model $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\boldsymbol{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$.
+We define the relationship between $\mathbf{y}$ and $\mathbf{z}$ via a forward diffusion process $r(\mathbf{y}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1}|\mathbf{z},\mathbf{x})\prod_{t=1}^{T-1}r(\mathbf{y}_{t+1}|\mathbf{y}_{t},\mathbf{x})$, which transforms $\mathbf{z}$—the latent whose intractable posterior we seek to approximate—into $\mathbf{y}_{T}$, whose posterior is easier to model (possibly conditioned on $\mathbf{x}$).
+Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes (Austin et al.
+2021).
+The model $q_{\boldsymbol{\phi}}$ is trained to approximately reverse this forward diffusion process.
+3.2 Learning Objective: A Markovian ELBO
+The standard approach to fit auxiliary-variable generative models (Maaløe et al. 2016) is to apply the ELBO twice:
+$\log p_{\boldsymbol{\theta}}(\mathbf{x})$ $\geq\log p_{\boldsymbol{\theta}}(\mathbf{x})-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$ (3)
+$\geq\log p_{\boldsymbol{\theta}}(\mathbf{x})-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$ (4)
+$~{}~{}~{}~{}-\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$
+$=~{}\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]$ (5)
+$~{}~{}~{}~{}-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$
+In Equation (3), we applied the ELBO over $\mathbf{z}$, and in Equation (4) we applied the ELBO over the latent $\mathbf{y}$ of $q$ (see Appendix J for the derivation).
+Notice that the gap between the ELBO and $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ is $\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))+\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$.
 Thus, if we correctly match $q$ and $r$, we will achieve a tight bound.
-Wake-Sleep Regularization
-Notice that optimizing $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{z},\mathbf{x})p(\mathbf{z}))$ in Equation 7 involves sampling from the approximate reverse process $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ to match the true reverse process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$: this is the opposite of diffusion training, where we would sample from $r$ to fit $q$.
-We have found the standard ELBO to insufficient to learn a good $q_{\bm{\phi}}$ that reverses the noising process $r$, as illustrated in Table 8.
-Instead, our strategy is to introduce a modified ELBO that yields diffusion-like training.
-This objective is the ELBO in Equation 7 augmented with an additional regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$.
-$\log p_{\bm{\theta}}(\mathbf{x})\geq\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / recons.
-term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]}_{\text{sleep term $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$}}$ (8)
-The optimization of the regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ is similar to the sleep phase of the wake-sleep algorithm, and closely resembles diffusion model training (see below).
-As in wake-sleep, $\mathcal{L}_{\text{sleep}}$ is optimized over ${\bm{\phi}}$ only, and the $\mathbf{x}$ are sampled from the model.
-3.3 Optimizing the Regularized ELBO
-Computing $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ still involves intractable distributions $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}),q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$.
-To optimize $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$, we introduce another lower bound $\mathcal{L}_{\text{diff}}({\bm{\phi}})$, which we call the denoising diffusion loss (for reasons that will become apparent shortly):
-$\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ $=-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x},\mathbf{z})}[\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})$ (9)
-$\geq\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x},\mathbf{z})}[\mathbb{E}_{r}[\log\frac{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{z},\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})=\mathcal{L}_{\text{diff}}({\bm{\phi}})$ (10)
-In Equation 10, we applied the ELBO with $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ playing the role of the variational posterior over the latent $\mathbf{y}$ in $q_{\bm{\phi}}$; $\bar{H}(p_{\bm{\theta}})$ is the expected conditional entropy of $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on ${\bm{\phi}}$.
-We define the DDVI learning objective $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ as the sum of the aforementioned terms:
-$\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})=\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\mathcal{L}_{\text{diff}}({\bm{\phi}})$ (11)
+Analyzing the ELBO
+Optimizing Equation (5) requires tractably dealing with the prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}}):=-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$, which we equivalently rewrite as:
+$\mathcal{L}_{\text{reg}}=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log(r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))]+H(q).$ (6)
+We can expand the first term by leveraging the Markov structure of $r,q$ to rewrite $\mathcal{L}_{\text{reg}}$ as the likelihood of samples from the reverse diffusion process $q$ under the forward process $r$.
+$\mathcal{L}_{\text{reg}}=\sum_{t=1}^{T}\mathbb{E}_{q}[\log(r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x})]+\mathbb{E}_{q}[\log p(\mathbf{z})]+H(q),$
+where $\mathbf{y}_{0}:=\mathbf{z}$.
+We refer to optimizing the Markovian ELBO as unregularized DDVI, the first instance of our method.
+The noise process $r$ defines prior regularization terms for each $\mathbf{y}_{t}$.
+This provides extra supervision for learning $q$ in the form of trajectories from latents $\mathbf{y}_{T}$ to $\mathbf{y}_{1}$; this extra supervision helps $q$ fit complex non-Gaussian posteriors.
+The term $H(q)=-\sum_{t=1}^{T+1}\mathbb{E}_{q}[\log q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]$ denotes the entropy.
+For example, when each term $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ is Gaussian, it is computed as
+$H(q)=$ $\sum_{t=1}^{T+1}\mathbb{E}_{q}\left[\frac{d}{2}\left(1+\log(2\pi)\right)+\frac{1}{2}\log\left|\Sigma_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x})\right|\right]$
+where $d$ is the dimension of $\mathbf{y}$ and we use the notation $\mathbf{y}_{T+1}=\mathbf{x}$.
+It is also common to leave the variance $\Sigma_{\boldsymbol{\phi}}$ fixed, in which case $H(q)$ is a constant.
+3.3 Refining the Objective: A Regularized ELBO
+Notice that optimizing $\mathcal{L}_{\text{reg}}$ involves sampling from the approximate reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ to match the true reverse process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$: this is the opposite of diffusion training, where we would sample from $r$ to fit $q$.
+This type of on-policy learning of $q$ has been studied in the context of approximate inference (Zhang and Chen 2021); however, it requires backpropagating through $T$ samples, which may hamper training, and it optimizes a mode-covering divergence that may struggle to fit complex $p(\mathbf{z})$.
+Adding Wake-Sleep Regularization to the ELBO
+We propose to further improve the ELBO via off-policy diffusion-like training.
+Our new objective is the ELBO in Equation (5) augmented with a regularizer $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$.
+$\log p_{\boldsymbol{\theta}}(\mathbf{x})\geq$ $\underbrace{\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / recons.
+term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}$ (7)
+$\underbrace{-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))}_{\text{priorregularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}$
+$\underbrace{-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))]}_{\text{sleep term $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$}}$
+The optimization of the regularizer $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ is similar to the sleep phase of wake-sleep, and closely resembles diffusion model training (see below).
+As in wake-sleep, $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ is optimized over ${\boldsymbol{\phi}}$ only, the $\mathbf{x}$ are sampled from $p$.
+From Wake-Sleep to Diffusion Regularization
+Computing $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ still involves intractable distributions $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}),q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$.
+To optimize $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$, we introduce another lower bound $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, which we call the denoising diffusion loss (for reasons that will become apparent shortly):
+$\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ $=-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))]$ (8)
+$=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}[\log{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\boldsymbol{\theta}})$
+$\geq\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}\left[\mathbb{E}_{r}[\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{z},\mathbf{x})}]\right]+\bar{H}(p_{\boldsymbol{\theta}})$
+$=\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$
+In Equation (8), we applied the ELBO with $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ playing the role of the variational posterior over the latent $\mathbf{y}$ in $q_{\boldsymbol{\phi}}$; $\bar{H}(p_{\boldsymbol{\theta}})$ is the expected conditional entropy of $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on ${\boldsymbol{\phi}}$.
+We can further simplify $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ by leveraging the Markov structure of the forward and reverse processes $r,q$.
+Recall that each $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ can be a vector of $T$ latents, which we also denote as $\mathbf{y}_{1:T}$, and that $r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x})$, where $\mathbf{y}_{0}=\mathbf{z}$ and also $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}).$
+We may use the Markov structure in $q,r$ to rewrite $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ as a sum of $T$ terms, one per Markov step.
+The derivation is identical to that used to obtain the ELBO of a diffusion model (Sohl-Dickstein et al.
+2015), and yields an expression of the same form:
+$\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})=$ $~{}\mathbb{E}_{r}\left[\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}\text{KL}(r_{t}||q_{t})\right]$ (9)
+$-\text{KL}(r(\mathbf{y}_{T}|\mathbf{z})||q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})).$
+where $r_{t},q_{t}$ denote the distributions $r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0},\mathbf{x})$ and $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ (see Appendix K for the derivation).
+Regularized DDVI Objective
+We define the full DDVI objective $\mathcal{L}_{\text{ddvi}}$ to be the sum of the aforementioned terms:
+$\mathcal{L}_{\text{ddvi}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})=\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$
 Terms $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{diff}}$ may be weighted by hyper-parameters $\beta_{\text{reg}},\beta_{\text{diff}}>0$, as in the $\beta$-VAE framework.
-In our experiments, assume $\beta_{\text{reg}}=\beta_{\text{diff}}=1$ unless otherwise specified.
-Note that since $\mathcal{L}_{\text{diff}}\leq\mathcal{L}_{\text{sleep}}\leq 0$, $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ is a valid lower bound on $\log p_{\bm{\theta}}(\mathbf{x})$ that is tight when $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})=p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$.
-Wake-Sleep Optimization
-We optimize our bound on $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ using gradient descent by alternating between ELBO optimization and taking sleep steps (see Section 5).
-Note that by maximizing $\mathcal{L}_{\text{diff}}$, we fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ via the forward KL divergence; similarly, by optimizing $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{reg}}$ (the ELBO), we fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ via the reverse KL divergence.
-Thus, optimizing $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ encourages $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to approximate $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, and when the two are equal, the bound on $\log p_{\bm{\theta}}(\mathbf{x})$ is tight.
-Simplifying Wake-Sleep
-We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$.
-In this case, $\mathcal{L}_{\text{diff}}$ requires only sampling from $p(\mathbf{z})$, and the entire loss $\mathcal{L}$ can be optimized end-to-end using gradient descent.
-This algorithm is a simpler (there is no separate sleep phase); however, $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may not perfectly approximate $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, hence $\mathcal{L}$ may no longer be a tight bound.
-We report results in Appendix B.
-3.4 Combining The Wake-Sleep Regularized ELBO With Diffusion Models
-Lastly, we can further simplify $\mathcal{L}_{\text{diff}}$ by leveraging the Markov structure of the forward and reverse processes $r,q$.
-Recall that each $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ can be a vector of $T$ latents, which we also denote as $\mathbf{y}_{1:T}$, and that $r(\mathbf{y}_{1:T}|\mathbf{z})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where $\mathbf{y}_{0}=\mathbf{z}$ and the $\mathbf{y}_{1:T}$ are increasingly noised versions of $\mathbf{y}_{0}$.
-Similarly, $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ is an approximate reverse diffusion process $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ that is trained to reverse $r$.
-In order to fit $q_{\bm{\phi}}$, we form a lower bound $\mathcal{L}_{\text{diff}}(\mathbf{x},{\bm{\phi}})$ on the sleep term $\mathbb{E}_{p(\mathbf{x},\mathbf{z})}\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}$ in Equation 9.
-This bound is identical to the ELBO of a diffusion model, and has the same derivation.
-$\mathcal{L}_{\text{diff}}=$ $~{}\mathbb{E}_{r}\left[\log q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}D_{\mathrm{KL}}(r_{t}||q_{t})\right]-D_{\mathrm{KL}}(r(\mathbf{y}_{T}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})).$ (12)
-where $r_{t},q_{t}$ denote the distributions $r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0})$ and $q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$.
-Parameterizing Diffusion-Based Encoders
-A common type of noising process compatible with this bound when $\mathbf{z}$ is continuous is Gaussian diffusion, where we define $r(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-{\bm{\alpha}}_{t}}\mathbf{y}_{t-1},{\bm{\alpha}}_{t}\mathbf{I})$ for a suitable schedule $({\bm{\alpha}}_{t})_{t=1}^{T}$.
-We then adopt the parameterization $q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})=\mathcal{N}(\mathbf{y}_{t-1};\mu_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t),\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t))$.
-It is then common to parameterize $q_{\bm{\phi}}$ with a noise prediction network ${\bm{\epsilon}}_{\bm{\phi}}$ [12]; the sum of KL divergences can be approximated by $\mathbb{E}_{t,{\bm{\epsilon}}_{t}\sim r(\mathbf{y}_{0},t)}||{\bm{\epsilon}}_{t}-{\bm{\epsilon}}_{\bm{\phi}}(\sqrt{\bar{{\bm{\alpha}}_{t}}}\mathbf{y}_{0}+\sqrt{1-\bar{{\bm{\alpha}}_{t}}}{\bm{\epsilon}}_{t},\mathbf{x},t)||^{2}$.
-Lastly, in order to use diffusion-based encoders with our ELBO-based objective, we need to show that we can tractably compute $-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$, which we equivalently rewrite as:
-$-D_{\mathrm{KL}}(q_{\bm{\phi}}||r\cdot p)=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log(r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))]+H(q).$
-The term $H(q)$ denotes the entropy. We approximate the first term using Monte-Carlo; the entropy is computed as
-$H(q)=-\sum_{t=1}^{T+1}\mathbb{E}_{q}[\log q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]=\sum_{t=1}^{T+1}\mathbb{E}_{q}\left[\frac{d}{2}\left(1+\log(2\pi)\right)+\frac{1}{2}\log\left|\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x})\right|\right]$ (13)
-where $d$ is the dimension of $\mathbf{y}$ and we use the notation $\mathbf{y}_{T+1}=\mathbf{x}$.
-The right-hand term can be approximated using Monte Carlo; it is also common to leave the variance $\Sigma_{\bm{\phi}}$ fixed (as we typically do in our experiments), in which case $H(q)$ is a constant.
+In our experiments, $\beta_{\text{reg}}=\beta_{\text{diff}}=1$ unless otherwise specified.
+Note that since $\mathcal{L}_{\text{diff}}\leq\mathcal{L}_{\text{sleep}}\leq 0$, $\mathcal{L}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ is a valid lower bound on $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ that is tight when $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})=p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$.
+3.4 Optimization: Extending Wake-Sleep
+We may optimize $\mathcal{L}_{\text{ddvi}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ using gradient descent by alternating between ELBO optimization and taking sleep steps (see Appendix A for full details):
+  1. 1.
+    Sample $\mathbf{x}$ from data, take gradient step on ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ optimizing $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ (the “wake” step);
+  2. 2.
+    Sample $\mathbf{z},\mathbf{x}$ from $p_{\boldsymbol{\theta}}$ and take a gradient step on ${\boldsymbol{\phi}}$ optimizing $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ (the “sleep” step).
+Again, terms may be weighted by $\beta_{\text{reg}},\beta_{\text{diff}}>0$.
+Note that by maximizing $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, we fit $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ via the forward KL divergence; similarly, by optimizing $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{reg}}$ (the ELBO), we fit $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ via the reverse KL divergence.
+Thus, optimizing $\mathcal{L}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ encourages $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to approximate $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$, and when the two are equal, the bound $\mathcal{L}_{\text{ddvi}}$ on $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ is tight.
+Simplified Wake-Sleep
+We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$.
+This scenario admits the following optimization procedure:
+  1. 1.
+    Sample $\mathbf{x}$ from data and compute gradient on ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ optimizing $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$.
+  2. 2.
+    Sample $\mathbf{z}$ from $p(\mathbf{z})$ and compute gradient on ${\boldsymbol{\phi}}$ optimizing $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$; take step on weighted sum of both gradients.
+In this case, $\mathcal{L}_{\text{diff}}$ requires only sampling from $p(\mathbf{z})$, and the entire loss $\mathcal{L}_{\text{ddvi}}$ can be optimized end-to-end using gradient descent.
+This algorithm is simpler (there is no separate sleep phase); however, $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ may not perfectly approximate $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ when $r(\mathbf{y}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$, hence $\mathcal{L}$ may no longer be a tight bound.
+Practical Considerations
+A common type of noising process compatible with this bound when $\mathbf{z}$ is continuous is Gaussian diffusion, where we define $r(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},{\boldsymbol{\alpha}}_{t}\mathbf{I})$ for a suitable schedule $({\boldsymbol{\alpha}}_{t})_{t=1}^{T}$.
+We then adopt the parameterization $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})=\mathcal{N}(\mathbf{y}_{t-1};\mu_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x},t),\Sigma_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x},t))$.
+It is then common to parameterize $q_{\boldsymbol{\phi}}$ with a noise prediction network ${\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}$ (Ho, Jain, and Abbeel 2020); the sum of KL divergences can be approximated by $\mathbb{E}_{t,{\boldsymbol{\epsilon}}_{t}\sim r(\mathbf{y}_{0},t)}||{\boldsymbol{\epsilon}}_{t}-{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\sqrt{\bar{{\boldsymbol{\alpha}}_{t}}}\mathbf{y}_{0}+\sqrt{1-\bar{{\boldsymbol{\alpha}}_{t}}}{\boldsymbol{\epsilon}}_{t},\mathbf{x},t)||^{2}$.
+Other extensions include discrete denoising processes (Austin et al.
+2021; Sahoo et al.
+2024; Schiff et al.
+2024).
+In the wake-sleep setting, we know both endpoints $\mathbf{y}_{T}\sim q(\cdot|\mathbf{x})$ and $\mathbf{y}_{0}=\mathbf{z}$ of the diffusion process, opening the possibility for applying optimal transport techniques (Cuturi 2013; De Bortoli et al.
+2021).
 4 Extensions
@@
 Following Makhzani et al.
-[20], we extend our algorithm to the semi-supervised learning setting where some data points have labels denoted by $l$.
-We assume the user provides a model of the form $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\bm{\theta}}(\mathbf{z}|l)p(l)$; we set the variational distributions to $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\bm{\phi}}(\mathbf{y}|\mathbf{x}),q_{\bm{\phi}}(l|\mathbf{x})$.
-In this setting, we consider two cases, depending on whether the label is observed [15].
-We extend Equation 8 to incorporate the label $l$ corresponding to a data point as follows:
-$\mathcal{L}_{\text{semi}}=$ $~{}\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},l)]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}|l))$ (14)
-$-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}\left[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x},l)||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},l))\right]$
-When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\bm{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\bm{{\bm{\theta}}}},{\bm{\phi}})+D_{\mathrm{KL}}(q_{\bm{\phi}}(l|\mathbf{x})||p(l))$.
-Therefore, we can conclude a marginal likelihood on our dataset as follows: ~$\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\bm{{\bm{\theta}}}},{\bm{{\bm{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\bm{{\bm{\theta}}}},{\bm{{\bm{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
-We also want to guarantee that all model parameters can be learned in all cases, including $q_{\bm{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference.
-Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: ~$\tilde{\mathcal{L}}_{\text{semi}_{\bm{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\bm{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\bm{\phi}}(l|\mathbf{x})\right]$
+(2015), we extend our algorithm to the semi-supervised learning setting where some data points have labels denoted by $l$.
+We assume the user provides a model of the form $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\boldsymbol{\theta}}(\mathbf{z}|l)p(l)$; we set the variational distributions to $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x}),q_{\boldsymbol{\phi}}(l|\mathbf{x})$.
+In this setting, we consider two cases, depending on whether the label is observed (Kingma et al.
+2014).
+We extend Equation (7) to incorporate the label $l$ corresponding to a data point as follows:
+$\mathcal{L}_{\text{semi}}=$ $~{}\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},l)]$ (10)
+$-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x},l)||p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z}|l))$
+$-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}\left[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x},l)||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x},l))\right]$
+When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\boldsymbol{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{\phi}})+\text{KL}(q_{\boldsymbol{\phi}}(l|\mathbf{x})||p(l))$.
+Therefore, we can conclude a marginal likelihood on our dataset as follows: ~$\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
+We also want to guarantee that all model parameters can be learned in all cases, including $q_{\boldsymbol{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference.
+Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: ~$\tilde{\mathcal{L}}_{\text{semi}_{\boldsymbol{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\boldsymbol{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\boldsymbol{\phi}}(l|\mathbf{x})\right]$
 4.2 Clustering
@@
 We propose two distinct strategies.
-In the first approach, we simply fir a model in which $p_{\bm{{\bm{\theta}}}}(\mathbf{z})$ as a mixture of desired priors.
-The means of these priors are characterized by ${\bm{{\bm{\theta}}}}$.
+In the first approach, we simply formulate a model in which $p_{\boldsymbol{{\boldsymbol{\theta}}}}(\mathbf{z})$ is a mixture of desired priors.
+The means of these priors are characterized by ${\boldsymbol{{\boldsymbol{\theta}}}}$.
 From these means, cluster membership, denoted as $\mathbf{c}$ can be deduced.
 This approach requires no alteration to the existing learning objective.
-Alternatively, the second method retains the original prior but introduces an additional cluster latent variable $\mathbf{c}$ where $\sum_{i}c_{i}=1$.
-Thus, the model can be specified as $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{c})=p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})p(\mathbf{c})$ with $p(\mathbf{c})=Dir({\bm{{\bm{\epsilon}}}})$.
-Consequently, the variational distributions become $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{c},\mathbf{x}),q_{\bm{\phi}}(\mathbf{y},\mathbf{c}|\mathbf{x})$.
-This reformulates the learning objective as:
-$\mathcal{L}_{\text{clus}}(\mathbf{x})=$ $~{}\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z},\mathbf{c}))-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}\left[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))\right]$ (15)
-Figure 2: Unsupervised visualization on MNIST using three different priors (pinwheel, swiss roll, and square). Each color indicates a class.
-Method Pinwheel Swiss Roll Square ELBO MMD Latent NLL ELBO MMD Latent NLL ELBO MMD Latent NLL AEVB $-12.13\pm 0.41$ $0.77\pm 0.04$ $1.68\pm 0.31$ $-14.80\pm 0.23$ $0.78\pm 0.17$ $5.65\pm 1.58$ $-7.85\pm 0.29$ $1.10\pm 0.66$ $2.78\pm 0.61$ AEVB-IAF $-4.19\pm 0.05$ ${0.77\pm 0.00}$ $1.64\pm 0.73$ $-5.10\pm 0.30$ $0.61\pm 0.15$ $4.43\pm 1.09$ $-3.97\pm 0.22$ $0.75\pm 0.12$ $1.68\pm 0.27$ AAEB $N/A$ $0.68\pm 0.02$ $1.54\pm 0.19$ $N/A$ $\bm{0.52\pm 0.03}$ ${3.34\pm 0.16}$ $N/A$ $0.80\pm 0.02$ $2.46\pm 0.46$ H-AEVB $-7.03\pm 3.13$ ${0.74\pm 0.02}$ $2.25\pm 3.02$ $-7.21\pm 4.62$ ${0.70\pm 0.22}$ $4.04\pm 4.62$ $-5.71\pm 3.05$ ${0.76\pm 0.21}$ ${2.22\pm 2.03}$ PIS $-7.83\pm 0.64$ $0.75\pm 0.14$ $6.50\pm 1.11$ $-9.83\pm 0.61$ $0.61\pm 0.03$ $\bm{2.40\pm 1.01}$ $-7.06\pm 0.06$ $0.77\pm 0.04$ $3.67\pm 0.08$ DDVI $\bm{-3.88\pm 0.96}$ $\bm{0.67\pm 0.04}$ $\bm{1.27\pm 0.21}$ $\bm{-5.03\pm 0.58}$ $0.62\pm 0.33$ $3.86\pm 0.17$ $\bm{-3.79\pm 0.14}$ $\bm{0.66\pm 0.07}$ $\bm{1.56\pm 0.09}$
+Alternatively, the second method retains the original prior, but introduces an additional latent cluster variable $\mathbf{c}$ where $\sum_{i}c_{i}=1$.
+Thus, the model can be specified as $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{c})=p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})r(\mathbf{y}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})p(\mathbf{c})$ with $p(\mathbf{c})=Dir({\boldsymbol{{\boldsymbol{\epsilon}}}})$.
+Consequently, the variational distributions become $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{c},\mathbf{x}),q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{c}|\mathbf{x})$.
+This yields the objective:
+$\mathcal{L}_{\text{clus}}(\mathbf{x})=$ $~{}\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},\mathbf{c})]$ (11)
+$-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z},\mathbf{c}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z},\mathbf{c}))$
+$-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}\left[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))\right]$
+Expectations over small numbers of classes $\mathbf{c}$ are done analytically; larger $\mathbf{c}$ require backpropagating through discrete sampling (Jang, Gu, and Poole 2016; Sahoo et al. 2023).
+Figure 2: Unsupervised visualization on MNIST using three priors (pinwheel, swiss roll, and square). Each color indicates a class.
+Method Pinwheel Swiss Roll Square ELBO MMD Latent NLL ELBO MMD Latent NLL ELBO MMD Latent NLL AEVB $-12.13\pm 0.41$ $0.77\pm 0.04$ $1.68\pm 0.31$ $-14.80\pm 0.23$ $0.78\pm 0.17$ $5.65\pm 1.58$ $-7.85\pm 0.29$ $1.10\pm 0.66$ $2.78\pm 0.61$ AEVB-IAF $-4.19\pm 0.05$ ${0.77\pm 0.00}$ $1.64\pm 0.73$ $-5.10\pm 0.30$ $0.61\pm 0.15$ $4.43\pm 1.09$ $-3.97\pm 0.22$ $0.75\pm 0.12$ $1.68\pm 0.27$ AAEB $N/A$ $0.68\pm 0.02$ $1.54\pm 0.19$ $N/A$ $\mathbf{0.52\pm 0.03}$ ${3.34\pm 0.16}$ $N/A$ $0.80\pm 0.02$ $2.46\pm 0.46$ H-AEVB $-7.03\pm 3.13$ ${0.74\pm 0.02}$ $2.25\pm 3.02$ $-7.21\pm 4.62$ ${0.70\pm 0.22}$ $4.04\pm 4.62$ $-5.71\pm 3.05$ ${0.76\pm 0.21}$ ${2.22\pm 2.03}$ PIS $-7.83\pm 0.64$ $0.75\pm 0.14$ $6.50\pm 1.11$ $-9.83\pm 0.61$ $0.61\pm 0.03$ $\mathbf{2.40\pm 1.01}$ $-7.06\pm 0.06$ $0.77\pm 0.04$ $3.67\pm 0.08$ DDVI $\mathbf{-3.88\pm 0.96}$ $\mathbf{0.67\pm 0.04}$ $\mathbf{1.27\pm 0.21}$ $\mathbf{-5.03\pm 0.58}$ $0.62\pm 0.33$ $3.86\pm 0.17$ $\mathbf{-3.79\pm 0.14}$ $\mathbf{0.66\pm 0.07}$ $\mathbf{1.56\pm 0.09}$
 Table 1: Unsupervised learning on MNIST. We report ELBO, MMD between generated images and test images, and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
 5 Experiments
-We compare DDVI with Auto-Encoding Variational Bayes (AEVB) [14], AEVB with inverse autoregressive flow posteriors (AEVB-IAF) [16], Adversarial Auto-Encoding Bayes (AAEB) [20], and Path Integral Sampler (PIS) [46] on MNIST [18] and CIFAR-10 [17] in unsupervised and semi-supervised learning settings, and also on the Thousand Genomes dataset [35].
-We also compare with Hierachical Auto-Encoding Variational Bayes (H-AEVB) [26, 38] in unsupervised setting.
+We compare DDVI with Auto-Encoding Variational Bayes (AEVB) (Kingma and Welling 2013), AEVB with inverse autoregressive flow posteriors (AEVB-IAF) (Kingma et al.
+2016), Adversarial Auto-Encoding Bayes (AAEB) (Makhzani et al.
+2015), and Path Integral Sampler (PIS) (Zhang and Chen 2021) on MNIST (Lecun et al.
+1998) and CIFAR-10 (Krizhevsky and Hinton 2009) in unsupervised and semi-supervised learning settings, and also on the Thousand Genomes dataset (Siva 2008).
+We also compare with Hierachical Auto-Encoding Variational Bayes (H-AEVB) (Ranganath, Tran, and Blei 2016; Vahdat and Kautz 2020) in unsupervised setting.
 We discuss the computational costs of all methods in Appendix D.
-The priors, model architecture, and training details can be founded in Appendix F, Appendix G, and Appendix H respectively.
+The priors, model architecture, and training details can also be founded in Appendix H.
 All results below are reported with 95% confidence interval using 3 different seeds.
@@
 We start with synthetic experiments that are aimed at benchmarking the expressivity of diffusion-based posteriors and their ability to improve fitting $p$, a distribution with a complex structured prior, like one might find in probabilistic programming, scientific analysis, or other applications.
-We fit a model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ on the MNIST and CIFAR-10 datasets with three priors $p(\mathbf{z})$: pinwheel, swiss roll, and square and report our results in Table 1 and Table 6.
-The model distribution $p_{\bm{\theta}}$ is instantiated by a deep Gaussian latent variable model (DGLVM) with multi-layer perceptrons (MLPs) on MNIST and convolutional neural networks (CNNs) on CIFAR-10.
-The details of model architecture are provided in Appendix G.
-Our first set of metrics (ELBO and MMD) seeks to evaluate the learned generative model $p_{\bm{\theta}}$ is good.
+We fit a model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ on the MNIST and CIFAR-10 datasets with three priors $p(\mathbf{z})$: pinwheel, swiss roll, and square and report our results in Table 1 and 7.
+The model distribution $p_{\boldsymbol{\theta}}$ is instantiated by a deep Gaussian latent variable model (DGLVM) with multi-layer perceptrons (MLPs) on MNIST and convolutional neural networks (CNNs) on CIFAR-10 (see the details of model architecture in Appendix G).
+Our first set of metrics (ELBO and MMD) seeks to evaluate the learned generative model $p_{\boldsymbol{\theta}}$ is good.
 In the ELBO calculation, we average the reconstruction loss across image pixels.
-We use MMD to measure sample quality: we generate images with the trained model and calculate MMD between the generated images and test images using a mixture of Gaussian kernel111with sigma equal to [2, 5, 10, 20, 40, 80].
+We use MMD to measure sample quality: we generate images with the trained model and calculate MMD between the generated images and test images using a mixture of Gaussian kernel with sigma equal to $[2,5,10,20,40,80]$.
 We only report MMD for MNIST, since CIFAR-10 generated samples are very low-quality for all methods because the latent dimension is 2.
@@
 We measure latent negative log-likelihood (Latent NLL) by fitting a kernel density estimator (KDE) on the latents produced by the model with test data as input and compute the log-likelihood of the latents sampled from the prior under the fitted KDE.
-From Tables 1 and 6, we see our method DDVI achieve best ELBO in all but one scenario, in which it still performs competitively.
-We also see strong results in Latent NLL and Acc in many scenarios, except for Swiss Roll where AAEB does well.
-We present visualizations of MNIST using the baseline methods and our method in Figure 2.
+From Table 1 and Table 7 in Appendix, we see our method DDVI achieve best ELBO in all but one scenario, in which it still performs competitively.
+We also see strong results in Latent NLL and k-nearest neighbors classification accuracy of the latents (Acc).
+in many scenarios, except for swiss roll where AAEB does well.
+We present visualizations on MNIST using the baselines and our method in Figure 2.
 Figure 3: Semi-supervised visualization on MNIST with 1,000 labels using three different priors (pinwheel, swiss roll, and square). Each a indicates one class.
-Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-11.15\pm 0.53$ $0.93\pm 0.01$ $1.36\pm 0.03$ $-15.29\pm 1.33$ $0.68\pm 0.01$ $4.60\pm 0.23$ $-10.26\pm 0.25$ $0.86\pm 0.01$ $1.68\pm 0.02$ AEVB-IAF $-2.10\pm 0.26$ $\bm{0.95\pm 0.00}$ $\bm{1.06\pm 0.03}$ $-5.38\pm 1.78$ $0.90\pm 0.02$ $2.75\pm 0.14$ $-2.67\pm 0.83$ $\bm{0.91\pm 0.01}$ $\bm{0.90\pm 0.02}$ AAEB $N/A$ $0.89\pm 0.01$ $1.55\pm 0.01$ $N/A$ $0.88\pm 0.01$ $3.07\pm 0.05$ $N/A$ $1.94\pm 0.38$ $0.76\pm 0.13$ DDVI $\bm{-0.24\pm 0.13}$ $\bm{0.95\pm 0.00}$ $\bm{1.06\pm 0.01}$ $\bm{-2.89\pm 0.33}$ $\bm{0.92\pm 0.01}$ $\bm{2.09\pm 0.00}$ $\bm{0.02\pm 0.09}$ $0.90\pm 0.01$ $1.49\pm 0.03$
+Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-11.15\pm 0.53$ $0.93\pm 0.01$ $1.36\pm 0.03$ $-15.29\pm 1.33$ $0.68\pm 0.01$ $4.60\pm 0.23$ $-10.26\pm 0.25$ $0.86\pm 0.01$ $1.68\pm 0.02$ AEVB-IAF $-2.10\pm 0.26$ $\mathbf{0.95\pm 0.00}$ $\mathbf{1.06\pm 0.03}$ $-5.38\pm 1.78$ $0.90\pm 0.02$ $2.75\pm 0.14$ $-2.67\pm 0.83$ $\mathbf{0.91\pm 0.01}$ $\mathbf{0.90\pm 0.02}$ AAEB $N/A$ $0.89\pm 0.01$ $1.55\pm 0.01$ $N/A$ $0.88\pm 0.01$ $3.07\pm 0.05$ $N/A$ $1.94\pm 0.38$ $0.76\pm 0.13$ DDVI $\mathbf{-0.24\pm 0.13}$ $\mathbf{0.95\pm 0.00}$ $\mathbf{1.06\pm 0.01}$ $\mathbf{-2.89\pm 0.33}$ $\mathbf{0.92\pm 0.01}$ $\mathbf{2.09\pm 0.00}$ $\mathbf{0.02\pm 0.09}$ $0.90\pm 0.01$ $1.49\pm 0.03$
 Table 2: Semi-supervised learning on MNIST (1,000 labels).
@@
 We also evaluate the performance of our method and the baselines under semi-supervised learning setting where some labels are observed (1,000 for MNIST and 10,000 for CIFAR-10) and the partitions of the priors are known.
-For this setting, we evaluate ELBO, latent negative log-likelihood (Latent NLL), and a k-nearest neighbors classification accuracy of the latents (Acc).
+For this setting, we evaluate ELBO, Latent NLL, and Acc.
 We choose classification accuracy since classification is a common downstream task for semi-supervised learning.
 We use the same set of priors and baselines.
-Details on how we partition each prior into $p_{\bm{{\bm{\theta}}}}(\mathbf{z}|\mathbf{x},l)$ can be founded in Appendix F.
+Details on how we partition each prior into $p_{\boldsymbol{{\boldsymbol{\theta}}}}(\mathbf{z}|\mathbf{x},l)$ can be founded in Appendix F.
 The partitions defined for our priors are local parts of the priors.
-We note that unlike unsupervised learning, we use the simplified sleep term from Appendix B in our objective for this setting, since $q_{\bm{\phi}}$ already gets extra information from $l$ here.
-The results are shown in Tables 2 and 7.
+We note that unlike unsupervised learning, we use the simplified sleep term in our objective for this setting (see Appendix B for details), since $q_{\boldsymbol{\phi}}$ already gets extra information from $l$ here.
+The results are shown in Table 2 and Table 8 in Appendix.
 DDVI mostly outperforms the baselines across different priors and metrics, especially on CIFAR-10 where DDVI is best across the board.
@@
 We also show the visualizations of the latents in Figure 3 where DDVI matches the prior almost perfectly.
+Method Cluster Purity Cluster Completeness NMI
+AEVB $0.28\pm 0.02$ $\mathbf{0.78\pm 0.16}$ $0.59\pm 0.08$
+AEVB-IAF $0.29\pm 0.04$ $\mathbf{0.73\pm 0.06}$ $0.55\pm 0.06$
+AAEB $0.37\pm 0.06$ $\mathbf{0.76\pm 0.11}$ ${0.63\pm 0.02}$
+DDVI $\mathbf{0.45\pm 0.03}$ $\mathbf{0.75\pm 0.05}$ $\mathbf{0.66\pm 0.04}$
+Table 3: Quantitative genotype clustering results.
 5.3 Clustering and Visualization for Genotype Analysis
@@
 The overall clustering performance can be captured with NMI.
-Method Cluster Cluster NMI
-Method Purity Completeness
-AEVB $0.28\pm 0.02$ $\bm{0.78\pm 0.16}$ $0.59\pm 0.08$
-AEVB-IAF $0.29\pm 0.04$ $\bm{0.73\pm 0.06}$ $0.55\pm 0.06$
-AAEB $0.37\pm 0.06$ $\bm{0.76\pm 0.11}$ ${0.63\pm 0.02}$
-DDVI $\bm{0.45\pm 0.03}$ $\bm{0.75\pm 0.05}$ $\bm{0.66\pm 0.04}$
-Figure 5: Quantitative genotype clustering results.
-In Figure 5, we see that DDVI attains the best performance on cluster purity and NMI.
+In Table 3, we see that DDVI attains the best performance on cluster purity and NMI.
 For cluster completeness, VAE and AAE have better means but much larger confidence interval.
-Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA [44], t-SNE [40], and UMAP [23]).
-The legend of Figure 4 can be founded at Figure 6.
+Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA (Wold, Esbensen, and Geladi 1987), t-SNE (Van der Maaten and Hinton 2008), and UMAP (McInnes, Healy, and Melville 2018)).
+The legend of Figure 4 can be founded at Figure 5 in Appendix.
 6 Discussion
 Diffusion vs. Normalizing Flows
-Our approach is most similar to flow-based approximators [28, 16]; in fact when $T\to\infty$, our diffusion-based posterior effectively becomes a continuous-time normalizing flow [36].
+Our approach is most similar to flow-based approximators (Rezende and Mohamed 2015; Kingma et al.
+2016); in fact when $T\to\infty$, our diffusion-based posterior effectively becomes a continuous-time normalizing flow (Song, Meng, and Ermon 2020).
 However, classical flow-based methods require invertible architectures for each flow layer: this constrains their expressivity and requires backpropagating through potentially a very deep network.
@@
 Diffusion vs. Other Generative Models
-Variational posteriors based on GANs [20] also admit expressive architectures and require only sample-based access to the prior $p(\mathbf{z})$.
+Variational posteriors based on GANs (Makhzani et al.
+2015) also admit expressive architectures and require only sample-based access to the prior $p(\mathbf{z})$.
 Our diffusion-based approach admits a more stable loss, and is potentially more expressive, as it effectively supports an infnite number of layers (with shared parameters when $T\to\infty$).
 Unlike GANs, our models also admit explicit likelihoods and allow us to compute the ELBO for model evaluation.
-Our approach is similar to variational MCMC [31]; however, we train with a better objective augmented with a diffusion loss, and we adopt improved architectures with shared weights across layers.
+Our approach is similar to variational MCMC (Salimans, Kingma, and Welling 2015); however, we train with a better objective augmented with a diffusion loss, and we adopt improved architectures with shared weights across layers.
 Diffusion for Approximate Inference
@@
 Our algorithms are also simpler.
-For example, diffusion sampling methods require backpropagating through a sampling process to minimize the reverse $\text{KL}(q_{\bm{\phi}}||p_{\bm{\theta}})$, which poses challenges with optimization and credit assignment.
-Some methods based on Schrodinger bridges require an iterative optimization process generalizing the sinkhorn algorithm or computationally expensive on-policy or off-policy [21] trajectory-based optimization.
-In contrast, DDVI optimizes the forward $\text{KL}(p_{\bm{\theta}}||q_{\bm{\phi}})$ using simple gradient-based optimization that directly emulates diffusion-based training.
+For example, diffusion sampling methods require backpropagating through a sampling process to minimize the reverse $\text{KL}(q_{\boldsymbol{\phi}}||p_{\boldsymbol{\theta}})$, which poses challenges with optimization and credit assignment.
+Some methods based on Schrodinger bridges require an iterative optimization process generalizing the sinkhorn algorithm or computationally expensive on-policy or off-policy (Malkin et al.
+2022) trajectory-based optimization.
+In contrast, DDVI optimizes the forward $\text{KL}(p_{\boldsymbol{\theta}}||q_{\boldsymbol{\phi}})$ using simple gradient-based optimization that directly emulates diffusion-based training.
 7 Related Work
 Latent Diffusion
-Vahdat et al.
-[39], Wehenkel and Louppe [43], Rombach et al.
-[30] perform diffusion in the latent space of a VAE to improve the efficiency of image generation.
+Vahdat, Kreis, and Kautz (2021); Wehenkel and Louppe (2021); Rombach et al.
+(2022) perform diffusion in the latent space of a VAE.
 Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors.
-Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
-Recent work [25, 47, 42] has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$.
+Our goal is different: we seek to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
+Recent work (Preechakul et al.
+2022; Zhang, Zhao, and Lin 2022; Wang et al.
+2023) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$.
 Cohen et al.
-[5] crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
+(2022) crafts a diffusion bridge linking a continuously coded vector to a non-informative prior distribution.
 Diffusion for Approximate Inference
-Diffusion sampling [3, 46, 41, 45, 29, 32, 1] mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function.
-These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) [46], denoising diffusion sampler (DDS) [41], and time-reversed diffusion sampler (DIS) [3], which have been unified by Richter et al.
-[29].
-Some other works [45, 1] use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration.
-Sendera et al.
-[32] benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps.
+Diffusion sampling (Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Zhang et al.
+2023; Richter, Berner, and Liu 2023; Sendera et al.
+2024; Akhound-Sadegh et al.
+2024) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function.
+These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift).
+Some other works (Zhang et al.
+2023; Akhound-Sadegh et al.
+2024; Sendera et al.
+2024) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration.
+While this work is applicable to our setting, it does not rely on the structure of $p(\mathbf{x},\mathbf{z})$ available to us, namely tractable sampling in $p$.
 Dimensionality Reduction
-Latent variable models in general are an attractive alternative to visualization methods like PCA, UMAP, and t-SNE [23, 40].
+Latent variable models in general are an attractive alternative to visualization methods like PCA, Sparse PCA, NMF, UMAP, and t-SNE (Wold, Esbensen, and Geladi 1987; Kuleshov 2013; Lee and Seung 2000; McInnes, Healy, and Melville 2018; Van der Maaten and Hinton 2008).
 Domain-specific knowledge can be injected through the prior, and deep neural networks can be utilized to achieve a more expressive mapping from the data space to the latent space.
@@
 While this paper focuses on applications of DDVI to dimensionality reduction and visualization, there exist other tasks for the algorithm, e.g., density estimation or sample quality.
-Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making [24, 6], meta-learning [27], or causal effect estimation [7].
+Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making (Nguyen and Grover 2022; Deshpande and Kuleshov 2023), meta-learning (Rastogi et al.
+2023), or causal effect estimation (Deshpande et al.
+2022).
 Since our learning objective differs from the ELBO (it adds a regularizer), we anticipate gains on models whose training benefits from regularization, but perhaps not on all models.
@@
 However, our ability to generate diverse samples and achieve class separation in latent space hints at the method’s potential on these tasks.
+Appendix
 Appendix A Pseudocode
-Here we provide a pseudocode to illustrate the training process of DDVI. We release anonymized code repo at https://anonymous.4open.science/r/vi-with-diffusion-D70F.
+Here we provide a pseudocode to illustrate the training process of DDVI.
 Algorithm 1 DDVI Pseudocode
-1:(Optional) Pre-train $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})$ and $q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})$ with DDVI but with unconditional diffusion model $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$
+1:(Optional) Pre-train $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})$ with DDVI but with unconditional diffusion model $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$
 2:for epoch $=1,\ldots,n$ do
 3:     for $\mathbf{x}_{1},\ldots,\mathbf{x}_{k}\sim p_{\mathcal{D}}(\mathbf{x})$ do
-4:         $\mathbf{y}_{i}\sim q_{\bm{\phi}}(\mathbf{y}|\mathbf{x}_{i})$ and $\mathbf{z}_{i}\sim q(\mathbf{z}|\mathbf{y}_{i},\mathbf{x}_{i})$ for $i=1,\ldots,k$
-5:         Optimize ${\bm{\theta}},{\bm{\phi}}$ with respect to a Monte Carlo estimate of $\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\left(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\|p_{\bm{\theta}}(\mathbf{y},\mathbf{z})\right)$ for each $\mathbf{x}_{i}$triangleright Standard ELBO training part
+4:         $\mathbf{y}_{i}\sim q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x}_{i})$ and $\mathbf{z}_{i}\sim q(\mathbf{z}|\mathbf{y}_{i},\mathbf{x}_{i})$ for $i=1,\ldots,k$
+5:         Optimize ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ with respect to a Monte Carlo estimate of $\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\|p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z})\right)$ for each $\mathbf{x}_{i}$triangleright Standard ELBO training part
 6:         for iteration $=1,\ldots,m$ do $\triangleright$ Do sleep for $m$ iterations
@@
 9:              $\mathbf{y}_{i}\sim r(\mathbf{y}|\mathbf{z}_{i})$ for $i=1,\ldots,k$triangleright Construct fantasy inputs
-10:              Optimize ${\bm{\phi}}$ using the standard diffusion noise prediction loss on $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{i},\hat{\mathbf{x}}_{i})$
+10:              Optimize ${\boldsymbol{\phi}}$ using the standard diffusion noise prediction loss on $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{i},\hat{\mathbf{x}}_{i})$
 11:         end for
@@
 Appendix B Simplifying Wake-Sleep
-In wake-sleep, sampling $\mathbf{x}$ from $p_{\bm{\theta}}$ to obtain gradients for the sleep term introduces computational overhead.
-To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation ^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ of $\mathcal{L}$:
-^$\hat{\mathcal{L}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ $=\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / reconstr.
-term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))}_{\text{latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$}}.$ (16)
-We have replaced $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ with a latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$, in which $\mathbf{x}$ is given, and we only seek to fit the true reverse noising process $r(\mathbf{z}|\mathbf{y})$ independently of $\mathbf{x}$.
+In wake-sleep, sampling $\mathbf{x}$ from $p_{\boldsymbol{\theta}}$ to obtain gradients for the sleep term introduces computational overhead.
+To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation ^$\hat{\mathcal{L}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ of $\mathcal{L}$:
+^$\hat{\mathcal{L}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ $=\underbrace{\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / reconstr.
+term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}\underbrace{-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}\underbrace{-\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))}_{\text{latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$}}.$ (12)
+We have replaced $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ with a latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$, in which $\mathbf{x}$ is given, and we only seek to fit the true reverse noising process $r(\mathbf{z}|\mathbf{y})$ independently of $\mathbf{x}$.
 We can similarly show that
-$\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$ $=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})}[\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})\geq\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})r(\mathbf{y}|\mathbf{z})}[\log({q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}/r(\mathbf{y}|\mathbf{z}))]+\bar{H}(p_{\bm{\theta}})$ (17)
-$=-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{z})}[D_{\mathrm{KL}}(r(\mathbf{y}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{z},\mathbf{x}))]-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z})||q(\mathbf{z}|\mathbf{x})),$ (18)
-where $\bar{H}(p_{\bm{\theta}})$ is an entropy term constant in ${\bm{\phi}}$.
-Thus, we minimize the forward KL divergence by sampling $\mathbf{z}$, and applying the noising process to get $\mathbf{y}$; the $q_{\bm{\phi}}$ is fit to denoise $\mathbf{z}$ from $\mathbf{y}$ as in Equation 10.
-We optimize our bound on ^$\hat{\mathcal{L}}(x,{\bm{\theta}},{\bm{\phi}})$ end-to-end using minibatch gradient descent over ${\bm{\theta}},{\bm{\phi}}$.
-While the wake term is a reconstruction loss as in wake-sleep, the sleep term generates latent samples $\mathbf{z},\mathbf{y}$ from $r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ (by analogy with $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ in normal wake-sleep); the denoiser $q_{\bm{\phi}}$ is trained to recover $\mathbf{z}$ from $\mathbf{y}$.
+$\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$ $=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{z})}[\log{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\boldsymbol{\theta}})\geq\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{z})r(\mathbf{y}|\mathbf{z})}[\log({q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}/r(\mathbf{y}|\mathbf{z}))]+\bar{H}(p_{\boldsymbol{\theta}})$ (13)
+$=-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{z})}[\text{KL}(r(\mathbf{y}|\mathbf{z})||q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{z},\mathbf{x}))]-\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z})||q(\mathbf{z}|\mathbf{x})),$ (14)
+where $\bar{H}(p_{\boldsymbol{\theta}})$ is an entropy term constant in ${\boldsymbol{\phi}}$.
+Thus, we minimize the forward KL divergence by sampling $\mathbf{z}$, and applying the noising process to get $\mathbf{y}$; the $q_{\boldsymbol{\phi}}$ is fit to denoise $\mathbf{z}$ from $\mathbf{y}$ as in Equation 8.
+We optimize our bound on ^$\hat{\mathcal{L}}(x,{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ end-to-end using minibatch gradient descent over ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$.
+While the wake term is a reconstruction loss as in wake-sleep, the sleep term generates latent samples $\mathbf{z},\mathbf{y}$ from $r(\mathbf{y}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ (by analogy with $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ in normal wake-sleep); the denoiser $q_{\boldsymbol{\phi}}$ is trained to recover $\mathbf{z}$ from $\mathbf{y}$.
 Thus, we perform wake-sleep in latent space, which obviates the need for alternating wake and sleep phases, and allows efficient end-to-end training.
-A limitation of this approximation is that the sleep term does not fit $q_{\bm{\phi}}$ to the true $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x},\mathbf{y})$, and as a consequence $\hat{L}$ is not a tight lower bound on $\log p_{\bm{\theta}}(\mathbf{x})$.
-We may think of $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\bm{\phi}})$ as a regularizer to the ELBO.
+A limitation of this approximation is that the sleep term does not fit $q_{\boldsymbol{\phi}}$ to the true $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x},\mathbf{y})$, and as a consequence $\hat{L}$ is not a tight lower bound on $\log p_{\boldsymbol{\theta}}(\mathbf{x})$.
+We may think of $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$ as a regularizer to the ELBO.
 Appendix C Comparision of Methods
 We provide a comprehensive comparison of different methods in Table 4.
-Vahdat et al.
-[39], Wehenkel and Louppe [43], Rombach et al.
-[30] perform diffusion in the latent space of a VAE to improve the efficiency of image generation.
+Vahdat, Kreis, and Kautz (2021); Wehenkel and Louppe (2021); Rombach et al.
+(2022) perform diffusion in the latent space of a VAE to improve the efficiency of image generation.
 Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors.
 Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
-Recent work [25, 47, 42] has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$.
+Recent work (Preechakul et al.
+2022; Zhang, Zhao, and Lin 2022; Wang et al.
+2023) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$.
 Cohen et al.
-[5] crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
+(2022) crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
 Model Training Objective Approximating Family Sample-based Prior Auxiliary Variable Tasks Simplified Graphical Illustration AEVB ELBO Diagonal Gaussian Density estimation $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ AEVB-IAF ELBO Normalizing flow Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{x}$ AAEB Adversarial training Adversarial generator Visualization $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ H-AEVB-(IAF) ELBO Factorial Normal / Normalizing flow Density estimation / High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ ADGM ELBO Non-Gaussian Density estimation $\mathbf{x}\rightarrow\mathbf{a}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ LDM ELBO Diagonal Gaussian High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ LSGM ELBO & score matching Diagonal Gaussian High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ DDVI ELBO & sleep term Denoising diffusion Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{T}(\mathbf{y})\rightarrow\mathbf{z}_{0}(\mathbf{z})\rightarrow\mathbf{x}$
@@
 Appendix E Connections to Diffusion Samplers
-Diffusion sampling [3, 46, 41, 45, 29, 32, 1] mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function.
-These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) [46], denoising diffusion sampler (DDS) [41], and time-reversed diffusion sampler (DIS) [3], which have been unified by Richter et al.
-[29].
-Some other works [45, 1] use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration.
+Diffusion sampling (Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Zhang et al.
+2023; Richter, Berner, and Liu 2023; Sendera et al.
+2024; Akhound-Sadegh et al.
+2024) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function.
+These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) (Zhang and Chen 2021), denoising diffusion sampler (DDS) (Vargas, Grathwohl, and Doucet 2023), and time-reversed diffusion sampler (DIS) (Berner, Richter, and Ullrich 2022), which have been unified by Richter, Berner, and Liu (2023).
+Some other works (Zhang et al.
+2023; Akhound-Sadegh et al.
+2024) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration.
 Sendera et al.
-[32] benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps.
+(2024) benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps.
 Overall, there are indeed some strong connections between these works and ours:
@@
 We aim to perform more accurate variational inference using an auxiliary variable model augmented by diffusion models to improve generative modeling.
-In our setting, $p_{\bm{\theta}}(z|x)$ is a moving target density, as we jointly learn ${\bm{\theta}}$ with ${\bm{\phi}}$, as opposed to a static target density that diffusion-structured samplers are designed to solve.
-  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy [46] or off-policy [21] trajectory-based optimization.
+In our setting, $p_{\boldsymbol{\theta}}(z|x)$ is a moving target density, as we jointly learn ${\boldsymbol{\theta}}$ with ${\boldsymbol{\phi}}$, as opposed to a static target density that diffusion-structured samplers are designed to solve.
+  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy (Zhang and Chen 2021) or off-policy (Malkin et al.
+2022) trajectory-based optimization.
 Instead, we introduce a wake-sleep optimization algorithm and its simplified version to alleviate the weak learning signal issue and optimize the evidence lower bound in a better way.
-  • In Equation 9, we are minimizing the forward KL divergence $\mathcal{D}_{KL}(p_{\bm{\theta}}||q_{\bm{\phi}})$, where diffusion samplers are minimizing the reverse $\mathcal{D}_{KL}(q_{\bm{\phi}}||p_{\bm{\theta}})$.
+  • In Equation 8, we are minimizing the forward KL divergence $\text{KL}(p_{\boldsymbol{\theta}}||q_{\boldsymbol{\phi}})$, where diffusion samplers are minimizing the reverse $\text{KL}(q_{\boldsymbol{\phi}}||p_{\boldsymbol{\theta}})$.
 We also summarize the connections and differences in the table below.
 Diffusion Samplers GFlowNet-based DDVI (ours)
-[3, 46, 41, 1] [45, 29, 32, 21]
+(Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Akhound-Sadegh et al.
+2024) (Zhang et al.
+2023; Richter, Berner, and Liu 2023; Sendera et al.
+2024; Malkin et al.
+2022)
 Tasks Sampling, density estimation Sampling, density estimation Learning, Sampling, Dimensionality reduction
@@
 Model Family for $q$ Markov chain Markov chain Markov chain
-Objective $D_{KL}(q||p)$ with regularizer Trajectory balance objective $D_{KL}(p||q)$ with ELBO
+Objective $\text{KL}(q||p)$ with regularizer Trajectory balance objective $\text{KL}(p||q)$ with ELBO
 Algorithm Gradient descent (with reference process), importance sampling RL-motivated off-policy optimization (replay buffers, Thompson sampling, etc.) Gradient descent with wake-sleep
@@
 Pinwheel.
-This distribution was used in [13].
+This distribution was used in (Johnson et al.
+2016).
 We define the number of clusters to be 10.
@@
 Swiss Roll.
-This distribution was used in [22].
+This distribution was used in (Marsland 2014).
 For semi-supervised learning experiments, this prior is partitioned into 10 partitions.
@@
 Appendix G Model Architecture
-All methods use the same architecture for encoder $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ and decoder $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})$, excluding the extra parts specific to each method which we describe below, for the same dataset.
+All methods use the same architecture for encoder $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ and decoder $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$, excluding the extra parts specific to each method which we describe below, for the same dataset.
 For MNIST, the encoder and decoder are multi-layer perceptron with two hidden layers, each with 1000 hidden units.
@@
 Appendix H Training Details
-For training, we update the parameters for each batch of inputs by alternating between the ELBO phase (optimizing ${\bm{\theta}}$ and ${\bm{\phi}}$ with respect to the ELBO, i.e., the reconstruction term and the prior matching term) and the sleep phase (optimizing ${\bm{\phi}}$ with respect to the sleep term).
+For training, we update the parameters for each batch of inputs by alternating between the ELBO phase (optimizing ${\boldsymbol{\theta}}$ and ${\boldsymbol{\phi}}$ with respect to the ELBO, i.e., the reconstruction term and the prior matching term) and the sleep phase (optimizing ${\boldsymbol{\phi}}$ with respect to the sleep term).
 We use Adam optimizer and latent size of 2 for all of our experiments.
@@
 The legend of the 1000 Genomes Visualization plot can be found at Figure 5.
-Method Pinwheel Swiss Roll Square ELBO Latent NLL ELBO Latent NLL ELBO Latent NLL AEVB $-12.96\pm 1.81$ $3.26\pm 0.60$ $-12.87\pm 4.55$ $6.25\pm 1.58$ $-7.91\pm 0.11$ $2.91\pm 0.17$ AEVB-IAF $-3.24\pm 0.16$ $1.71\pm 0.84$ $-4.03\pm 0.73$ $5.51\pm 0.51$ $\bm{-2.10\pm 0.31}$ $1.71\pm 0.77$ AAEB $N/A$ $1.70\pm 0.41$ $N/A$ $\bm{3.18\pm 0.22}$ $N/A$ $1.67\pm 0.17$ H-AEVB $-4.42\pm 0.46$ $\bm{1.69\pm 0.17}$ $-5.36\pm 0.77$ $5.74\pm 0.55$ $-2.86\pm 0.11$ $1.64\pm 0.09$ PIS ${-2.92\pm 1.23}$ ${3.61\pm 0.62}$ $-4.14\pm 0.49$ $7.14\pm 0.14$ $-4.85\pm 0.06$ $3.91\pm 0.06$ DDVI $\bm{-1.38\pm 0.44}$ ${1.75\pm 0.53}$ $\bm{-3.05\pm 0.65}$ ${5.66\pm 2.63}$ $-2.47\pm 0.30$ $\bm{1.58\pm 0.09}$
+Method Pinwheel Swiss Roll Square
+ELBO Latent NLL ELBO Latent NLL ELBO Latent NLL
+AEVB $-12.96\pm 1.81$ $3.26\pm 0.60$ $-12.87\pm 4.55$ $6.25\pm 1.58$ $-7.91\pm 0.11$ $2.91\pm 0.17$
+AEVB-IAF $-3.24\pm 0.16$ $1.71\pm 0.84$ $-4.03\pm 0.73$ $5.51\pm 0.51$ $\mathbf{-2.10\pm 0.31}$ $1.71\pm 0.77$
+AAEB $N/A$ $1.70\pm 0.41$ $N/A$ $\mathbf{3.18\pm 0.22}$ $N/A$ $1.67\pm 0.17$
+H-AEVB $-4.42\pm 0.46$ $\mathbf{1.69\pm 0.17}$ $-5.36\pm 0.77$ $5.74\pm 0.55$ $-2.86\pm 0.11$ $1.64\pm 0.09$
+PIS ${-2.92\pm 1.23}$ ${3.61\pm 0.62}$ $-4.14\pm 0.49$ $7.14\pm 0.14$ $-4.85\pm 0.06$ $3.91\pm 0.06$
+DDVI $\mathbf{-1.38\pm 0.44}$ ${1.75\pm 0.53}$ $\mathbf{-3.05\pm 0.65}$ ${5.66\pm 2.63}$ $-2.47\pm 0.30$ $\mathbf{1.58\pm 0.09}$
 Table 7: Unsupervised learning on CIFAR-10. We report ELBO and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
-Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $-17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $-13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $-5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $-5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $-4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\bm{-1.60\pm 0.29}$ $\bm{0.49\pm 0.01}$ $\bm{1.09\pm 0.05}$ $\bm{-4.13\pm 1.51}$ $\bm{0.47\pm 0.09}$ $\bm{2.29\pm 0.08}$ $\bm{-1.73\pm 0.64}$ $\bm{0.49\pm 0.01}$ $\bm{1.48\pm 0.02}$
+Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $-17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $-13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $-5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $-5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $-4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\mathbf{-1.60\pm 0.29}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.09\pm 0.05}$ $\mathbf{-4.13\pm 1.51}$ $\mathbf{0.47\pm 0.09}$ $\mathbf{2.29\pm 0.08}$ $\mathbf{-1.73\pm 0.64}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.48\pm 0.02}$
 Table 8: Semi-supervised learning on CIFAR-10 (10,000 labels).
 We report ELBO, accuracy using KNN (K=20) classifier (Acc), and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
-Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB -- -- -- H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\bm{1.27\pm 0.21}$ $\bm{3.86\pm 1.17}$ $\bm{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
+Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB -- -- -- H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\mathbf{1.27\pm 0.21}$ $\mathbf{3.86\pm 1.17}$ $\mathbf{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
 Table 9: Unsupervised learning on MNIST, including the results of DDVI without the sleep term.
-Appendix J Data Licenses
-Yann LeCun and Corinna Cortes hold the copyright of MNIST dataset, which is a derivative work from original NIST datasets.
-CIFAR-10 is under MIT license.
-1000 Genomes dataset is under the creative commons Attribution-NonCommercial-ShareAlike 3.0 Unported licence.
+Appendix J ELBO for Auxiliary-Variable Generative Models
+We aim to derive the lower bound for the log-likelihood $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ by introducing auxiliary variables and applying the Evidence Lower Bound (ELBO) twice.
+Introduce the latent variable $\mathbf{z}$ and apply the ELBO:
+[…]
+[…]
+$=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\text{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\right)$ (17)
+Then, introduce an auxiliary variable $\mathbf{y}$ and apply the ELBO again:
+$\log p_{\boldsymbol{\theta}}(\mathbf{x})$ $\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\text{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\right)$ (19)
+$=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})+\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\log p(\mathbf{z})$ (20)
+$\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]\right]-\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})}\left[\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{x},\mathbf{z})}\right]\right]+\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})}\left[\log p(\mathbf{z})\right]\right]$ (21)
+$=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\text{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\,\|\,r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z})\right)$ (22)
+Appendix K Diffusion Regularization
+We begin with the definition of $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, derived as a lower bound on $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$:
+$\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}\left[\mathbb{E}_{r}\left[\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{z},\mathbf{x})}\right]\right]+\bar{H}(p_{\boldsymbol{\theta}}).$
+To simplify $\mathcal{L}_{\text{diff}}$, we leverage the Markov structure of the forward process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ and the reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$.
+The forward process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ is decomposed as:
+$r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x}),$
+where $\mathbf{y}_{0}=\mathbf{z}$.
+The reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ is decomposed as:
+$q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}),$
+where $\mathbf{y}_{0}=\mathbf{z}$ and $\mathbf{y}_{1:T}$ are increasingly noisy versions of $\mathbf{z}$.
+Substituting the factorizations of $r$ and $q_{\boldsymbol{\phi}}$ into the definition of $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, and rewriting the logarithm and rearranging terms, we have:
+$\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ $=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}\Bigg{[}\mathbb{E}_{r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})}\Bigg{[}\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})}{\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x})}\Bigg{]}\Bigg{]}+\bar{H}(p_{\boldsymbol{\theta}})$
+$=\mathbb{E}_{r(\mathbf{y}_{1:T},\mathbf{z},\mathbf{x})}\left[\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}\text{KL}(r_{t}||q_{t})\right]-\text{KL}(r(\mathbf{y}_{T}|\mathbf{z},\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x}))+\bar{H}(p_{\boldsymbol{\theta}})$
+Here, $\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})$ corresponds to the reconstruction term for the initial latent state $\mathbf{z}$, while the summation represents the KL divergence between the forward and reverse processes at each intermediate step.
+$r_{t}=r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{z},\mathbf{x})$ is the conditional forward distribution at step $t$, $q_{t}=q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ is the reverse process distribution at step $t$.
+$\bar{H}(p_{\boldsymbol{\theta}})$ is the expected conditional entropy of $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on ${\boldsymbol{\phi}}$.
+This form mirrors the ELBO derivation for diffusion models (Sohl-Dickstein et al.
+2015), where each step in the Markov chain contributes a KL divergence term, and the reconstruction term arises from the connection between the noisy latent variables and the original data.

### v5

@@
 in which we approximate the marginal likelihood $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ of a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ with an evidence lower bound $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ that is a function of an approximate variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$.
-We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks, topic models (Blei, Ng, and Jordan 2003), variational autoencoders (VAEs), and broad classes of $p_{\boldsymbol{\theta}}$ defined via modern probabilistic programming frameworks (Gordon et al.
-2014).
+We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks, topic models (blei2003latent), variational autoencoders (VAEs), and broad classes of $p_{\boldsymbol{\theta}}$ defined via modern probabilistic programming frameworks (gordon2014probabilistic).
 Maximizing $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ over $\boldsymbol{\phi}$} yields a variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that approximates $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ as well as a tight bound on $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ that serves as a learning objective for $p_{\boldsymbol{\theta}}$.
 The approximation gap $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})-\max_{\boldsymbol{\phi}}\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ equals precisely $\min_{\boldsymbol{\phi}}\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$, which motivates the design of expressive classes of posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that reduce this gap.
-Recent efforts leverage modern generative models—including normalizing flows (Rezende and Mohamed 2015; Kingma et al.
-2016) and generative adversarial networks (Goodfellow et al.
-2014; Makhzani et al.
-2015)—as expressive model families for $q_{\boldsymbol{\phi}}$ that tighten the ELBO.
-This work seeks to further improve variational inference via expressive posteriors based on diffusion models (Ho, Jain, and Abbeel 2020; Song et al.
-2020).
-Diffusion methods have become the de-facto standard for high-quality image synthesis (Rombach et al.
-2022; Gokaslan et al.
-2024).
+Recent efforts leverage modern generative models—including normalizing flows (rezende2015variational; kingma2016improved) and generative adversarial networks (goodfellow2014generative; makhzani2015adversarial)—as expressive model families for $q_{\boldsymbol{\phi}}$ that tighten the ELBO.
+This work seeks to further improve variational inference via expressive posteriors based on diffusion models (ho2020denoising; song2020score).
+Diffusion methods have become the de-facto standard for high-quality image synthesis (rombach2022high; gokaslan2024commoncanvas).
 Here, we use diffusion in latent space to parameterize $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$.
-We train this distribution with a denoising diffusion-like objective that does not involve adversarial training (Makhzani et al.
-2015) or constrained invertible normalizing flow architectures (Kingma et al.
-2016).
+We train this distribution with a denoising diffusion-like objective that does not involve adversarial training (makhzani2015adversarial) or constrained invertible normalizing flow architectures (kingma2016improved).
 Samples from $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ are obtained via iterative refinement of $\mathbf{z}$, starting from a Gaussian distribution, and gradually forming one that is multi-modal and complex.
-[…]
-Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion and a learning objective inspired by the wake-sleep algorithm (Hinton et al.
-1995) that implements regularized variational inference.
+Our work expands upon existing diffusion-based approximate inference methods (berner2022optimal; zhang2021path; vargas2023denoising; zhang2023diffusion; richter2023improved; sendera2024diffusion; akhound2024iterated) that focus on the task of drawing samples from unnormalized distributions $\tilde{p}(\mathbf{z})$ and estimating the partition function $Z=\int_{\mathbf{z}}\tilde{p}(\mathbf{z})d\mathbf{z}$.
+While these methods are applicable in our setting—we set the unnormalized $\tilde{p}(\mathbf{z})$ to $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ such that $Z=p_{\boldsymbol{\theta}}(\mathbf{x})$—they do not make use of characteristics of $p_{\boldsymbol{\theta}}$ that are common in many types of models (VAEs, Bayes networks, etc.), namely the factorization $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and efficient sampling.
+We find that leveraging these properties yields simpler algorithms that avoid backpropagating through a sampling process, and that are fast enough to perform learning in addition to inference.
+Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion and a learning objective inspired by the wake-sleep algorithm (hinton1995wake) that implements regularized variational inference.
 We also derive extensions of our method to semi-supervised learning and clustering.
@@
 We evaluate DDVI on synthetic benchmarks and on a real problem in biological data analysis—inferring human ancestry from genetic data.
-Our method outperforms strong baselines on 1000 Genomes dataset (Siva 2008) and learns a low-dimensional latent space that preserves biologically meaningful structure (Haghverdi, Buettner, and Theis 2015).
+Our method outperforms strong baselines on 1000 Genomes dataset (siva20081000) and learns a low-dimensional latent space that preserves biologically meaningful structure (haghverdi2015diffusion).
 Contributions.
@@
 which serves as a tractable surrogate for the marginal log-likelihood (MLL).
-The gap between the MLL and the ELBO equals precisely $\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO (Kingma and Welling 2013).
-Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods (Maaløe et al.
-2016), MCMC-based methods (Salimans, Kingma, and Welling 2015), normalizing flows (Rezende and Mohamed 2015)—or improved learning objectives—e.g., adversarial or sample-based losses (Makhzani et al.
-2015; Zhao, Song, and Ermon 2017; Si, Bishop, and Kuleshov 2022; Si et al.
-2023).
-The wake-sleep algorithm (Hinton et al. 1995) optimizes an alternative objective
+The gap between the MLL and the ELBO equals precisely $\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO (kingma2013auto).
+Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods (maaloe2016auxiliary), MCMC-based methods (salimans2015mcmc), normalizing flows (rezende2015variational)—or improved learning objectives—e.g., adversarial or sample-based losses (makhzani2015adversarial; zhao2017infovae; si2022autoregressive; pmlr-v202-si23a).
+The wake-sleep algorithm (hinton1995wake) optimizes an alternative objective
 $\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]-\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})),$
@@
 We obtain $\mathbf{y}_{1:T}$ by applying a Markov chain $q(\mathbf{y}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where we define $\mathbf{y}_{0}=\mathbf{x}_{0}$ for convenience.
-When $\mathbf{x}_{0}$ is a continuous vector, a standard choice of transition kernel is $q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{I})$, which is a Gaussian centered around a copy of $\mathbf{y}_{t-1}$ to which we added noise following a schedule $10<{\boldsymbol{\alpha}}_{1}<{\boldsymbol{\alpha}}_{2}<...<{\boldsymbol{\alpha}}_{T}=$ 1.
+When $\mathbf{x}_{0}$ is a continuous vector, a standard choice of transition kernel is $q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{I})$, which is a Gaussian centered around a copy of $\mathbf{y}_{t-1}$ to which we added noise following a schedule $0<{\boldsymbol{\alpha}}_{1}<{\boldsymbol{\alpha}}_{2}<...<{\boldsymbol{\alpha}}_{T}=1$.
 A diffusion model can then be represented as a latent variable distribution $p(\mathbf{x}_{0},\mathbf{y}_{1:T})$ that factorizes as $p(\mathbf{x}_{0},\mathbf{y}_{1:T})=p(\mathbf{y}_{T})\prod_{t=0}^{T-1}p_{\boldsymbol{\theta}}(\mathbf{y}_{t}\mid\mathbf{y}_{t+1})$ (again using $\mathbf{y}_{0}$ as shorthand for $\mathbf{x}_{0}$).
@@
 The goal of DDVI is to fit a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$.
-We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) (Kingma and Welling 2013).
+We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) (kingma2013auto).
 Our approach is comprised of three components:
@@
 3.1 Modeling Family: Diffusion-Based Posteriors
-[…]
+DDVI performs variational inference using a family of approximate posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})=\int_{\mathbf{y}}q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{x})q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})d\mathbf{y}$, which themselves contain latent variables $\mathbf{y}\in\mathcal{Y}$.
+The models $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y},\mathbf{x}),q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})$ must have tractable densities and support gradient-based optimization over $\boldsymbol{\phi}$}.
 We choose the latent $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ to be a vector of $T$ variables that represent progressively simplified versions of $\mathbf{z}$, with $\mathbf{y}_{T}$ corresponding to a simple distribution (e.g., a Gaussian).
@@
 We define the relationship between $\mathbf{y}$ and $\mathbf{z}$ via a forward diffusion process $r(\mathbf{y}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1}|\mathbf{z},\mathbf{x})\prod_{t=1}^{T-1}r(\mathbf{y}_{t+1}|\mathbf{y}_{t},\mathbf{x})$, which transforms $\mathbf{z}$—the latent whose intractable posterior we seek to approximate—into $\mathbf{y}_{T}$, whose posterior is easier to model (possibly conditioned on $\mathbf{x}$).
-Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes (Austin et al.
-2021).
+Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes (austin2021structured).
 The model $q_{\boldsymbol{\phi}}$ is trained to approximately reverse this forward diffusion process.
 3.2 Learning Objective: A Markovian ELBO
-The standard approach to fit auxiliary-variable generative models (Maaløe et al. 2016) is to apply the ELBO twice:
+The standard approach to fit auxiliary-variable generative models (maaloe2016auxiliary) is to apply the ELBO twice:
 $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ $\geq\log p_{\boldsymbol{\theta}}(\mathbf{x})-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$ (3)
@@
 Notice that optimizing $\mathcal{L}_{\text{reg}}$ involves sampling from the approximate reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ to match the true reverse process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$: this is the opposite of diffusion training, where we would sample from $r$ to fit $q$.
-This type of on-policy learning of $q$ has been studied in the context of approximate inference (Zhang and Chen 2021); however, it requires backpropagating through $T$ samples, which may hamper training, and it optimizes a mode-covering divergence that may struggle to fit complex $p(\mathbf{z})$.
+This type of on-policy learning of $q$ has been studied in the context of approximate inference (zhang2021path); however, it requires backpropagating through $T$ samples, which may hamper training, and it optimizes a mode-covering divergence that may struggle to fit complex $p(\mathbf{z})$.
 Adding Wake-Sleep Regularization to the ELBO
@@
 We may use the Markov structure in $q,r$ to rewrite $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ as a sum of $T$ terms, one per Markov step.
-The derivation is identical to that used to obtain the ELBO of a diffusion model (Sohl-Dickstein et al.
-2015), and yields an expression of the same form:
+The derivation is identical to that used to obtain the ELBO of a diffusion model (sohl2015deep), and yields an expression of the same form:
 $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})=$ $~\mathbb{E}_{r}\left[\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}\text{KL}(r_{t}||q_{t})\right]$ (9)
-$-\text{KL}(r(\mathbf{y}_{T}|\mathbf{z})||q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})).$
+$-\text{KL}(r(\mathbf{y}_{T}|\mathbf{z},\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x}))+\bar{H}(p_{\boldsymbol{\theta}}).$
 where $r_{t},q_{t}$ denote the distributions $r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0},\mathbf{x})$ and $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ (see Appendix K for the derivation).
@@
 Simplified Wake-Sleep
-We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$.
+We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z})$ do not depend on $\mathbf{x}$.
 This scenario admits the following optimization procedure:
@@
 Practical Considerations
-A common type of noising process compatible with this bound when $\mathbf{z}$ is continuous is Gaussian diffusion, where we define $r(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},{\boldsymbol{\alpha}}_{t}\mathbf{I})$ for a suitable schedule $({\boldsymbol{\alpha}}_{t})_{t=1}^{T}$.
-We then adopt the parameterization $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})=\mathcal{N}(\mathbf{y}_{t-1};\mu_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x},t),\Sigma_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x},t))$.
-It is then common to parameterize $q_{\boldsymbol{\phi}}$ with a noise prediction network ${\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}$ (Ho, Jain, and Abbeel 2020); the sum of KL divergences can be approximated by $\mathbb{E}_{t,{\boldsymbol{\epsilon}}_{t}\sim r(\mathbf{y}_{0},t)}||{\boldsymbol{\epsilon}}_{t}-{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\sqrt{\bar{{\boldsymbol{\alpha}}_{t}}}\mathbf{y}_{0}+\sqrt{1-\bar{{\boldsymbol{\alpha}}_{t}}}{\boldsymbol{\epsilon}}_{t},\mathbf{x},t)||^{2}$.
-Other extensions include discrete denoising processes (Austin et al.
-2021; Sahoo et al.
-2024; Schiff et al.
-2024).
-In the wake-sleep setting, we know both endpoints $\mathbf{y}_{T}\sim q(\cdot|\mathbf{x})$ and $\mathbf{y}_{0}=\mathbf{z}$ of the diffusion process, opening the possibility for applying optimal transport techniques (Cuturi 2013; De Bortoli et al.
-2021).
+[…]
 4 Extensions
 4.1 Semi-Supervised Learning
-Following Makhzani et al.
-(2015), we extend our algorithm to the semi-supervised learning setting where some data points have labels denoted by $l$.
+Following makhzani2015adversarial, we extend our algorithm to the semi-supervised learning setting where some data points have labels denoted by $l$.
 We assume the user provides a model of the form $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z},l)=p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z},l)r(\mathbf{y}|\mathbf{z},l)p_{\boldsymbol{\theta}}(\mathbf{z}|l)p(l)$; we set the variational distributions to $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y},l),q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x}),q_{\boldsymbol{\phi}}(l|\mathbf{x})$.
-In this setting, we consider two cases, depending on whether the label is observed (Kingma et al.
-2014).
+In this setting, we consider two cases, depending on whether the label is observed (kingma2014semi).
 We extend Equation (7) to incorporate the label $l$ corresponding to a data point as follows:
@@
 When the label $c$ cannot be observed, we treat it as a latent variable and modify the learning objective $\mathcal{U}_{\text{semi}}=\sum_{c}q_{\boldsymbol{\phi}}(l|\mathbf{x})\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{\phi}})+\text{KL}(q_{\boldsymbol{\phi}}(l|\mathbf{x})||p(l))$.
-Therefore, we can conclude a marginal likelihood on our dataset as follows: ~$\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
+Therefore, we can conclude a marginal likelihood on our dataset as follows: $\tilde{\mathcal{L}}_{\text{semi}}=\sum_{(\mathbf{x},l)\in L}\mathcal{L}_{\text{semi}}(\mathbf{x},l,{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}})+\sum_{\mathbf{x}\in U}\mathcal{U}_{\text{semi}}(\mathbf{x},{\boldsymbol{{\boldsymbol{\theta}}}},{\boldsymbol{{\boldsymbol{\phi}}}}).$ where $L$ and $U$ are the sets of data with and without labels, respectively.
 We also want to guarantee that all model parameters can be learned in all cases, including $q_{\boldsymbol{\phi}}(l|\mathbf{x})$, such that this posterior can be applied as a classifier during inference.
-Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: ~$\tilde{\mathcal{L}}_{\text{semi}_{\boldsymbol{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\boldsymbol{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\boldsymbol{\phi}}(l|\mathbf{x})\right]$
+Thus, we combine the marginal likelihood with a classification loss to form an extended learning objective: $\tilde{\mathcal{L}}_{\text{semi}_{\boldsymbol{\alpha}}}=\tilde{\mathcal{L}}_{\text{semi}}+{\boldsymbol{\alpha}}\cdot\mathbb{E}_{\tilde{p}(\mathbf{x},l)}\left[-\log q_{\boldsymbol{\phi}}(l|\mathbf{x})\right]$
 4.2 Clustering
@@
 $-\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x})}\left[\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))\right]$
-Expectations over small numbers of classes $\mathbf{c}$ are done analytically; larger $\mathbf{c}$ require backpropagating through discrete sampling (Jang, Gu, and Poole 2016; Sahoo et al. 2023).
+Expectations over small numbers of classes $\mathbf{c}$ are done analytically; larger $\mathbf{c}$ require backpropagating through discrete sampling (jang2016categorical; sahoo2023backpropagation).
 Figure 2: Unsupervised visualization on MNIST using three priors (pinwheel, swiss roll, and square). Each color indicates a class.
-Method Pinwheel Swiss Roll Square ELBO MMD Latent NLL ELBO MMD Latent NLL ELBO MMD Latent NLL AEVB $-12.13\pm 0.41$ $0.77\pm 0.04$ $1.68\pm 0.31$ $-14.80\pm 0.23$ $0.78\pm 0.17$ $5.65\pm 1.58$ $-7.85\pm 0.29$ $1.10\pm 0.66$ $2.78\pm 0.61$ AEVB-IAF $-4.19\pm 0.05$ ${0.77\pm 0.00}$ $1.64\pm 0.73$ $-5.10\pm 0.30$ $0.61\pm 0.15$ $4.43\pm 1.09$ $-3.97\pm 0.22$ $0.75\pm 0.12$ $1.68\pm 0.27$ AAEB $N/A$ $0.68\pm 0.02$ $1.54\pm 0.19$ $N/A$ $\mathbf{0.52\pm 0.03}$ ${3.34\pm 0.16}$ $N/A$ $0.80\pm 0.02$ $2.46\pm 0.46$ H-AEVB $-7.03\pm 3.13$ ${0.74\pm 0.02}$ $2.25\pm 3.02$ $-7.21\pm 4.62$ ${0.70\pm 0.22}$ $4.04\pm 4.62$ $-5.71\pm 3.05$ ${0.76\pm 0.21}$ ${2.22\pm 2.03}$ PIS $-7.83\pm 0.64$ $0.75\pm 0.14$ $6.50\pm 1.11$ $-9.83\pm 0.61$ $0.61\pm 0.03$ $\mathbf{2.40\pm 1.01}$ $-7.06\pm 0.06$ $0.77\pm 0.04$ $3.67\pm 0.08$ DDVI $\mathbf{-3.88\pm 0.96}$ $\mathbf{0.67\pm 0.04}$ $\mathbf{1.27\pm 0.21}$ $\mathbf{-5.03\pm 0.58}$ $0.62\pm 0.33$ $3.86\pm 0.17$ $\mathbf{-3.79\pm 0.14}$ $\mathbf{0.66\pm 0.07}$ $\mathbf{1.56\pm 0.09}$
+Method Pinwheel Swiss Roll Square ELBO MMD Latent NLL ELBO MMD Latent NLL ELBO MMD Latent NLL AEVB $12.13\pm 0.41$ $0.77\pm 0.04$ $1.68\pm 0.31$ $14.80\pm 0.23$ $0.78\pm 0.17$ $5.65\pm 1.58$ $7.85\pm 0.29$ $1.10\pm 0.66$ $2.78\pm 0.61$ AEVB-IAF $4.19\pm 0.05$ $0.77\pm 0.00$} $1.64\pm 0.73$ $5.10\pm 0.30$ $0.61\pm 0.15$ $4.43\pm 1.09$ $3.97\pm 0.22$ $0.75\pm 0.12$ $1.68\pm 0.27$ AAEB $N/A$ $0.68\pm 0.02$ $1.54\pm 0.19$ $N/A$ $\mathbf{0.52\pm 0.03}$ $3.34\pm 0.16$} $N/A$ $0.80\pm 0.02$ $2.46\pm 0.46$ H-AEVB $7.03\pm 3.13$ $0.74\pm 0.02$} $2.25\pm 3.02$ $7.21\pm 4.62$ $0.70\pm 0.22$} $4.04\pm 4.62$ $5.71\pm 3.05$ $0.76\pm 0.21$} $2.22\pm 2.03$} PIS $7.83\pm 0.64$ $0.75\pm 0.14$ $6.50\pm 1.11$ $9.83\pm 0.61$ $0.61\pm 0.03$ $\mathbf{2.40\pm 1.01}$ $6\pm 0.06 0.77±0$$7\pm 0.04$ $3.67\pm 0.08$ DDVI $\mathbf{-3.88\pm 0.96}$ $\mathbf{0.67\pm 0.04}$ $\mathbf{1.27\pm 0.21}$ $\mathbf{-5.03\pm 0.58}$ $0.62\pm 0.33$ $3.86\pm 0.17$ $\mathbf{-3.79\pm 0.14}$ $\mathbf{0.66\pm 0.07}$ $\mathbf{1.56\pm 0.09}$
 Table 1: Unsupervised learning on MNIST. We report ELBO, MMD between generated images and test images, and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
 5 Experiments
-We compare DDVI with Auto-Encoding Variational Bayes (AEVB) (Kingma and Welling 2013), AEVB with inverse autoregressive flow posteriors (AEVB-IAF) (Kingma et al.
-2016), Adversarial Auto-Encoding Bayes (AAEB) (Makhzani et al.
-2015), and Path Integral Sampler (PIS) (Zhang and Chen 2021) on MNIST (Lecun et al.
-1998) and CIFAR-10 (Krizhevsky and Hinton 2009) in unsupervised and semi-supervised learning settings, and also on the Thousand Genomes dataset (Siva 2008).
-We also compare with Hierachical Auto-Encoding Variational Bayes (H-AEVB) (Ranganath, Tran, and Blei 2016; Vahdat and Kautz 2020) in unsupervised setting.
+We compare DDVI with Auto-Encoding Variational Bayes (AEVB) (kingma2013auto), AEVB with inverse autoregressive flow posteriors (AEVB-IAF) (kingma2016improved), Adversarial Auto-Encoding Bayes (AAEB) (makhzani2015adversarial), and Path Integral Sampler (PIS) (zhang2021path) on MNIST (lecun1998mnist) and CIFAR-10 (krizhevsky2009cifar) in unsupervised and semi-supervised learning settings, and also on the Thousand Genomes dataset (siva20081000).
+We also compare with Hierachical Auto-Encoding Variational Bayes (H-AEVB) (ranganath2016hierarchical; vahdat2020nvae) in unsupervised setting.
 We discuss the computational costs of all methods in Appendix D.
@@
 Figure 3: Semi-supervised visualization on MNIST with 1,000 labels using three different priors (pinwheel, swiss roll, and square). Each a indicates one class.
-Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-11.15\pm 0.53$ $0.93\pm 0.01$ $1.36\pm 0.03$ $-15.29\pm 1.33$ $0.68\pm 0.01$ $4.60\pm 0.23$ $-10.26\pm 0.25$ $0.86\pm 0.01$ $1.68\pm 0.02$ AEVB-IAF $-2.10\pm 0.26$ $\mathbf{0.95\pm 0.00}$ $\mathbf{1.06\pm 0.03}$ $-5.38\pm 1.78$ $0.90\pm 0.02$ $2.75\pm 0.14$ $-2.67\pm 0.83$ $\mathbf{0.91\pm 0.01}$ $\mathbf{0.90\pm 0.02}$ AAEB $N/A$ $0.89\pm 0.01$ $1.55\pm 0.01$ $N/A$ $0.88\pm 0.01$ $3.07\pm 0.05$ $N/A$ $1.94\pm 0.38$ $0.76\pm 0.13$ DDVI $\mathbf{-0.24\pm 0.13}$ $\mathbf{0.95\pm 0.00}$ $\mathbf{1.06\pm 0.01}$ $\mathbf{-2.89\pm 0.33}$ $\mathbf{0.92\pm 0.01}$ $\mathbf{2.09\pm 0.00}$ $\mathbf{0.02\pm 0.09}$ $0.90\pm 0.01$ $1.49\pm 0.03$
+Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $11.15\pm 0.53$ $0.93\pm 0.01$ $1.36\pm 0.03$ $15.29\pm 1.33$ $0.68\pm 0.01$ $4.60\pm 0.23$ $10.26\pm 0.25$ $0.86\pm 0.01$ $1.68\pm 0.02$ AEVB-IAF $2.10\pm 0.26$ $\mathbf{0.95\pm 0.00}$ $\mathbf{1.06\pm 0.03}$ $5.38\pm 1.78$ $0.90\pm 0.02$ $2.75\pm 0.14$ $2.67\pm 0.83$ $\mathbf{0.91\pm 0.01}$ $\mathbf{0.90\pm 0.02}$ AAEB $N/A$ $0.89\pm 0.01$ $1.55\pm 0.01$ $N/A$ $0.88\pm 0.01$ $3.07\pm 0.05$ $N/A$ $1.94\pm 0.38$ $0.76\pm 0.13$ DDVI $\mathbf{-0.24\pm 0.13}$ $\mathbf{0.95\pm 0.00}$ $\mathbf{1.06\pm 0.01}$ $\mathbf{-2.89\pm 0.33}$ $\mathbf{0.92\pm 0.01}$ $\mathbf{2.09\pm 0.00}$ $\mathbf{0.02\pm 0.09}$ $0.90\pm 0.01$ $1.49\pm 0.03$
 Table 2: Semi-supervised learning on MNIST (1,000 labels).
@@
 For cluster completeness, VAE and AAE have better means but much larger confidence interval.
-Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA (Wold, Esbensen, and Geladi 1987), t-SNE (Van der Maaten and Hinton 2008), and UMAP (McInnes, Healy, and Melville 2018)).
+Furthermore, we visualize our genotype clustering results in latent space, shown in Figure 4, and also report results from classical dimensionality reduction and visualization methods that do not perform clustering (PCA (wold1987principal), t-SNE (van2008visualizing), and UMAP (mcinnes2018umap)).
 The legend of Figure 4 can be founded at Figure 5 in Appendix.
@@
 Diffusion vs. Normalizing Flows
-Our approach is most similar to flow-based approximators (Rezende and Mohamed 2015; Kingma et al.
-2016); in fact when $T\to\infty$, our diffusion-based posterior effectively becomes a continuous-time normalizing flow (Song, Meng, and Ermon 2020).
+Our approach is most similar to flow-based approximators (rezende2015variational; kingma2016improved); in fact when $T\to\infty$, our diffusion-based posterior effectively becomes a continuous-time normalizing flow (song2020denoising).
 However, classical flow-based methods require invertible architectures for each flow layer: this constrains their expressivity and requires backpropagating through potentially a very deep network.
@@
 Diffusion vs. Other Generative Models
-Variational posteriors based on GANs (Makhzani et al.
-2015) also admit expressive architectures and require only sample-based access to the prior $p(\mathbf{z})$.
+Variational posteriors based on GANs (makhzani2015adversarial) also admit expressive architectures and require only sample-based access to the prior $p(\mathbf{z})$.
 Our diffusion-based approach admits a more stable loss, and is potentially more expressive, as it effectively supports an infnite number of layers (with shared parameters when $T\to\infty$).
 Unlike GANs, our models also admit explicit likelihoods and allow us to compute the ELBO for model evaluation.
-Our approach is similar to variational MCMC (Salimans, Kingma, and Welling 2015); however, we train with a better objective augmented with a diffusion loss, and we adopt improved architectures with shared weights across layers.
+Our approach is similar to variational MCMC (salimans2015mcmc); however, we train with a better objective augmented with a diffusion loss, and we adopt improved architectures with shared weights across layers.
 Diffusion for Approximate Inference
-[…]
+Existing diffusion-based approximate inference methods (berner2022optimal; zhang2021path; vargas2023denoising; zhang2023diffusion; richter2023improved; sendera2024diffusion; akhound2024iterated) focus on the task of drawing samples from unnormalized distributions $\tilde{p}(\mathbf{z})$ and estimating the partition function $Z=\int_{\mathbf{z}}\tilde{p}(\mathbf{z})d\mathbf{z}$.
+While these methods are applicable in our setting—we set the unnormalized $\tilde{p}(\mathbf{z})$ to $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ such that $Z=p_{\boldsymbol{\theta}}(\mathbf{x})$—they also tackle a more challenging problem (drawing samples from energy-based models) in more general classes of models (arbitrary unnormalized distributions).
+In contrast, we focus on restricted but still important classes of models (VAEs, Bayes networks, etc.), and we solve more challenging sets of tasks (e.g., maximum-likelihood learning) by using properties of $p_{\boldsymbol{\theta}}$ (the factorization $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and efficient sampling from $p_{\boldsymbol{\theta}}$).
 Our algorithms are also simpler.
 For example, diffusion sampling methods require backpropagating through a sampling process to minimize the reverse $\text{KL}(q_{\boldsymbol{\phi}}||p_{\boldsymbol{\theta}})$, which poses challenges with optimization and credit assignment.
-Some methods based on Schrodinger bridges require an iterative optimization process generalizing the sinkhorn algorithm or computationally expensive on-policy or off-policy (Malkin et al.
-2022) trajectory-based optimization.
+Some methods based on Schrodinger bridges require an iterative optimization process generalizing the sinkhorn algorithm or computationally expensive on-policy or off-policy (malkin2022trajectory) trajectory-based optimization.
 In contrast, DDVI optimizes the forward $\text{KL}(p_{\boldsymbol{\theta}}||q_{\boldsymbol{\phi}})$ using simple gradient-based optimization that directly emulates diffusion-based training.
@@
 Latent Diffusion
-Vahdat, Kreis, and Kautz (2021); Wehenkel and Louppe (2021); Rombach et al.
-(2022) perform diffusion in the latent space of a VAE.
+vahdat2021score; wehenkel2021diffusion; rombach2022high perform diffusion in the latent space of a VAE.
 Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors.
 Our goal is different: we seek to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
-Recent work (Preechakul et al.
-2022; Zhang, Zhao, and Lin 2022; Wang et al.
-2023) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$.
-Cohen et al.
-(2022) crafts a diffusion bridge linking a continuously coded vector to a non-informative prior distribution.
+Recent work (preechakul2022diffusion; zhang2022unsupervised; wang2023infodiffusion) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$.
+cohen2022diffusion crafts a diffusion bridge linking a continuously coded vector to a non-informative prior distribution.
 Diffusion for Approximate Inference
-Diffusion sampling (Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Zhang et al.
-2023; Richter, Berner, and Liu 2023; Sendera et al.
-2024; Akhound-Sadegh et al.
-2024) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function.
+Diffusion sampling (berner2022optimal; zhang2021path; vargas2023denoising; zhang2023diffusion; richter2023improved; sendera2024diffusion; akhound2024iterated) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function.
 These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift).
-Some other works (Zhang et al.
-2023; Akhound-Sadegh et al.
-2024; Sendera et al.
-2024) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration.
+Some other works (zhang2023diffusion; akhound2024iterated; sendera2024diffusion) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration.
 While this work is applicable to our setting, it does not rely on the structure of $p(\mathbf{x},\mathbf{z})$ available to us, namely tractable sampling in $p$.
 Dimensionality Reduction
-Latent variable models in general are an attractive alternative to visualization methods like PCA, Sparse PCA, NMF, UMAP, and t-SNE (Wold, Esbensen, and Geladi 1987; Kuleshov 2013; Lee and Seung 2000; McInnes, Healy, and Melville 2018; Van der Maaten and Hinton 2008).
+Latent variable models in general are an attractive alternative to visualization methods like PCA, Sparse PCA, NMF, UMAP, and t-SNE (wold1987principal; kuleshov2013fast; NIPS2000_f9d11525; mcinnes2018umap; van2008visualizing).
 Domain-specific knowledge can be injected through the prior, and deep neural networks can be utilized to achieve a more expressive mapping from the data space to the latent space.
@@
 While this paper focuses on applications of DDVI to dimensionality reduction and visualization, there exist other tasks for the algorithm, e.g., density estimation or sample quality.
-Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making (Nguyen and Grover 2022; Deshpande and Kuleshov 2023), meta-learning (Rastogi et al.
-2023), or causal effect estimation (Deshpande et al.
-2022).
+Accurate variational inference has the potential to improve downstream applications of generative modeling, e.g., decision making (DBLP:conf/icml/NguyenG22; deshpande2023calibrated), meta-learning (rastogi2023semi), or causal effect estimation (deshpande2022deep).
 Since our learning objective differs from the ELBO (it adds a regularizer), we anticipate gains on models whose training benefits from regularization, but perhaps not on all models.
@@
 1:(Optional) Pre-train $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$ and $q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x})$ with DDVI but with unconditional diffusion model $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y})$
-2:for epoch $=1,\ldots,n$ do
+2:for epoch $1,\ldots,n$ do
 3:  for $\mathbf{x}_{1},\ldots,\mathbf{x}_{k}\sim p_{\mathcal{D}}(\mathbf{x})$ do
 4:   $\mathbf{y}_{i}\sim q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x}_{i})$ and $\mathbf{z}_{i}\sim q(\mathbf{z}|\mathbf{y}_{i},\mathbf{x}_{i})$ for $i=1,\ldots,k$
-5:         Optimize ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ with respect to a Monte Carlo estimate of $\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\|p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z})\right)$ for each $\mathbf{x}_{i}$triangleright Standard ELBO training part
-6:         for iteration $=1,\ldots,m$ do $\triangleright$ Do sleep for $m$ iterations
-7:              $\mathbf{z}_{1},\ldots,\mathbf{z}_{k}\sim p(\mathbf{z})$triangleright Batch-sample latents from prior
-8:              $\hat{\mathbf{x}}_{i}\sim p(\mathbf{x}|\mathbf{z}_{i})$ for $i=1,\ldots,k$triangleright Construct fantasy inputs
-9:              $\mathbf{y}_{i}\sim r(\mathbf{y}|\mathbf{z}_{i})$ for $i=1,\ldots,k$triangleright Construct fantasy inputs
+5:   Optimize ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$ with respect to a Monte Carlo estimate of $\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\|p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z})\right)$ for each $\mathbf{x}_{i}$ $\triangleright$ Standard ELBO training part
+6:   for iteration $1,\ldots,m$ do $\triangleright$ Do sleep for $m$ iterations
+7:     $\mathbf{z}_{1},\ldots,\mathbf{z}_{k}\sim p(\mathbf{z})$ $\triangleright$ Batch-sample latents from prior
+8:     $\hat{\mathbf{x}}_{i}\sim p(\mathbf{x}|\mathbf{z}_{i})$ for $i=1,\ldots,k$ $\triangleright$ Construct fantasy inputs
+9:     $\mathbf{y}_{i}\sim r(\mathbf{y}|\mathbf{z}_{i})$ for $i=1,\ldots,k$ $\triangleright$ Construct fantasy inputs
 10:     Optimize $\boldsymbol{\phi}$} using the standard diffusion noise prediction loss on $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{i},\hat{\mathbf{x}}_{i})$
@@
 In wake-sleep, sampling $\mathbf{x}$ from $p_{\boldsymbol{\theta}}$ to obtain gradients for the sleep term introduces computational overhead.
-To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation ^$\hat{\mathcal{L}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ of $\mathcal{L}$:
-^$\hat{\mathcal{L}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ $=\underbrace{\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / reconstr.
-term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}\underbrace{-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{y},\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$}}\underbrace{-\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}))}_{\text{latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$}}.$ (12)
+To address this issue, we propose wake-sleep in latent space, an algorithm that optimizes an approximation $\hat{\mathcal{L}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ of $\mathcal{L}$:
+[…]
 We have replaced $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$ with a latent sleep term $\mathcal{L}_{\text{sleep}}(\mathbf{x},{\boldsymbol{\phi}})$, in which $\mathbf{x}$ is given, and we only seek to fit the true reverse noising process $r(\mathbf{z}|\mathbf{y})$ independently of $\mathbf{x}$.
@@
 Thus, we minimize the forward KL divergence by sampling $\mathbf{z}$, and applying the noising process to get $\mathbf{y}$; the $q_{\boldsymbol{\phi}}$ is fit to denoise $\mathbf{z}$ from $\mathbf{y}$ as in Equation 8.
-We optimize our bound on ^$\hat{\mathcal{L}}(x,{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ end-to-end using minibatch gradient descent over ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$.
+We optimize our bound on $\hat{\mathcal{L}}(x,{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ end-to-end using minibatch gradient descent over ${\boldsymbol{\theta}},{\boldsymbol{\phi}}$.
 While the wake term is a reconstruction loss as in wake-sleep, the sleep term generates latent samples $\mathbf{z},\mathbf{y}$ from $r(\mathbf{y}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ (by analogy with $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ in normal wake-sleep); the denoiser $q_{\boldsymbol{\phi}}$ is trained to recover $\mathbf{z}$ from $\mathbf{y}$.
@@
 We provide a comprehensive comparison of different methods in Table 4.
-Vahdat, Kreis, and Kautz (2021); Wehenkel and Louppe (2021); Rombach et al.
-(2022) perform diffusion in the latent space of a VAE to improve the efficiency of image generation.
+vahdat2021score; wehenkel2021diffusion; rombach2022high perform diffusion in the latent space of a VAE to improve the efficiency of image generation.
 Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors.
 Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
-Recent work (Preechakul et al.
-2022; Zhang, Zhao, and Lin 2022; Wang et al.
-2023) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$.
-Cohen et al.
-(2022) crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
+Recent work (preechakul2022diffusion; zhang2022unsupervised; wang2023infodiffusion) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$.
+cohen2022diffusion crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
 Model Training Objective Approximating Family Sample-based Prior Auxiliary Variable Tasks Simplified Graphical Illustration AEVB ELBO Diagonal Gaussian Density estimation $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ AEVB-IAF ELBO Normalizing flow Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{x}$ AAEB Adversarial training Adversarial generator Visualization $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ H-AEVB-(IAF) ELBO Factorial Normal / Normalizing flow Density estimation / High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ ADGM ELBO Non-Gaussian Density estimation $\mathbf{x}\rightarrow\mathbf{a}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ LDM ELBO Diagonal Gaussian High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ LSGM ELBO & score matching Diagonal Gaussian High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ DDVI ELBO & sleep term Denoising diffusion Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{T}(\mathbf{y})\rightarrow\mathbf{z}_{0}(\mathbf{z})\rightarrow\mathbf{x}$
@@
 Appendix E Connections to Diffusion Samplers
-Diffusion sampling (Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Zhang et al.
-2023; Richter, Berner, and Liu 2023; Sendera et al.
-2024; Akhound-Sadegh et al.
-2024) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function.
-These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) (Zhang and Chen 2021), denoising diffusion sampler (DDS) (Vargas, Grathwohl, and Doucet 2023), and time-reversed diffusion sampler (DIS) (Berner, Richter, and Ullrich 2022), which have been unified by Richter, Berner, and Liu (2023).
-Some other works (Zhang et al.
-2023; Akhound-Sadegh et al.
-2024) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration.
-Sendera et al.
-(2024) benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps.
+Diffusion sampling (berner2022optimal; zhang2021path; vargas2023denoising; zhang2023diffusion; richter2023improved; sendera2024diffusion; akhound2024iterated) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function.
+These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) (zhang2021path), denoising diffusion sampler (DDS) (vargas2023denoising), and time-reversed diffusion sampler (DIS) (berner2022optimal), which have been unified by richter2023improved.
+Some other works (zhang2023diffusion; akhound2024iterated) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration.
+sendera2024diffusion benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps.
 Overall, there are indeed some strong connections between these works and ours:
@@
 In our setting, $p_{\boldsymbol{\theta}}(z|x)$ is a moving target density, as we jointly learn $\boldsymbol{\theta}$} with $\boldsymbol{\phi}$}, as opposed to a static target density that diffusion-structured samplers are designed to solve.
-  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy (Zhang and Chen 2021) or off-policy (Malkin et al.
-2022) trajectory-based optimization.
+  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy (zhang2021path) or off-policy (malkin2022trajectory) trajectory-based optimization.
 Instead, we introduce a wake-sleep optimization algorithm and its simplified version to alleviate the weak learning signal issue and optimize the evidence lower bound in a better way.
@@
 Diffusion Samplers GFlowNet-based DDVI (ours)
-(Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Akhound-Sadegh et al.
-2024) (Zhang et al.
-2023; Richter, Berner, and Liu 2023; Sendera et al.
-2024; Malkin et al.
-2022)
+(berner2022optimal; zhang2021path; vargas2023denoising; akhound2024iterated) (zhang2023diffusion; richter2023improved; sendera2024diffusion; malkin2022trajectory)
 Tasks Sampling, density estimation Sampling, density estimation Learning, Sampling, Dimensionality reduction
@@
 Pinwheel.
-This distribution was used in (Johnson et al.
-2016).
+This distribution was used in (johnson2016svae).
 We define the number of clusters to be 10.
@@
 Swiss Roll.
-This distribution was used in (Marsland 2014).
+This distribution was used in (marsland2014swissroll).
 For semi-supervised learning experiments, this prior is partitioned into 10 partitions.
@@
 ELBO Latent NLL ELBO Latent NLL ELBO Latent NLL
-AEVB $-12.96\pm 1.81$ $3.26\pm 0.60$ $-12.87\pm 4.55$ $6.25\pm 1.58$ $-7.91\pm 0.11$ $2.91\pm 0.17$
-AEVB-IAF $-3.24\pm 0.16$ $1.71\pm 0.84$ $-4.03\pm 0.73$ $5.51\pm 0.51$ $\mathbf{-2.10\pm 0.31}$ $1.71\pm 0.77$
+AEVB $12.96\pm 1.81$ $3.26\pm 0.60$ $12.87\pm 4.55$ $6.25\pm 1.58$ $7.91\pm 0.11$ $2.91\pm 0.17$
+AEVB-IAF $3.24\pm 0.16$ $1.71\pm 0.84$ $4.03\pm 0.73$ $5.51\pm 0.51$ $\mathbf{-2.10\pm 0.31}$ $1.71\pm 0.77$
 AAEB $N/A$ $1.70\pm 0.41$ $N/A$ $\mathbf{3.18\pm 0.22}$ $N/A$ $1.67\pm 0.17$
-H-AEVB $-4.42\pm 0.46$ $\mathbf{1.69\pm 0.17}$ $-5.36\pm 0.77$ $5.74\pm 0.55$ $-2.86\pm 0.11$ $1.64\pm 0.09$
-PIS ${-2.92\pm 1.23}$ ${3.61\pm 0.62}$ $-4.14\pm 0.49$ $7.14\pm 0.14$ $-4.85\pm 0.06$ $3.91\pm 0.06$
-DDVI $\mathbf{-1.38\pm 0.44}$ ${1.75\pm 0.53}$ $\mathbf{-3.05\pm 0.65}$ ${5.66\pm 2.63}$ $-2.47\pm 0.30$ $\mathbf{1.58\pm 0.09}$
+H-AEVB $4.42\pm 0.46$ $\mathbf{1.69\pm 0.17}$ $5.36\pm 0.77$ $5.74\pm 0.55$ $2.86\pm 0.11$ $1.64\pm 0.09$
+PIS $2.92\pm 1.23$} $3.61\pm 0.62$} $4.14\pm 0.49$ $7.14\pm 0.14$ $4.85\pm 0.06$ $3.91\pm 0.06$
+DDVI $\mathbf{-1.38\pm 0.44}$ $1.75\pm 0.53$} $\mathbf{-3.05\pm 0.65}$ $5.66\pm 2.63$} $2.47\pm 0.30$ $\mathbf{1.58\pm 0.09}$
 Table 7: Unsupervised learning on CIFAR-10. We report ELBO and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
-Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $-17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $-13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $-5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $-5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $-4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\mathbf{-1.60\pm 0.29}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.09\pm 0.05}$ $\mathbf{-4.13\pm 1.51}$ $\mathbf{0.47\pm 0.09}$ $\mathbf{2.29\pm 0.08}$ $\mathbf{-1.73\pm 0.64}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.48\pm 0.02}$
+Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\mathbf{-1.60\pm 0.29}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.09\pm 0.05}$ $\mathbf{-4.13\pm 1.51}$ $\mathbf{0.47\pm 0.09}$ $\mathbf{2.29\pm 0.08}$ $\mathbf{-1.73\pm 0.64}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.48\pm 0.02}$
 Table 8: Semi-supervised learning on CIFAR-10 (10,000 labels).
 We report ELBO, accuracy using KNN (K=20) classifier (Acc), and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
-Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB -- -- -- H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\mathbf{1.27\pm 0.21}$ $\mathbf{3.86\pm 1.17}$ $\mathbf{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
+Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB - - - H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\mathbf{1.27\pm 0.21}$ $\mathbf{3.86\pm 1.17}$ $\mathbf{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
 Table 9: Unsupervised learning on MNIST, including the results of DDVI without the sleep term.
@@
 Introduce the latent variable $\mathbf{z}$ and apply the ELBO:
-[…]
-[…]
+$\log p_{\boldsymbol{\theta}}(\mathbf{x})$ $=\log\int p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})\,d\mathbf{z}$ (15)
+$\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\right]\quad\text{(by Jensen's inequality)}$ (16)
 $=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\text{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\right)$ (17)
@@
 $\bar{H}(p_{\boldsymbol{\theta}})$ is the expected conditional entropy of $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on $\boldsymbol{\phi}$}.
-This form mirrors the ELBO derivation for diffusion models (Sohl-Dickstein et al.
-2015), where each step in the Markov chain contributes a KL divergence term, and the reconstruction term arises from the connection between the noisy latent variables and the original data.
+This form mirrors the ELBO derivation for diffusion models (sohl2015deep), where each step in the Markov chain contributes a KL divergence term, and the reconstruction term arises from the connection between the noisy latent variables and the original data.